### v2

# When does a Gaussian process have its paths in a reproducing kernel Hilbert space?

Abstract
We investigate for which Gaussian processes there do or do not exist reproducing kernel Hilbert spaces (RKHSs) that contain almost all of their paths. In particular, we establish a new result that makes it possible to exclude the existence of such RKHSs in many cases. Moreover, we combine this negative result with some known techniques to establish positive results. Here it turns out that for many classical families of Gaussian processes we can fully characterize for which members of these families there exist RKHSs containing the paths. Similar characterizations are obtained for Gaussian processes, for which the RKHSs of their covariance functions are Sobolev spaces or Sobolev spaces of mixed smoothness.
1 Introduction
Given a centered Gaussian process $X:=(X_{t})_{t\in T}$ on some (complete) probability space $(\Omega,{\cal A},\mathrm{P})$, a classical task is to describe its path properties by function spaces such as spaces of continuous functions. One particular instance of this general task asks, whether the paths $X_{\bullet}$ of the process are almost surely contained in a suitable reproducing kernel Hilbert space (RKHS) $H$ on $T$. Clearly, the most obvious choice of such a space is the RKHS $H_{X}$ of the the covariance function $k_{X}(s,t):=\mathbb{E}(X_{s}X_{t})$ of $X$. In this case, a classical result, which goes back to Parzen [26], Kallianpur [16], and LePage [25] shows that
$\mathrm{P}(\{X_{\bullet}\in H_{X}\})=1$
if and only if $\dim H_{X}<\infty$. In other words, in basically no situation of interest is $H_{X}$ an RKHS containing the paths of $X$. However, $H_{X}$ is, of course, by no means the only RKHS on $T$ and beginning with Driscoll [9], and subsequently Fortet [13], more general $H$ have been considered. The most general investigation in this direction has been conducted by Lukić and Beder [22] and since their results provide a basis of our analysis, we quickly recall their findings. To this end, we need the following notion.
Definition 1.1.
Let $H_{1}$ and $H_{2}$ be RKHSs on $T$. Then we write $H_{1}\ll H_{2}$ and speak of nuclear dominance of $H_{2}$ over $H_{1}$, if $H_{1}\subset H_{2}$ and the resulting embedding map $J:H_{1}\to H_{2}$ is a Hilbert-Schmidt operator.
Note that formally, [22] defines nuclear dominance differently, see the discussion following Theorem 6.2 below, but Lemma 6.3 shows that our definition is actually equivalent to the one used in [22]. With these preparations, [22, Theorem 7.4] provides the following characterization for a fixed RKHS $H$.
Theorem 1.2.
Let $X:=(X_{t})_{t\in T}$ be a centered Gaussian process on a complete probability space $(\Omega,{\cal A},\mathrm{P})$ and $H$ be an RKHS on $T$. Then the following statements hold true:
  1. i)
    If $H_{X}\ll H$, then there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
  2. ii)
    If $H_{X}\not\ll H$, then for all versions $Y$ of $X$ we have $\mathrm{P}(\{Y_{\bullet}\in H\})=0$.
Here we note that strictly speaking [22, Theorem 7.4] only shows ii) for $Y=X$. However, if we have an arbitrary version $Y$ of $X$, then $Y$ is a centered Gaussian process with $k_{Y}=k_{X}$, which in turn gives $H_{Y}=H_{X}$. Applying the original result [22, Theorem 7.4] to $Y$ then yields ii).
Informally speaking, given a fixed “candidate” RKHS $H$, Theorem 1.2 shows that we have either
$\mathrm{P}(\{X_{\bullet}\in H\})=1\qquad\qquad\mbox{ or }\qquad\qquad\mathrm{P}(\{X_{\bullet}\in H\})=0\,,$
and in addition both situations are characterized by $H_{X}\ll H$, respectively $H_{X}\not\ll H$. In this sense, it provides a “test” for each RKHS $H$ on $T$. However, it does not tell us us how to find an $H$ with $H_{X}\ll H$, nor whether $H_{X}\ll H$ is even possible for suitable RKHSs $H$.
In this respect we note that positive, constructive results can sometimes be derived if we already have some knowledge on the paths of $X$. To illustrate this, let $E$ be a Banach space of functions (BSF) on $T$, that is, $E$ is a Banach space consisting of functions $f:T\to\mathbb{R}$ such that the point evaluations $\delta_{t}:E\to\mathbb{R}$ given by $\delta_{t}(f):=f(t)$ for all $t\in T$ and $f\in E$, are continuous. If $(\Omega,{\cal A},\mathrm{P})$ is complete and we know that there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in E\})=1$, then for every “surrounding” RKHS $H$ on $T$, that is $E\subset H$, we obviously also have $\mathrm{P}(\{Y_{\bullet}\in H\})=1$. Unfortunately, however, this technique is limited to “sufficiently benign” $E$, since there are various important BSFs $E$, for which no such surrounding $H$ exists. For example, for the somewhat natural candidate $E=C(T)$, where $C(T)$ denotes the space of continuous functions $T\to\mathbb{R}$ on a compact metric space $T$, we cannot find a surrounding $H$ as soon as $T$ is not countable, see e.g. [37]. Moreover, [34] has recently shown that even for many spaces $E$ of smoother functions no such surrounding $H$ with bounded kernel may exist. For example, if $\operatorname{H\ddot{o}l}^{\beta}([0,1])$ denotes the space of all $\beta$-Hölder continuous functions $f:[0,1]\to\mathbb{R}$, then there exists an RKHS $H$ with bounded kernel and $\operatorname{H\ddot{o}l}^{\beta}([0,1])\subset H$ if $\beta>1/2$, while there is no such RKHS if $\beta<1/2$, see [34, Section 4.5, Theorem 22, and Lemma 24, respectively Theorem 12]. In particular, the approach of taking a surrounding RKHS does not work for the Wiener process, while for the fractional Brownian motions with Hurst index $>1/2$ it does work, see also Example 5.4. In summary, the approach of using a surrounding RKHS can only be used if we already know that the paths are contained in a sufficiently small BSF $E$.
A different approach for constructing an RKHS $H$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ for some suitable version $Y$ of $X$ has been taken in [36]. In a nutshell, this approach is based on a refined analysis of the Karhunen-Loève expansion for $X$. Unfortunately, this approach, as well as its recent refinement in [17], again fails to produce positive results in many situations. In particular, it does not provide an RKHS containing the paths of the Wiener process.
The example of the Wiener process suggests that for some Gaussian processes $X$ it may be impossible to find an RKHS containing the paths of $X$. To the best of our knowledge, however, such general impossibility results, even for concrete processes, have not been established yet, despite the fact that Gaussian processes with paths in an RKHS have attracted interests in diverse areas of applied mathematics, see e.g. [30, 12, 15, 41] and the references mentioned in [17]. In particular, Theorem 1.2 cannot be directly used, since we would have to check $H_{X}\ll H$ for all RKHS $H$ on $T$, or at least for a sufficiently rich class of such RKHSs. Similarly, a failure of the described two approaches for finding an $H$ containing the paths of the considered process does, of course, not imply a general impossibility result.
The main goal of this paper is close this gap by characterizing when for a Gaussian process $X$ there exists an RKHS $H$ and a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$. To this end, we will first present a necessary condition for the existence of $H$ with $H_{X}\ll H$. Combining this necessary condition with the two approaches discussed above for finding $Y$ and $H$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ we can then characterize for various families of Gaussian processes, when such $Y$ and $H$ exist.
The rest of this paper is organized as follows: In Section 2 we collect the notions and notations that are necessary for understanding our results. The general theory is then presented in Section 3, and in Section 4 we apply our methods to Gaussian processes $X:=(X_{t})_{t\in T}$ whose RKHSs $H_{X}$ are essentially equal to a fractional Sobolev space on $T\subset\mathbb{R}^{d}$. In addition, the case of Sobolev spaces with dominating mixed smoothness is also investigated. Moreover, some concrete examples are presented in Section 5. Finally, all proofs as well as some additional material needed for these proofs can be found in Section 6.
2 Preliminaries
For sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}\subset(0,\infty)$ we write $a_{n}\sim b_{n}$ if there exist constants $c_{1},c_{2}>0$ with $c_{1}a_{n}\leq b_{n}\leq c_{2}a_{n}$ for all $n\geq 1$.
If $E$ and $F$ are Banach spaces, we write $E\hookrightarrow F$, if $E\subset F$ and the resulting embedding map is continuous. In addition, we write $E\circeq F$ if $E\subset F\subset E$, and we write $E\cong F$ if $E\circeq F$ and the norms of $E$ and $F$ are equivalent. In other words, we have $E\cong F$ if and only if $E\hookrightarrow F\hookrightarrow E$. Obviously $E\cong F$ implies $E\circeq F$, and if $E$ and $F$ are BSFs, then the converse implication is also true by a simple application of the closed graph theorem, see e.g. [34, Lemma 23]. For RKHSs $H_{1}$ and $H_{2}$ on $T$ we thus have $H_{1}\circeq H_{2}$ if and only if $H_{1}\cong H_{2}$.
Given a measurable space $(T,{\cal B})$, we write ${{\cal L}_{0}({\cal B})}$ for the space of measurable functions $T\to\mathbb{R}$, where we equip $\mathbb{R}$ with the Borel $\sigma$-algebra $\mathfrak{B}$. Analogously, ${{\cal L}_{\infty}({\cal B})}$ stands for the Banach space of all bounded, measurable functions $T\to\mathbb{R}$ equipped with the supremums-norm $\|\cdot\|_{\infty}$. Furthermore, if $\nu$ is some measure on $(T,{\cal B})$, then ${L_{2}(\nu)}$ denotes the usual Hilbert space of $\nu$-equivalence classes $[f]_{\sim}$ of square $\nu$-integrable functions $f$, and ${L_{\infty}(\nu)}$ stands for the Banach space of $\nu$-equivalence classes $[f]_{\sim}$ of measurable, $\nu$-almost surely bounded functions $f$. Note that the map
$I_{\nu}:{{\cal L}_{\infty}({\cal B})}$ $\to{L_{\infty}(\nu)}$ (1)
$f$ $\mapsto[f]_{\sim}\,,$
is linear and bounded with $\|[f]_{\sim}\|_{\infty}\leq\|f\|_{\infty}$ for all $f\in{{\cal L}_{\infty}({\cal B})}$, where $\|[f]_{\sim}\|_{\infty}$ denotes the usual essential supremum norm of $[f]_{\sim}$ in ${L_{\infty}(\nu)}$. Finally, if $\lambda^{d}$ is the Lebesgue measure on $\mathbb{R}^{d}$ and $T\subset\mathbb{R}^{d}$ is measurable with $\lambda^{d}(T)>0$, then we often write ${L_{2}(T)}:={L_{2}(\lambda^{d}_{|T})}$, where $\lambda^{d}_{|T}$ is the Lebesgue measure on $T$. Analogously, we write ${{\cal L}_{\infty}(T)}:={{\cal L}_{\infty}(\mathfrak{B}^{d}_{|T})}$, where $\mathfrak{B}^{d}_{|T}$ denotes the restriction of the Borel $\sigma$-algebra $\mathfrak{B}^{d}$ on $\mathbb{R}^{d}$ to $T$.
Let us now assume that we have an RKHS $H$ on $T$ with kernel $k$, where we refer to e.g. [3, 4], and [38, Chapter 4] for some background information on these spaces. Then $k$ is bounded if and only if all functions in $H$ are bounded, see e.g. [38, Lemma 4.23]. Moreover, if ${\cal B}$ is some $\sigma$-algebra on $T$ for which $k$ is ${\cal B}\otimes{\cal B}$-measurable, then $H\subset{{\cal L}_{0}({\cal B})}$, see e.g. [38, Lemma 4.24], and the converse implication is true for separable $H$, see e.g. [38, Lemmas 4.24 and 4.25]. Finally, let us assume that we have a finite measure space $(T,{\cal B},\nu)$ and an RKHS $H$ on $T$ with $H\subset{{\cal L}_{\infty}({\cal B})}$. As mentioned above this yields $H\hookrightarrow{{\cal L}_{\infty}({\cal B})}$, and therefore the linear map
$I_{k,\nu}:H$ $\to{L_{2}(\nu)}$ (2)
$h$ $\mapsto[h]_{\sim}$
is well-defined and continuous. Finally, we say that $\nu$ is $H$-positive, if $I_{k,\nu}$ is injective, that is, $\nu(\{h\neq 0\})>0$ for all $h\in H$ with $h\neq 0$. For example, if $H$ consists of continuous functions and $\nu(U)>0$ for all non-empty, open $U\subset T$, then $\nu$ is $H$-positive. On the other hand, if e.g. $\nu$ is an empirical measure based upon $n$ points and $\dim H=\infty$, then $I_{k,\nu}$ cannot be injective.
Recall that a probability measure $\mathrm{P}$ on $\mathbb{R}$ is a Gaussian measure, if there are $\mu\in\mathbb{R}$ and $\sigma\in[0,\infty)$ such that the characteristic function of $\mathrm{P}$ is given by
$\varphi_{\mathrm{P}}(t)=\mathrm{e}^{\mathrm{i}t\mu}\cdot\exp\biggl{(}-\frac{\sigma^{2}t^{2}}{2}\biggr{)}\,,\qquad\qquad t\in\mathbb{R}.$
Clearly, if $\sigma>0$, then $\mathrm{P}$ is a normal distribution with expectation $\mu$ and variance $\sigma^{2}$. In the case $\sigma=0$, the probability measure $\mathrm{P}$ is a point mass on $\mu$, that is $\mathrm{P}(\{\mu\})=1$.
Now let $X=(X_{t})_{t\in T}$ be an $\mathbb{R}$-valued, stochastic process on some probability space $(\Omega,{\cal A},\mathrm{P})$. As usual, we call the map $X_{\bullet}(\omega):t\mapsto X_{t}(\omega)$, where $\omega\in\Omega$ is fixed, a path of $X$. Moreover, $X$ is called a Gaussian process, if for all $t_{1},\dots,t_{n}\in T$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$, the distribution of $\alpha_{1}X_{t_{1}}+\dots+\alpha_{n}X_{t_{n}}$ is a Gaussian measure on $\mathbb{R}$. Furthermore, the Gaussian process is centered, if $\mathbb{E}X_{t}=0$ for all $t\in T$. In this case, its covariance function $k_{X}:T\times T\to\mathbb{R}$ is given by
$k_{X}(t_{1},t_{2}):=\mathbb{E}(X_{t_{1}}X_{t_{2}})\,,\qquad\qquad t_{1},t_{2}\in T.$
It is well known and easy to verify that $k_{X}$ is a symmetric and positive definite function, and thus a (reproducing) kernel. We write $H_{X}$ for the corresponding RKHS of $k_{X}$ and we further use the short hand $I_{X,\nu}:=I_{k_{X},\nu}$.
3 General Results
In this section we present and investigate necessary and sufficient conditions for the existence of RKHSs containing the paths of a given Gaussian process.
To begin with, let us assume that we have a centered Gaussian process $X:=(X_{t})_{t\in T}$ whose RKHS satisfies $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$, where ${\cal B}$ is some $\sigma$-algebra on $T$. Then $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ defined in (2) is a Hilbert-Schmidt operator for all finite measures $\nu$ on $(T,{\cal B})$, see Lemma 6.4. Our first theorem shows that $I_{X,\nu}$ is even nuclear if the paths of a suitable version $Y$ of $X$ are contained in some RKHS $H$ with bounded, measurable kernel.
Theorem 3.1.
Let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space, $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ with RKHS $H_{X}$, and ${\cal B}$ be a $\sigma$-algebra on $T$. If there exist an RKHS $H$ on $T$ with $H\subset{{\cal L}_{\infty}({\cal B})}$ and a version $Y$ of $X$ such that
$\mathrm{P}(\{Y_{\bullet}\in H\})>0$
then, for all finite measures $\nu$ on $(T,{\cal B})$, the operator $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is nuclear.
Note that there is a conceptual difference between Theorems 1.2 and 3.1. Indeed, Theorem 1.2 provides a “test” for every single candidate RKHS $H$ as discussed in the introduction. In contrast, Theorem 3.1 provides a general impossibility result: If for a given $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$ we find a finite measure $\nu$ on $(T,{\cal B})$ for which $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is not nuclear, then there exist no RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ and no version $Y$ of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
While at first glance finding such a $\nu$ may seem to be a hard task, we will see in Sections 4 and 5 that for e.g. bounded, measurable $T\subset\mathbb{R}^{d}$ the Lebesgue measure is typically a sufficient choice. In fact, we will leverage there from well-established results on the “embedding” $I_{X,\nu}:H_{X}\to{L_{2}(T)}$ if $H_{X}$ belongs to particular classes of RKHSs such as fractional Sobolev spaces.
In some other situations, however, the RKHS $H_{X}$ and the operator $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ may not be as amenable as desired. In the following, we will thus present two equivalent characterizations of $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ being nuclear, which sometimes make it easier to work with this property.
We will begin with a simple yet useful characterization that makes it possible to alter the space $H_{X}$ by some finite dimensional space. To this end, we need the following definition.
Definition 3.2.
Let $H_{1}$ and $H_{2}$ be RKHSs on $T$. Then we say that $H_{1}$ and $H_{2}$ are essentially equivalent, if there exist RKHSs $H_{1}^{\star}$ and $H_{2}^{\star}$ on $T$ with $\dim H_{1}^{\star}<\infty$, $\dim H_{2}^{\star}<\infty$, and
$H_{1}+H_{1}^{\star}\circeq H_{2}+H_{2}^{\star}\,.$ (3)
It is straightforward to check that the notion of being essentially equivalent gives an equivalence relation on the set of all RKHSs on $T$, but in the following we do not need this. In addition, another simple calculation shows that if $H_{1}$ and $H_{2}$ are essentially equivalent, then both $H_{1}$ and $H_{2}$ are also essentially equivalent to $H_{1}+H_{2}$. Furthermore, RKHSs $H_{1}$ and $H_{2}$ with $H_{1}\circeq H_{2}$ are obviously also essentially equivalent. Finally, the following lemma shows that in some situations the essential equivalence can be easily verified with the help of the kernels of the involved RKHSs.
Lemma 3.3.
Let $H_{1}$ and $H_{2}$ be RKHSs on $T$ with kernels $k_{1}$ and $k_{2}$, and $k_{1}^{\star}$ and $k_{2}^{\star}$ be kernels on $T$ with finite dimensional RKHSs. Then $H_{1}$ and $H_{2}$ are essentially equivalent, if we have
$k_{1}-k_{2}=k_{2}^{\star}-k_{1}^{\star}\,.$
Now, the following result shows that $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is nuclear if and only if $I_{H,\nu}:H\to{L_{2}(\nu)}$ is nuclear for all RKHS $H$ that are essentially equivalent to $H_{X}$. This will make it possible to switch from $H_{X}$ to an essentially equivalent $H$ for which $I_{H,\nu}:H\to{L_{2}(\nu)}$ is better understood.
Theorem 3.4.
Let $(T,{\cal B},\nu)$ be a finite measure space and $H_{1}$ and $H_{2}$ be essentially equivalent RKHSs on $T$ with $H_{1},H_{2}\subset{{\cal L}_{\infty}({\cal B})}$. Then $I_{H_{1},\nu}:H_{1}\to{L_{2}(\nu)}$ is nuclear if and only if $I_{H_{2},\nu}:H_{2}\to{L_{2}(\nu)}$ is nuclear.
Theorem 3.4 can be used to establish non-existence results with the help of Theorem 3.1, see e.g. Example 5.2. Combining the next result with Theorem 1.2 shows that in some situations the essential equivalence of RKHSs can also be used to find RKHSs containing the paths of a given Gaussian process, see e.g. Theorem 4.1.
Theorem 3.5.
Let $X:=(X_{t})_{t\in T}$ be a centered Gaussian process on a probability space $(\Omega,{\cal A},\mathrm{P})$ and $H$ be an RKHS that is essentially equivalent to $H_{X}$. If there exists an RKHS $H^{\dagger}$ on $T$ with $H\ll H^{\dagger}$, then there also exists an RKHS $H_{X}^{\dagger}$ on $T$ that is essentially equivalent to $H^{\dagger}$ and satisfies both $H_{X}^{\dagger}\subset H_{X}+H^{\dagger}$ and
$H_{X}\ll H_{X}^{\dagger}\,.$
Our next goal is to present another equivalent way to express that $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is nuclear as well as to construct an $H$ with $H_{X}\ll H$ in some situations. To this end, we assume $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$, where we note that in the situation of Theorem 3.1, we do have $H_{X}\subset H\subset{{\cal L}_{\infty}({\cal B})}$ by Theorem 1.2, and therefore the assumption $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$ fits to our previous results. For a given, finite measure $\nu$ on $(T,{\cal B})$, let us now consider the operator $T_{X,\nu}:=I_{X,\nu}\circ I_{X,\nu}^{*}:{L_{2}(\nu)}\to{L_{2}(\nu)}$, where $I_{X,\nu}^{*}:{L_{2}(\nu)}\to H_{X}$ denotes the adjoint of $I_{X,\nu}$. Then some elementary calculations, see Lemma 6.4, show that $T_{X,\nu}$ is given by
$T_{X,\nu}f=\int_{T}k_{X}(\,\cdot\,,t)f(t)\,\mathrm{d}\nu(t)\,,\qquad\qquad f\in{L_{2}(\nu)},$
where the right-hand side is viewed as a $\nu$-equivalence class in ${L_{2}(\nu)}$. By construction, $T_{X,\nu}$ is positive and self-adjoint, and it is also compact, since $I_{X,\nu}$ is Hilbert-Schmidt by our assumption $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$, see again see Lemma 6.4. Consequently, the spectral theorem, see Section 6.1, shows that $T_{X,\nu}$ has at most countably many non-zero eigenvalues $(\mu_{i}(T_{X,\nu}))_{i\in I}$, where we have set either $I:=\{1,\dots,n\}$ or $I:=\mathbb{N}$, and where we have included the geometric multiplicities. Moreover, these eigenvalues are positive and there exists an orthonormal system (ONS) $([f_{i}]_{\sim})_{i\in I}\subset{L_{2}(\nu)}$ such that each $[f_{i}]_{\sim}$ is an eigenfunction of $T_{X,\nu}$ with eigenvalue $\mu_{i}:=\mu_{i}(T_{X,\nu})$. We now define
$e_{i}:=\mu_{i}^{-1}I_{X,\nu}^{*}[f_{i}]_{\sim}\,,\qquad\qquad i\in I.$ (4)
This gives $e_{i}\in H_{X}$ and $[e_{i}]_{\sim}=I_{X,\nu}e_{i}=\mu_{i}^{-1}T_{X,\nu}[f_{i}]_{\sim}=[f_{i}]_{\sim}$ for all $i\in I$. In other words, we may choose representatives $e_{i}$ of $[f_{i}]_{\sim}$ that are contained in $H$. Finally, $T_{X,\nu}$ is nuclear since $I_{X,\nu}$ is Hilbert-Schmidt, see Lemma 6.1, and therefore $(\mu_{i}(T_{X,\nu}))_{i\in I}$ is summable.
Now recall from [39] that in some cases the eigenfunctions and -values can be used to construct RKHSs surrounding $H_{X}$. To briefly describe this construction, let us fix a $\beta>0$ for which we have
$\sum_{i\in I}\mu_{i}^{\beta}\,e_{i}^{2}(t)<\infty\,,\qquad\qquad t\in T.$ (5)
Then [39, Lemma 2.6] shows that
$k_{X}^{\beta}(t,t^{\prime}):=\sum_{i\in I}\mu_{i}^{\beta}e_{i}(t)e_{i}(t^{\prime})\,,\qquad\qquad t,t^{\prime}\in T$ (6)
defines a kernel on $T$ whose RKHS is given by
$H_{X}^{\beta}:=\biggl{\{}\sum_{i\in I}a_{i}\mu_{i}^{\beta/2}e_{i}:(a_{i})_{i\in I}\in\ell_{2}(I)\biggr{\}}\,,$
where the series converges pointwise and $\ell_{2}(I)$ denotes the space of all 2-summable families $(a_{i})_{i\in I}\subset\mathbb{R}$. Moreover, $(\mu_{i}^{\beta/2}e_{i})_{i\in I}$ is an ONB of $H_{X}^{\beta}$ and we have $H_{X}^{\beta}\subset{{\cal L}_{0}({\cal B})}$. Obviously, Condition (5) is monotone in the sense that if it holds for some $\beta$ then it also holds for all exponents $\alpha>\beta$. In addition, for such pairs we have $H_{X}^{\alpha}\subset H_{X}^{\beta}$. Finally, if $\nu$ is $H$-positive, then $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ is an ONB of $H_{X}$, see [39, Theorem 3.1], and thus we have $H_{X}^{1}=H_{X}$. Consequently, if in this case we have found a $\beta\in(0,1)$ with (5) then $H_{X}\subset H_{X}^{\beta}$.
With these preparations we can now present our last result of this section, where we note that ii) essentially simplifies similar results of [36, Section 5].
Theorem 3.6.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,{\cal B},\nu)$ be a finite measure space, and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ such that $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$. Moreover, let $(\mu_{i})_{i\in I}$ be the non-zero eigenvalues of $T_{X,\nu}$ and $(e_{i})_{i\in I}\subset H_{X}$ be a family such that $([e_{i}]_{\sim})_{i\in I}\subset{L_{2}(\nu)}$ is an ONS in ${L_{2}(\nu)}$ and each $[e_{i}]_{\sim}$ is an eigenfunction of $T_{X,\nu}$ to the eigenvalue $\mu_{i}$. Then the following statements hold true:
  1. i)
    The operator $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is nuclear if and only if
    $\sum_{i\in I}\sqrt{\mu_{i}}<\infty.$ (7)
  2. ii)
    Assume that $\nu$ is $H_{X}$-positive and we have a $\beta\in(0,1)$ satisfying (5). Then $H_{X}\ll H_{X}^{\beta}$ if and only if
    $\sum_{i\in I}\mu_{i}^{1-\beta}<\infty.$ (8)
    In this case, there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H_{X}^{\beta}\})=1$.
To illustrate Theorem 3.6 we assume for a moment that $\nu$ is $H_{X}$-positive and $\sup_{i\in I}\|e_{i}\|_{\infty}<\infty$. Then, if (7) is not satisfied, Theorem 3.6 shows that $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is not nuclear, and therefore Theorem 3.1 shows that there exist no RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ and no version $Y$ of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
Conversely, if (7) is satisfied, then both (5) and (8) hold true for $\beta:=1/2$, and hence there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H_{X}^{1/2}\})=1$. Furthermore, note that in this case $k_{X}^{1/2}$ is bounded, and hence we have $H_{X}^{1/2}\subset{{\cal L}_{\infty}({\cal B})}$. In other words, if we have a $H_{X}$-positive, finite measure $\nu$ with $\sup_{i\in I}\|e_{i}\|_{\infty}<\infty$, then (7) characterizes the existence of an RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ containing the paths of a suitable version of $X$. We refer to Example 5.7 for such a situation.
Finally, recall from [39, Theorem 4.6] that the spaces $H_{X}^{\beta}$ can in some sense be identified with the interpolation spaces $[{L_{2}(\nu)},[H_{X}]_{\sim}]_{\beta,2}$ of the real method. In particular, if under some mild technical assumptions we have $[{L_{2}(\nu)},[H_{X}]_{\sim}]_{1/2,2}\hookrightarrow{L_{\infty}(\nu)}$, then $H_{X}^{1/2}$ exists and is an RKHS version of $[{L_{2}(\nu)},[H_{X}]_{\sim}]_{1/2,2}$, see [36, Theorem 5.5] in combination with [39, (36) and Theorem 4.6], while [39, Theorem 5.3] ensures (7). In such situations it may thus be possible to construct $H_{X}^{1/2}$ modulo equivalent norms without using the eigenfunctions and -values. In this respect note that for certain kernels $k_{X}$ [12] presents another way to find $H_{X}^{1/2}$ without the eigenfunctions and -values.
Finally, [36, Section 1] lists various articles in which the eigenfunctions and -values are computed for specific Gaussian processes. For such processes, we can then apply Theorem 3.6 directly.
4 Results for RKHSs of Sobolev Type
In this section we investigate how our main results can be used to investigate Gaussian processes whose RKHSs are essentially equivalent to a fractional Sobolev space, or equal a Sobolev space of dominating mixed smoothness. To this end, recall that for $s\in(0,\infty)$, the fractional Sobolev spaces, or Bessel potential spaces, are defined by
$H^{s}(\mathbb{R}^{d}):=\bigl{\{}f\in{L_{2}(\mathbb{R}^{d})}:(1+\|\cdot\|^{2})^{s/2}\hat{f}\in{L_{2}(\mathbb{R}^{d})}\bigr{\}}\,,$
where $\hat{f}$ is the Fourier transformation of $f$, see e.g. [23, Section 3.1] or [42, Section 2.2.2]. By setting
$\|f\|_{H^{s}(\mathbb{R}^{d})}:=\bigl{\|}\,(1+\|\cdot\|^{2})^{s/2}\hat{f}\,\bigr{\|}_{{L_{2}(\mathbb{R}^{d})}}$
the space $H^{s}(\mathbb{R}^{d})$ becomes a Hilbert space and for $m\in\mathbb{N}$ we have $H^{m}(\mathbb{R}^{d})\cong W^{m}_{2}(\mathbb{R}^{d})$, where $W^{m}_{2}(\mathbb{R}^{d})$ denotes the classical Sobolev space with smoothness $m$, see e.g. [42, (3) in Chapter 2.5.6] or [1, Chapter 7, Paragraph 7.62]. Moreover, the definition of the fractional Sobolev spaces immediately ensures $H^{s}(\mathbb{R}^{d})\subset H^{r}(\mathbb{R}^{d})$ for all $0<r<s<\infty$, and for $s>d/2$ we have
$H^{s}(\mathbb{R}^{d})\hookrightarrow C_{b}(\mathbb{R}^{d})$ (9)
by Sobolev’s embedding theorem, see the beginning of Section 6.6 for details and references. In other words, in the case $s>d/2$ the space $H^{s}(\mathbb{R}^{d})$ can be viewed as an RKHS with bounded and continuous kernel. Finally, for bounded, measurable $T\subset\mathbb{R}^{d}$ and $s>d/2$ we define $H^{s}(T):=\{f_{|T}:f\in H^{s}(\mathbb{R}^{d})\}$ and equip this space with the norm
$\|h\|_{H^{s}(T)}:=\inf\bigl{\{}\|f\|_{H^{s}(\mathbb{R}^{d})}:f\in H^{s}(\mathbb{R}^{d})\mbox{ with }f_{|T}=h\bigr{\}}\,.$ (10)
Here, we followed [10, Chapter 2.5.1] and in particular we refer to the discussion there, when such restricted spaces also have an intrinsic description. In any case, as inclusions between BSFs $E$ and $F$ on some set $T_{0}$ remain unchanged when both spaces are restricted to some subset $T\subset T_{0}$, see e.g. [34, Proposition 3] for a general statement, the inclusions above lead to $H^{s}(T)\hookrightarrow H^{r}(T)\hookrightarrow C_{b}(T)\hookrightarrow{{\cal L}_{\infty}(T)}$ for all $s>r>d/2$. In this case, both $H^{s}(T)$ and $H^{r}(T)$ can thus be viewed to be RKHSs.
With these preparations, we can now exactly describe those Gaussian processes that have an RKHS that is essentially equivalent to a fractional Sobolev space and whose paths are contained in some RKHS.
Theorem 4.1.
Let $T\subset\mathbb{R}^{d}$ be bounded and measurable with non-empty interior. Moreover, let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ whose RKHS $H_{X}$ is essentially equivalent to $H^{s}(T)$ for some $s>d/2$. Then the following statements are equivalent:
  1. i)
    We have $s>d$.
  2. ii)
    There exist an RKHS $H\subset{{\cal L}_{\infty}(T)}$ and a version $Y$ of $X$ such that $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
In this case, for all $r\in(d/2,s-d/2)$ there actually exists such an $H$ that is essentially equivalent to $H^{r}(T)$.
As the implication from ii) to i) is based upon Theorem 3.1 we see that the necessary condition derived there is sharp in the setup of Theorem 4.1. Below in Example 5.6 we apply Theorem 4.1 to the class of Matérn kernels that are popular in many applications, see e.g. [32, 31].
Theorem 4.1 shows that we require a strong, dimension dependent regularity to ensure that the paths are contained in an RKHS. Interestingly, such a dimension dependent regularity disappears if we switch to (tensor) product kernels, which are used in e.g. efficient implementations of Gaussian processes for machine learning [14, 11]. and whose RKHSs
Theorem 4.2.
Let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space, $T=(0,1)^{d}$, and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ whose covariance function is of the form
$k_{X}(t,t^{\prime})=\prod_{i=1}^{d}k_{0}(t_{i},t_{i}^{\prime})\,,\qquad\qquad t,t^{\prime}\in T,$
where $k_{0}:(0,1)\times(0,1)\to\mathbb{R}$ is a kernel with RKHS $H_{0}\circeq H^{s}(0,1)$ for some $s>1/2$. Then the following statements are equivalent:
  1. i)
    We have $s>1$.
  2. ii)
    There exist an RKHS $H\subset{{\cal L}_{\infty}(T)}$ and a version $Y$ of $X$ such that $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
In this case, we may choose a Sobolev space of dominating mixed smoothness $H\cong B^{r}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{r}_{2,2}(0,1)$ for any $r\in(1/2,s-1/2)$.
5 Some Concrete Examples
In this section we present some explicit examples of Gaussian processes that do or do not have their paths in a suitable RKHS. To this end, we assume throughout this section that the considered Gaussian processes are defined on some complete probability space $(\Omega,{\cal A},\mathrm{P})$. Moreover, $\lambda$ and $\lambda^{d}$ denote the Lebesgue measures on the considered subsets $T\subset\mathbb{R}$, respectively $T\subset\mathbb{R}^{d}$.
With these preparations, our first example considers the Wiener process.
Example 5.1.
Let $T:=[0,1]$ and $W:=(W_{t})_{t\in T}$ be the Wiener process. Then $I_{W,\lambda}:H_{W}\to{L_{2}(T)}$ is not nuclear and consequently, there exists no RKHS $H\subset{{\cal L}_{\infty}(T)}$ that contains almost all paths of the Wiener process.
The proof of Example 5.1 shows that $I_{W,\lambda}:H_{W}\to{L_{2}(T)}$ fails to be nuclear by a thin margin. This observation will be further investigated in Example 5.4, which considers the fractional Brownian motion.
Since the Brownian bridge $(B_{t})_{t\in T}$ on $T:=[0,1]$ can be defined by $B_{t}:=(1-t)W_{t}$, it is not surprising that the negative result of Example 5.1 carries over to the Brownian bridge. The details are formulated in the following example, whose proof also illustrates the usefulness of the notion of essentially equivalent RKHSs.
Example 5.2.
Let $T:=[0,1]$. Then there exists no RKHS $H\subset{{\cal L}_{\infty}(T)}$ that contains almost all paths of the Brownian bridge.
Like the Brownian bridge, the Ornstein-Uhlenbeck process is closely related to the Wiener process, and hence it is again not surprising that the negative result of Example 5.1 carries over to the Ornstein-Uhlenbeck process. The following example presents the details for two variants of this process.
Example 5.3.
Let $T:=[0,1]$ and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process whose covariance function is given by either
$k_{X}^{(1)}(t_{1},t_{2})$ $:=a\exp(-\sigma|t_{1}-t_{2}|)\,,$
or
$k_{X}^{(2)}(t_{1},t_{2})$ $:=a\exp(-\sigma|t_{1}-t_{2}|)-a\exp(-\sigma(t_{1}+t_{2}))\,,$
where in both cases $a,\sigma>0$ are some constants. Then there exist no RKHS $H\subset{{\cal L}_{\infty}(T)}$ and no version $Y$ of $X$ such that $\mathrm{P}(\{Y_{\bullet}\in H\})>0$.
Our next example generalizes Example 5.1 to the fractional Brownian motion. Here it turns out that the results depend on the Hurst parameter.
Example 5.4.
Let $T:=[0,1]$ and $\alpha\in(0,1)$. Moreover, let $B^{(\alpha)}=(B^{(\alpha)}_{t})_{t\in T}$ be the fractional Brownian motion with Hurst index $\alpha$, that is, the centered Gaussian process with continuous paths and covariance function
$k_{B^{(\alpha)}}(t_{1},t_{2}):=\frac{1}{2}\bigl{(}|s|^{2\alpha}+|t|^{2\alpha}-|t-s|^{2\alpha}\bigr{)}\,,\qquad\qquad t_{1},t_{2}\in[0,1]\,.$
Then there exists an RKHS $H\subset{{\cal L}_{\infty}(T)}$ with $\mathrm{P}(\{B_{\bullet}^{(\alpha)}\in H\})=1$ if and only if $\alpha>1/2$.
More precisely, in the case $\alpha\leq 1/2$, the map $I_{H_{B^{(\alpha)}},\lambda}:H_{B^{(\alpha)}}\to{L_{2}(T)}$ is not nuclear and in the case $\alpha>1/2$ we may choose $H=H^{s}(T)$ for any $s\in(1/2,\alpha)$.
The fractional Brownian motion is closely related to the Riemann-Liouville process, see e.g. [21, Example 3.4] or [27], which is considered in the following example.
Example 5.5.
Let $T:=[0,1]$ and $\alpha\in(0,1)$. Moreover, let $R^{(\alpha)}=(R^{(\alpha)}_{t})_{t\in T}$ be the Riemann-Liouville process with index $\alpha$, that is, the centered Gaussian process with continuous paths given by the Itô-Integral
$R^{(\alpha)}_{t}:=\frac{1}{\Gamma{(\alpha+1/2})}\int_{0}^{t}(t-s)^{\alpha-1/2}\,\mathrm{d}W_{s}\,,\qquad\qquad t\in[0,1].$
Then there exists an RKHS $H\subset{{\cal L}_{\infty}(T)}$ with $\mathrm{P}(\{R_{\bullet}^{(\alpha)}\in H\})=1$ if and only if $\alpha>1/2$. Moreover, in the case $\alpha>1/2$ we may choose $H=H^{s}(T)$ for any $s\in(1/2,\alpha)$.
The next example, which considers Gaussian processes on a higher dimensional $T$, uses our result for RKHSs of Sobolev type formulated in Theorem 4.1. Analogous results can thus be obtained for other kernels of Sobolev type, such as the Wendland kernels, which are often attractive from a computational point of view, see [43, Chapter 9].
Example 5.6.
Let $T\subset\mathbb{R}^{d}$ be bounded and measurable with non-empty interior. Moreover, let $X:=(X_{t})_{t\in T}$ be a centered Gaussian process whose covariance function is a Matérn kernel of order $\alpha>0$ and width $\sigma>0$, that is
$k_{X}(t_{1},t_{2})=\bigl{(}\sigma\|t_{1}-t_{2}\|\bigr{)}^{\alpha}h_{\alpha}\bigl{(}\sigma\|t_{1}-t_{2}\|\bigr{)}\,,\qquad\qquad t_{1},t_{2}\in T,$
where $h_{\alpha}$ denotes the modified Bessel function of the second type of order $\alpha$. Then there exists an RKHS $H\subset{{\cal L}_{\infty}(T)}$ and a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ if and only if $\alpha>d/2$. Moreover, in the case $\alpha>d/2$ we may choose $H=H^{r}(T)$ for any $r\in(d/2,\alpha)$.
Our last example considers stationary processes on compact Abelian groups $T$ with Haar measure $\lambda$. It mostly relies on the fact that for translation invariant, $\mathbb{C}$-valued kernels the characters of $T$ form on ONB of ${L_{2}(\lambda,\mathbb{C})}$ consisting of eigenvectors of $T_{k,\lambda}$.
Example 5.7.
Let $(T,+)$ be a compact Abelian group with Haar measure $\lambda$ and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process whose covariance function $k_{X}$ is measurable and satisfies
$k_{X}(t_{1},t_{2})=k_{X}(t_{1}-t_{2},0)\,,\qquad\qquad t_{1},t_{2}\in T.$
Then there exist an RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ and a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ if and only if
$\sum_{i\in I}\sqrt{\mu_{i}(T_{X,\lambda})}<\infty.$
Moreover, in this case we may choose $H=H_{X}^{1/2}$ and we have $H_{X}^{1/2}\subset C(T)$.
Since Example 5.7 is an application of Theorem 3.6, it can be easily extended to situations in which the spectral properties of $k_{X}$ are well understood in the sense of the discussion following Theorem 3.6. Such situations include e.g. isotropic Gaussian random fields on $\mathbb{S}^{2}$, and with some more work also on $\mathbb{S}^{d}$, if $\lambda$ is the surface measure. We refer to [20] for the details on the required spectral properties.
6 Proofs
6.1 Hilbert-Schmidt and Nuclear Operators
In this subsection we briefly recall some aspects of compact operators between Hilbert spaces. To this end, let $H$ be a Hilbert space and $R:H\to H$ be a compact, positive, and self-adjoint operator. Then the spectral theorem for such operators, e.g. [18, Chapter V.2.3] or [44, Theorem VI.3.2], shows that we have at most countably many non-zero eigenvalues $(\mu_{i}(R))_{i\in I}$, where we have set either $I:=\{1,\dots,n\}$ or $I:=\mathbb{N}$, and where we have included the geometric multiplicities. Moreover, there exists an ONS $(e_{i})_{i\in I}$ in $H$ such that $e_{i}$ is an eigenvector of $R$ to the eigenvalue $\mu_{i}(R)$ for all $i\in I$. Finally, these eigenvalues are positive and in the case $I=\mathbb{N}$ they can be ordered to form a decreasing sequence converging to 00. In the following, we always assume a decreasing order.
If $S:H_{1}\to H_{2}$ is a compact operator between two Hilbert spaces $H_{1}$ and $H_{2}$, we write $S^{*}:H_{2}\to H_{1}$ for the adjoint of $S$. Then the operator $S\circ S^{*}:H_{2}\to H_{2}$ is compact, positive, and self-adjoint, and hence we can apply the spectral theorem to $S\circ S^{*}$. By doing so, we can define the singular numbers of $S$ by
$s_{i}(S):=\sqrt{\mu_{i}(S\circ S^{*})}\,,\qquad\qquad i\in I,$
and $s_{i}(S):=0$ if $I:=\{1,\dots,n\}$ and $i>n$. Recall that for compact, positive, and self-adjoint $R:H\to H$ we have
$s_{i}(R)=\sqrt{\mu_{i}(R\circ R^{*})}=\sqrt{\mu_{i}(R^{2})}=\mu_{i}(R)\,,$ (11)
where the last identity directly follows from the spectral series representation, see e.g. [18, Equation (2.21) in Chapter V.2.3] or [44, Theorem VI.3.2]. For compact $S:H_{1}\to H_{2}$ this implies
$s_{i}^{2}(S)=\mu_{i}(S\circ S^{*})=s_{i}(S\circ S^{*})\,,$ (12)
and in addition, there exist ONSs $(e_{i})_{i\in I}$ in $H_{1}$ and $(f_{i})_{i\in I}$ in $H_{2}$ with
$Sx=\sum_{i\in I}s_{i}(S)\langle x,e_{i}\rangle_{H_{1}}f_{i}\,,\qquad\qquad x\in H_{1},$ (13)
see e.g. [18, Equation (2.23) in Chapter V.2.3] or [44, Theorem VI.3.6]. Finally, recall that a compact $S:H_{1}\to H_{2}$ is nuclear if and only if we have
$\sum_{i\in I}s_{i}(S)<\infty\,,$ (14)
see e.g. [5, Chapter 11.2]. Moreover, by definition $S$ is a Hilbert-Schmidt operator if $\sum_{j\in J}\|Se_{j}\|_{H_{2}}^{2}<\infty$ for some (and then all) ONB(s) $(e_{j})_{j\in J}$ of $H_{1}$, or equivalently, if and only if
$\sum_{i\in I}s_{i}^{2}(S)<\infty\,,$ (15)
see e.g. [5, Chapter 11.3, Theorem 1]. By (12) we then immediately obtain the following elementary yet useful lemma.
Lemma 6.1.
Let $S:H_{1}\to H_{2}$ be a compact operator between two Hilbert spaces $H_{1}$ and $H_{2}$. Then $S$ is a Hilbert-Schmidt operator if and only if $S\circ S^{*}$ is nuclear.
Finally, it is sometimes helpful to compare the singular numbers with some other $s$-number scales in the sense of [29, Chapter 2]. To this end, let us first recall that for a given bounded linear operator $S:E\to F$ between two Banach spaces $E$ and $F$, the $n$-th approximation number is defined by
$a_{n}(S):=\inf\bigl{\{}\|S-A\|\,\bigl{|}\,A:E\to F\mbox{ boundedand linear with }\operatorname{rank}A<n\bigr{\}}\,.$ (16)
Furthermore, the $n$-th Kolmogorov number is defined by
$d_{n}(S):=\inf\bigl{\{}\varepsilon>0:\exists F_{\varepsilon}\mbox{ subspace of $F$ with $\dim F_{\varepsilon}<n$ and }SB_{E}\subset F_{\varepsilon}+\varepsilon B_{F}\bigr{\}}\,,$ (17)
where, $B_{E}$ and $B_{F}$ denote the closed unit balls in $E$ and $F$ and we followed the exposition in [6, Chapter 2.2]. It is well known that the approximation and Kolmogorov numbers form an $s$-scale in the sense of [29, Chapter 2.2.1]. We refer to [29, Chapters 2.3 and 2.5] and [6, Chapters 2.1 and 2.2] for corresponding results. In addition, the Kolmogorov numbers are a surjective $s$-scale, see e.g. [29, Chapter 2.5]. Moreover, $S:E\to F$ is compact if and only if $d_{n}(S)\to 0$, see e.g. [6, page 49]. Furthermore, for operators $S:H_{1}\to H_{2}$ between Hilbert spaces we have $a_{n}(S)=d_{n}(S)$, see e.g. [28, Theorem 11.3.4], and if $S:H_{1}\to H_{2}$ is compact we even have
$s_{n}(S)=a_{n}(S)=d_{n}(S)\,,\qquad\qquad n\geq 1,$ (18)
by [28, Proposition 11.3.3] as (13) is a so-called Schmidt representation in the sense of [28, D.3.2]. In other words, we may freely switch between these quantities as soon as we know that $S$ is compact.
6.2 Nuclear Dominance
In this subsection we show that the notion of nuclear dominance used in [22] is equivalent to ours. To this end, we begin with recalling [22, Theorem 1.1].
Theorem 6.2.
Let $H_{1}$ and $H_{2}$ be RKHSs over $T$ with $H_{1}\subset H_{2}$. Then the resulting embedding map
$J:H_{1}$ $\to H_{2}$ (19)
$h_{1}$ $\mapsto h_{1}$
is linear and bounded. Moreover, there exists a unique linear $D_{H_{1},H_{2}}:H_{2}\to H_{2}$ that satisfies both $\operatorname{ran}D_{H_{1},H_{2}}\subset H_{1}$ and
$\langle h_{2},h_{1}\rangle_{H_{2}}=\langle D_{H_{1},H_{2}}h_{2},h_{1}\rangle_{H_{1}}\,,\qquad\qquad h_{1}\in H_{1},h_{2}\in H_{2}.$ (20)
In addition, $D_{H_{1},H_{2}}$ is bounded, positive and self-adjoint.
Recall that [22] speaks of nuclear dominance if $D_{H_{1},H_{2}}$ is nuclear. The following simple lemma shows that this notion is equivalent to our Definition 1.1.
Lemma 6.3.
Let $H_{1}$ and $H_{2}$ be RKHSs over $T$ with $H_{1}\subset H_{2}$ and $J:H_{1}\to H_{2}$ be the corresponding embedding map (19). Then we have
$D_{H_{1},H_{2}}=J\circ J^{*}\,.$
In particular, $D_{H_{1},H_{2}}$ is nuclear if and only if $J$ is a Hilbert-Schmidt operator.
Proof of Lemma 6.3: Let us write $R:=J\circ J^{*}:H_{2}\to H_{2}$. By construction, its range satisfies $\operatorname{ran}R\subset H_{1}$. Moreover, for $h_{1}\in H_{1}$ and $h_{2}\in H_{2}$ we have
$\langle h_{2},h_{1}\rangle_{H_{2}}=\langle h_{2},Jh_{1}\rangle_{H_{2}}=\langle J^{*}h_{2},h_{1}\rangle_{H_{1}}=\langle(J\circ J^{*})h_{2},h_{1}\rangle_{H_{1}}=\langle Rh_{2},h_{1}\rangle_{H_{1}}\,.$
Consequently, $R$ satisfies (20), and by uniqueness stated in Theorem 20 we can thus conclude $R=D_{H_{1},H_{2}}$. The last assertion is a direct consequence of Lemma 6.1 and $D_{H_{1},H_{2}}=J\circ J^{*}$. ∎
6.3 Proof of Theorem 3.1
Before we can prove Theorem 3.1 we need to establish an auxiliary result investigating the operator $I_{k,\nu}:H\to{L_{2}(\nu)}$ in more detail.
Lemma 6.4.
Let $k:T\times T\to\mathbb{R}$ be a kernel with RKHS $H$ and ${\cal B}$ be a $\sigma$-algebra on $T$. If we have $H\subset{{\cal L}_{\infty}({\cal B})}$, then for every finite measure $\nu$ on $(T,{\cal B})$, the “embedding” operator
$I_{k,\nu}:H$ $\to{L_{2}(\nu)}$ (21)
$h$ $\mapsto[h]_{\sim}$
is well-defined, linear, and Hilbert-Schmidt. Moreover, its adjoint $I^{*}_{k,\nu}:{L_{2}(\nu)}\to H$ is given by
$(I_{k,\nu}^{*}f)(s)=\int_{T}k(s,t)f(t)\,\mathrm{d}\nu(t)\,,\qquad\qquad f\in{L_{2}(\nu)},\,s\in T.$
Proof of Lemma 6.4: Since we have assumed $H\subset{{\cal L}_{\infty}({\cal B})}$, we know by e.g. [38, Lemma 4.23] that the resulting embedding map $J:H\to{{\cal L}_{\infty}({\cal B})}$ is bounded. Let us define $I_{k,\nu,\infty}:H\to{L_{\infty}(\nu)}$ by $I_{k,\nu,\infty}:=I_{\nu}\circ J$, where $I_{\nu}:{{\cal L}_{\infty}({\cal B})}\to{L_{\infty}(\nu)}$ is given by (1). Then $I_{k,\nu,\infty}$ is bounded and linear, and since $\nu$ is finite, we also have the continuous embedding $J:{L_{\infty}(\nu)}\hookrightarrow{L_{2}(\nu)}$. Consequently, the linear map (21) enjoys the factorization
[…]
By [8, Corollary 4.12 and Theorem 3.2] we can thus conclude that $I_{k,\nu}$ is a Hilbert-Schmidt operator.
Let us finally derive the formula for the adjoint operator. To this end, we fix an $f\in{L_{2}(\nu)}$ and an $s\in T$. We then obtain
$(I_{k,\nu}^{*}f)(s)=\bigl{\langle}I_{k,\nu}^{*}f,k(s,\,\cdot\,)\bigr{\rangle}_{H}=\bigl{\langle}f,I_{k,\nu}(k(s,\,\cdot\,))\bigr{\rangle}_{{L_{2}(\nu)}}=\bigl{\langle}f,[k(s,\,\cdot\,)]_{\sim}\bigr{\rangle}_{{L_{2}(\nu)}}=\int_{T}k(s,t)f(t)\,\mathrm{d}\nu(t)\,,$
where in the first step we used the reproducing property $k$. ∎
Proof of Theorem 3.1: We first observe that we have $H_{X}\ll H$: To this end, we assume that $H_{X}\not\ll H$ was true. Then we would have
$\mathrm{P}(\{Y_{\bullet}\in H\})=0$
by Theorem 1.2, which contradicts our assumption on $H$ and $Y$. Consequently, the embedding map $J:H_{X}\to H$ is a Hilbert-Schmidt operator. Moreover, we have the natural factorization
[…]
where $k$ is the kernel of $H$ and $I_{k,\nu}$ is defined by (21). Now, $H\subset{{\cal L}_{\infty}({\cal B})}$ in combination with Lemma 6.4 ensures that $I_{k,\nu}$ is also a Hilbert-Schmidt operator. Consequently, both factors $J$ and $I_{k,\nu}$ in the diagram above are Hilbert-Schmidt operators, and therefore their product $I_{X,\nu}$ is nuclear, see e.g. [5, Chapter 11.3, Theorem 3]. ∎
6.4 Proofs Related to Essentially Equivalent RKHSs
In this subsection we prove Lemma 3.3 as well as Theorems 3.4 and 3.5.
To begin with, let us assume that $H_{1}$ and $H_{2}$ are RKHSs on $T$ with kernels $k_{1}$ and $k_{2}$. Then their sum $H_{1}+H_{2}$ can be equipped with the natural norm
$\|h\|_{H_{1}+H_{2}}^{2}=\inf\bigl{\{}\|h_{1}\|_{H_{1}}^{2}+\|h_{2}\|_{H_{2}}^{2}:h_{1}\in H_{1},h_{2}\in H_{2}\mbox{ with }h=h_{1}+h_{2}\bigr{\}}\,.$ (22)
It turns out that $H_{1}+H_{2}$ equipped with this norm is the RKHS of the kernel $k_{1}+k_{2}$, see e.g. [3, Theorem on page 353]. Consequently, (3) implies $H_{1}+H_{1}^{\star}\cong H_{2}+H_{2}^{\star}$ as noted at the beginning of Section 2. With these preparations, the proof of Lemma 3.3 is straightforward:
Proof of Lemma 3.3: By our assumption we have $k_{1}+k_{1}^{\star}=k_{2}+k_{2}^{\star}$. As discussed above this gives $H_{1}+H_{1}^{\star}=H_{2}+H_{2}^{\star}$, and hence also $H_{1}+H_{1}^{\star}\circeq H_{2}+H_{2}^{\star}$. ∎
In the following, we adopt the usual notation $H_{1}\oplus H_{2}:\circeq H_{1}+H_{2}$ if we additionally have $H_{1}\cap H_{2}=\{0\}$. Moreover, if we equip $H_{1}\oplus H_{2}$ with the norm given by (22), then the resulting RKHS is denoted by $H_{1}\oplus_{2}H_{2}$. In this case, taking the infimum in (22) is, of course, superfluous, and $H_{1}$ and $H_{2}$ are closed subspaces of $H_{1}\oplus_{2}H_{2}$ with $H_{1}^{\perp}=H_{2}$ and $H_{2}^{\perp}=H_{1}$.
Now, our first auxiliary result shows that sums of RKHSs are suitable direct sums if one of the involved spaces is finite dimensional.
Lemma 6.5.
Let $H$ and $H_{0}$ be RKHSs on $T$ with $\dim H_{0}<\infty$. Then there exists an RKHS $H^{\star}$ on $T$ with $H^{\star}\subset H_{0}$ and
$H\oplus H^{\star}\circeq H+H_{0}\,.$ (23)
Proof of Lemma 6.5: Since $H\cap H_{0}$ is a subspace of $H_{0}$ and $\dim H_{0}<\infty$, there exists a subspace $H^{\star}\subset H_{0}$ with
$H_{0}\circeq(H\cap H_{0})\oplus H^{\star}\,.$ (24)
Moreover, $\dim H_{0}<\infty$ shows that $H^{\star}$ is closed in $H_{0}$ and thus $H^{\star}$ is an RKHS.
We first show that $H\cap H^{\star}=\{0\}$. This this end, we pick an $h\in H\cap H^{\star}$. By $H^{\star}\subset H_{0}$ we find $h\in H\cap H_{0}$ and hence we have $h\in(H\cap H_{0})\cap H^{\star}$. By (24) we then obtain $h=0$.
For the verification of (23) we first note that $H^{\star}\subset H_{0}$ implies $H\oplus H^{\star}\subset H+H_{0}$. For the converse inclusion, we fix some $h\in H$ and $h_{0}\in H_{0}$. By (24) there then exist a $g\in H\cap H_{0}$ and an $h^{\star}\in H^{\star}$ with $h_{0}=g+h^{\star}$. This gives
$h+h_{0}=h+h_{0}-h^{\star}+h^{\star}\in H+H^{\star}\circeq H\oplus H^{\star}\,,$
where in the second step we used $h_{0}-h^{\star}=g\in H\cap H_{0}\subset H$. ∎
The next auxiliary lemma shows that we may use direct sums in the definition of essentially equivalent RKHSs and that we have some additional control over the spaces $H_{1}^{\star}$ and $H_{2}^{\star}$.
Lemma 6.6.
Let $H_{1}$ and $H_{2}$ be essentially equivalent RKHSs on $T$. If $E$ is a BSF on $T$ with $H_{1}\cup H_{2}\subset E$, then there exist finite dimensional RKHSs $H_{1}^{\star}$ and $H_{2}^{\star}$ on $T$ with $H_{1}^{\star}\cup H_{2}^{\star}\subset E$ and
$H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}\,.$
Proof of Lemma 6.6: By our assumption and Lemma 6.5 there exist finite dimensional RKHSs $H_{1}^{\times}$ and $H_{2}^{\times}$ on $T$ with
$H_{1}\oplus H_{1}^{\times}\circeq H_{2}\oplus H_{2}^{\times}\,.$ (25)
We equip both direct sums with the $\oplus_{2}$-norm and consider the two orthogonal projections $P_{i}:H_{i}\oplus_{2}H_{i}^{\times}\to H_{i}\oplus_{2}H_{i}^{\times}$ onto $H_{i}$ as well as the two orthogonal projections $Q_{i}:H_{i}\oplus_{2}H_{i}^{\times}\to H_{i}\oplus_{2}H_{i}^{\times}$ onto $H_{i}^{\times}$. Moreover, we define $H:\circeq H_{1}\oplus H_{1}^{\times}\circeq H_{2}\oplus H_{2}^{\times}$ and the subspace
$H^{\dagger}:\circeq\bigl{\{}(Q_{1}-Q_{2})h:h\in H\bigr{\}}\circeq\bigl{\{}(Q_{2}-Q_{1})h:h\in H\bigr{\}}\,.$
Clearly, $H^{\dagger}$ is a subspace of $H_{1}^{\times}+H_{2}^{\times}$ and thus a finite dimensional RKHS on $T$. Moreover, for $h\in H$ our construction yields
$P_{1}h+Q_{1}h=h=P_{2}h+Q_{2}h\,.$ (26)
In particular, for $h_{1}\in H_{1}$ we find $h_{1}=P_{1}h_{1}=P_{2}h_{1}+(Q_{2}-Q_{1})h_{1}\in H_{2}+H^{\dagger}$, and hence $H_{1}\subset H_{2}+H^{\dagger}$. The latter implies $H_{1}+H^{\dagger}\subset H_{2}+H^{\dagger}$. Since analogously we find $H_{2}+H^{\dagger}\subset H_{1}+H^{\dagger}$ by the symmetry of $H^{\dagger}$ in $Q_{1}$ and $Q_{2}$, we have
$H_{1}+H^{\dagger}\circeq H_{2}+H^{\dagger}\,.$
Moreover, (26) also shows $(Q_{1}-Q_{2})h=P_{2}h-P_{1}h\in H_{2}+H_{1}\subset E$, and hence we obtain $H^{\dagger}\subset E$. Applying Lemma 6.5 on both sides then yields the assertion. ∎
Lemma 6.7.
Let $H_{1},H_{2}$ and $H_{1}^{\star},H_{2}^{\star}$ be RKHSs on $T$ with $H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}$ and $m_{i}:=\dim H_{i}^{\star}<\infty$. For $i=1,2$ we write $\overline{H}_{i}:=H_{i}\oplus_{2}H_{i}^{\star}$ and consider the embedding maps $J_{i}:H_{i}\to\overline{H}_{i}$. Finally, let $E$ be a Banach space and $S_{i}:\overline{H}_{i}\to E$ be bounded and linear operators with $S_{1}h=S_{2}h$ for all $h\in\overline{H}_{1}\circeq\overline{H}_{2}$. Then there exists a constant $c\in(0,\infty)$ with
$d_{n+m_{2}}(S_{1}\circ J_{1})$ $\leq c\,d_{n}(S_{2}\circ J_{2})\,,$
$d_{n+m_{1}}(S_{2}\circ J_{2})$ $\leq c\,d_{n}(S_{1}\circ J_{1})$
for all $n\geq 1$.
Proof of Lemma 6.7: We first note that we have $\overline{H}_{1}\cong\overline{H}_{2}$ as mentioned at the beginning of Section 2, and hence there exists a constant $c\in(0,\infty)$ with
$c^{-1}\,d_{n}(S_{1})\leq d_{n}(S_{2})\leq c\,d_{n}(S_{1})\,,\qquad\qquad n\geq 1.$ (27)
Let us further write $P_{i}:\overline{H}_{i}\to\overline{H}_{i}$ for the orthogonal projection onto $H_{i}$ and $P^{\perp}_{i}:\overline{H}_{i}\to\overline{H}_{i}$ for the orthogonal projection onto its complement $H_{i}^{\perp}=H_{i}^{\star}$, that is $P^{\perp}_{i}=\operatorname{id}_{\overline{H}_{i}}-P_{i}$, where $\operatorname{id}_{\overline{H}_{i}}$ is the identity map on $\overline{H}_{i}$. Moreover, we define $Q_{i}:\overline{H}_{i}\to H_{i}$ by $Q_{i}h:=P_{i}h$ for all $h\in\overline{H}_{i}$. In other words, the only difference between $P_{i}$ and $Q_{i}$ are the spaces the operators formally map into. Now, by construction the operators $Q_{i}$ are metric surjections and we have
$P_{i}=J_{i}\circ Q_{i}\qquad\qquad\mbox{ and }\qquad\qquad Q_{i}\circ J_{i}=\operatorname{id}_{H_{i}}\,.$
By the additivity of the Kolmogorov numbers we then find for all $n\geq 1$
$d_{n+m_{i}}(S_{i})\leq d_{m_{i}+1}(S_{i}-S_{i}\circ P_{i})+d_{n}(S_{i}\circ P_{i})$ $=d_{m_{i}+1}(S_{i}\circ P^{\perp}_{i})+d_{n}(S_{i}\circ J_{i}\circ Q_{i})$
$\leq\|S_{i}\|\cdot d_{m_{i}+1}(P^{\perp}_{i})+d_{n}(S_{i}\circ J_{i}\circ Q_{i})$
$=d_{n}(S_{i}\circ J_{i})\,,$
where in the last step we used $\operatorname{rank}P_{i}^{\perp}=m_{i}<m_{i}+1$ and the surjectivity of the Kolmogorov numbers. Combining this inequality with $\|J_{i}\|=1$ and (27) yields
$d_{n+m_{2}}(S_{1}\circ J_{1})\leq d_{n+m_{2}}(S_{1})\leq c\,d_{n+m_{2}}(S_{2})\leq c\,d_{n}(S_{2}\circ J_{2})\,,$
and the second inequality can be shown analogously. ∎
Proof of Theorem 3.4: Since $H_{1}$ and $H_{2}$ be essentially equivalent with $H_{1},H_{2}\subset{{\cal L}_{\infty}({\cal B})}$, Lemma 6.6 shows that there are finite dimensional RKHSs $H_{1}^{\star}\subset{{\cal L}_{\infty}({\cal B})}$ and $H_{2}^{\star}\subset{{\cal L}_{\infty}({\cal B})}$ on $T$ with
$H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}\,.$
We write $m_{i}:=\dim H_{i}^{\star}$ and observe that $\overline{H}_{i}:=H_{i}\oplus_{2}H_{i}^{\star}\subset{{\cal L}_{\infty}({\cal B})}$ for $i=1,2$. Consequently, we can consider the operators $S_{i}:=I_{\overline{H}_{i},\nu}:\overline{H}_{i}\to{L_{2}(\nu)}$. Clearly this shows $S_{1}h=S_{2}h$ for all $h\in\overline{H}_{1}\circeq\overline{H}_{2}$ and for the embedding maps $J_{i}:H_{i}\to\overline{H}_{i}$ we have
$S_{i}\circ J_{i}=I_{H_{i},\nu}\,.$
Lemma 6.7 thus gives a constant $c\in(0,\infty)$ with
$d_{n+m_{2}}(I_{H_{1},\nu})\leq c\,d_{n}(I_{H_{2},\nu})\qquad\qquad\mbox{ and }\qquad\qquad d_{n+m_{1}}(I_{H_{2},\nu})\leq c\,d_{n}(I_{H_{1},\nu})$ (28)
for all $n\geq 1$. Let us now assume that $I_{H_{1},\nu}$ is nuclear. Then $I_{H_{1},\nu}$ is compact and (18) together with the discussion around it shows both $s_{n}(I_{H_{1},\nu})=d_{n}(I_{H_{1},\nu})$ and $d_{n}(I_{H_{1},\nu})\to 0$. By (28) we conclude that $d_{n}(I_{H_{2},\nu})\to 0$ and hence $I_{H_{2},\nu}$ is also compact as discussed around (18). Consequently, we have $s_{n}(I_{H_{2},\nu})=d_{n}(I_{H_{2},\nu})$, and therefore the second inequality in (28) shows
$s_{n+m_{1}}(I_{H_{2},\nu})\leq c\,s_{n}(I_{H_{1},\nu})$
for all $n\geq 1$. Since we assumed that $I_{H_{1},\nu}$ is nuclear, we then see that $I_{H_{2},\nu}$ is also nuclear. The converse implication can be shown analogously. ∎
Proposition 6.8.
Let $H_{1},H_{2}$ be RKHS on $T$ and $H_{1}^{\star},H_{2}^{\star}$ be finite dimensional RKHSs on $T$ with $H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}$. Moreover, let $H$ be an RKHS on $T$ with $H_{1}\ll H$. Then there exists an RKHS $H_{1}^{\dagger}\subset H_{1}^{\star}$ with
$H_{2}\ll H\oplus_{2}H_{1}^{\dagger}\,.$
Proof of Proposition 6.8: For $i=1,2$ we consider the embedding maps $J_{i}:H_{i}\to H_{i}\oplus_{2}H_{i}^{\star}$. Now, we first note that by Lemma 6.5 there exists an $H_{1}^{\dagger}\subset H_{1}^{\star}$ with
$H\oplus H_{1}^{\dagger}\circeq H+H_{1}^{\star}\,.$
Moreover, by our assumption $H_{1}\ll H$ we have $H_{1}\subset H$ and the resulting embedding $J:H_{1}\to H$ is Hilbert-Schmidt. Since $H_{1}\subset H$ gives
$H_{2}\oplus H_{2}^{\star}\circeq H_{1}\oplus H_{1}^{\star}\subset H+H_{1}^{\star}\circeq H\oplus H_{1}^{\dagger}\,,$
we can thus consider the two resulting embedding maps $S_{i}:H_{i}\oplus_{2}H_{i}^{\star}\to H\oplus_{2}H_{1}^{\dagger}$. Clearly, this gives $S_{i}h=S_{2}h$ for all $h\in H_{1}\oplus_{2}H_{1}^{\star}\circeq H_{2}\oplus_{2}H_{2}^{\star}$. Moreover, the operators $S_{i}$ are continuous since embedding maps between RKHSs are automatically continuous by a simple application of the closed graph theorem, see e.g. [34, Lemma 23]. By Lemma 6.7 there then exists a constant $c\in(0,\infty)$ such that for $m_{1}:=\dim H_{1}^{\dagger}$ we have
$d_{n+m_{1}}(S_{2}\circ J_{2})$ $\leq c\,d_{n}(S_{1}\circ J_{1})\,,\qquad\qquad n\geq 1.$
Let us now consider the embedding map $I:H\to H\oplus_{2}H_{1}^{\dagger}$, which by construction satisfies $\|I\|\leq 1$. Clearly, our construction gives
and hence we obtain
$d_{n+m_{1}}(S_{2}\circ J_{2})\leq c\,d_{n}(S_{1}\circ J_{1})=c\,d_{n}(I\circ J)\leq c\,d_{n}(J)$
for all $n\geq 1$. Now, $J$ is Hilbert-Schmidt, and hence compact. This shows both $d_{n}(J)\to 0$ and $d_{n}(J)=s_{n}(J)$ as discussed around (18). Consequently, we find $d_{n}(S_{2}\circ J_{2})\to 0$ and thus also $s_{n}(S_{2}\circ J_{2})=d_{n}(S_{2}\circ J_{2})$, where for the second assertion we again used the properties mentioned around (18). In summary, we thus have $s_{n+m_{1}}(S_{2}\circ J_{2})\leq c\,s_{n}(J)$, and since $J$ is Hilbert-Schmidt, so ist $S_{2}\circ J_{2}$. However, $S_{2}\circ J_{2}:H_{2}\to H\oplus_{2}H_{1}^{\dagger}$ is the embedding map, and hence we have found $H_{2}\ll H\oplus_{2}H_{1}^{\dagger}$. ∎
Proof of Theorem 3.5: Let us consider the RKHS $E:=H^{\dagger}+H_{X}$. By our assumptions we have $H_{X}\cup H\subset H_{X}\cup H^{\dagger}\subset E$, and therefore Lemma 6.6 provides us with finite dimensional RKHSs $H^{\star}$ and $H_{X}^{\star}$ on $T$ with $H^{\star}\cup H_{X}^{\star}\subset E$ and
$H\oplus H^{\star}\circeq H_{X}\oplus H_{X}^{\star}\,.$
Combining this with the assumed $H\ll H^{\dagger}$, Proposition 6.8 gives us an RKHS $H^{\ddagger}\subset H^{\star}$ with
$H_{X}\ll H^{\dagger}\oplus_{2}H^{\ddagger}\,.$
In addition, note that $H_{X}^{\dagger}:=H^{\dagger}\oplus_{2}H^{\ddagger}$ is essentially equivalent to $H^{\dagger}$ and our construction ensures $H_{X}^{\dagger}=H^{\dagger}\oplus_{2}H^{\ddagger}\subset H^{\dagger}+H^{\star}\subset H^{\dagger}+H^{\dagger}+H_{X}=H^{\dagger}+H_{X}$, where for the second inclusion we used $H^{*}\subset E=H^{\dagger}+H_{X}$. ∎
6.5 Proof of Theorem 3.6
Proof of Theorem 3.6: i). We first recall that $I_{X,\nu}$ is compact by our assumption $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$ and Lemma 6.4. The definition of singular numbers then yields
$s_{i}(I_{X,\nu})=\sqrt{\mu_{i}(I_{X,\nu}\circ I_{X,\nu}^{*})}=\sqrt{\mu_{i}(T_{X,\nu})}\,.$
By (14) we thus obtain the equivalence.
ii). Let $J:H_{X}\to H_{X}^{\beta}$ be the embedding map. For $i\in I$ we then have
$\|J(\sqrt{\mu_{i}}e_{i})\|_{H_{X}^{\beta}}=\mu_{i}^{1/2-\beta/2}\|\mu_{i}^{\beta/2}e_{i}\|_{H_{X}^{\beta}}=\mu_{i}^{1/2-\beta/2}\,,$
and this in turn shows
$\sum_{i\in I}\|J(\sqrt{\mu_{i}}e_{i})\|_{H_{X}^{\beta}}^{2}=\sum_{i\in I}\mu_{i}^{1-\beta}\,.$
Since the $H_{X}$-positivity of $\nu$ ensures that $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ is an ONB of $H_{X}$, see [39, Theorem 3.1], we can then conclude that (8) indeed characterizes the Hilbert-Schmidt property of $J$, see the discussion around (15). The second assertion directly follows from $H_{X}\ll H_{X}^{\beta}$ and Theorem 1.2. ∎
6.6 Proofs Related to RKHSs of Sobolev Type
Before we can prove Theorem 4.1 some preparations are required. We begin by providing details for (9). To this end, let $B^{s}_{p,q}(\mathbb{R}^{d})$ and $F^{s}_{p,q}(\mathbb{R}^{d})$ denote the Besov spaces, respectively Triebel-Lizorkin spaces, see e.g. [1, 10]. Then for $s>d/2$ we have
$H^{s}(\mathbb{R}^{d})\cong F_{2,2}^{s}(\mathbb{R}^{d})\cong B_{2,2}^{s}(\mathbb{R}^{d})\hookrightarrow C_{b}(\mathbb{R}^{d})\,,$ (29)
where the first identity follows from e.g. [1, Chapter 7, Paragraphs 7.62, 7.65, and 7.57], the second identity is shown in e.g. [1, Chapter 7, Paragraph 7.67], and the inclusion can be found in e.g. [1, Theorem 7.34].
In addition, Triebel-Lizorkin and Besov spaces on open Euclidean balls $T\subset\mathbb{R}^{d}$ are defined analogously to (10), that is $A^{s}_{p,q}(T):=\{f_{|T}:f\in A^{s}_{p,q}(\mathbb{R}^{d})\}$ with norm
$\|h\|_{A^{s}_{p,q}(T)}:=\inf\bigl{\{}\|f\|_{A^{s}_{p,q}(\mathbb{R}^{d})}:f\in A^{s}_{p,q}(\mathbb{R}^{d})\mbox{ with }f_{|T}=h\bigr{\}}\,$
where $A^{s}_{p,q}$ stands for either $F^{s}_{p,q}$ or $B^{s}_{p,q}$, see e.g. [10, Chapter 2.5]. Note that if $F_{p,q}^{s}(\mathbb{R}^{d})$ or $B_{p,q}^{s}(\mathbb{R}^{d})$ are BSFs then the same is true for $F_{p,q}^{s}(T)$ or $B_{p,q}^{s}(T)$. In addition, continuous inclusions between such spaces on $\mathbb{R}^{d}$ carry over to the restricted spaces on $T$, see e.g. [10, Chapter 2.5] or [34, Proposition 3] for a general BSF statement. For $s>d/2$, Equation (29) thus implies
$H^{s}(T)\cong F_{2,2}^{s}(T)\cong B_{2,2}^{s}(T)\hookrightarrow C_{b}(T)\,.$ (30)
Finally, for the proof of Theorem 4.1 we need the following auxiliary result that investigates both the operators $I_{H^{s}(T),\lambda^{d}}:H^{s}(T)\to{L_{2}(T)}$ and the nuclear dominance between fractional Sobolev spaces.
Lemma 6.9.
Let $T\subset\mathbb{R}^{d}$ be bounded and measurable with non-empty interior and $s>d/2$. Then the following statements hold true:
  1. i)
    The operator $I_{H^{s}(T),\lambda^{d}}:H^{s}(T)\to{L_{2}(T)}$ is nuclear if and only if $s>d$.
  2. ii)
    If $s>d$ and $r>d/2$, then $H^{s}(T)\ll H^{r}(T)$ if and only if $r<s-d/2$.
Proof of Lemma 6.9: i). We first consider the case, when $T$ is an open Euclidean ball. Then we have $B_{2,2}^{0}(\mathbb{R}^{d})\cong F_{2,2}^{0}(\mathbb{R}^{d})\cong{L_{2}(\mathbb{R}^{d})}$, where the first identity follows from e.g. [10, (1) in Chapter 2.3.3] and the second identity can be found in e.g. [42, (5) in Chapter 2.5.6]. Restricting to $T$ leads to $B_{2,2}^{0}(T)\cong{L_{2}(T)}$, and therefore [10, (10) in Chapter 2.5.1] shows that the embedding $B_{2,2}^{s}(T)\hookrightarrow{L_{2}(T)}$ is compact. By (30) we can thus conclude that $I_{H^{s}(T),\lambda^{d}}:H^{s}(T)\to{L_{2}(T)}$ is also compact, and by (18) we hence find
$a_{n}(I_{H^{s}(T),\lambda^{d}})=d_{n}(I_{H^{s}(T),\lambda^{d}})=s_{n}(I_{H^{s}(T),\lambda^{d}})\,,\qquad\qquad n\geq 1.$ (31)
Moreover, the already observed $H^{s}(T)\cong B_{2,2}^{s}(T)$ and $B_{2,2}^{0}(T)\cong{L_{2}(T)}$ in combination with the results in [10, Chapter 3.3.4] show that
$a_{n}(I_{H^{s}(T),\lambda^{d}})\sim n^{-s/d}\,.$ (32)
Combining (31) with (32) then yields the characterization.
To prepare the proof of the general case, let us assume that we have measurable $T_{1}\subset T_{2}\subset\mathbb{R}^{d}$ with $\lambda^{d}(T_{1}),\lambda^{d}(T_{2})\in(0,\infty)$. Moreover, let $Q:H^{s}(T_{2})\to H^{s}(T_{1})$ and $R:{L_{2}(T_{2})}\to{L_{2}(T_{1})}$ be the restriction operators. Then $Q$ and $R$ are metric surjections and we have the commutative diagram
[…]
By the surjectivity of the Kolmogorov numbers we thus find
$d_{n}(I_{H^{s}(T_{1}),\lambda^{d}})=d_{n}(I_{H^{s}(T_{1}),\lambda^{d}}\circ Q)=d_{n}(R\circ I_{H^{s}(T_{2}),\lambda^{d}})\leq d_{n}(I_{H^{s}(T_{2}),\lambda^{d}})\,,\qquad\qquad n\geq 1.$ (33)
With these preparations we can now establish the assertion in the general case. To this end, we first note that by our assumptions on $T$ there exist open balls $B_{0}\subset T\subset B_{1}$. Now, we have already seen at the very beginning that $I_{H^{s}(B_{i}),\lambda^{d}}$ is compact, and hence we have $d_{n}(I_{H^{s}(B_{i}),\lambda^{d}})\to 0$ for $i=0,1$. Applying (33) we further find $d_{n}(I_{H^{s}(T),\lambda^{d}})\to 0$, and therefore $I_{H^{s}(T),\lambda^{d}}$ is also compact. By (18) we conclude that
$s_{n}(I_{H^{s}(T),\lambda^{d}})=d_{n}(I_{H^{s}(T),\lambda^{d}})\,,\qquad\qquad n\geq 1,$
while our initial considerations (31) with (32) showed
$d_{n}(I_{H^{s}(B_{i}),\lambda^{d}})=a_{n}(I_{H^{s}(B_{i}),\lambda^{d}})\sim n^{-s/d}\,,\qquad\qquad n\geq 1$
for $i=0,1$. Combining these identities with (33) yields $s_{n}(I_{H^{s}(T),\lambda^{d}})\sim n^{-s/d}$.
ii). Again, we first consider the case, when $T$ is an open ball. Here, we note that $H^{s}(T)\cong B_{2,2}^{s}(T)$ and $H^{r}(T)\cong B_{2,2}^{r}(T)$ in combination with the results in [10, Chapter 3.3.4] show that
$a_{n}(J)\sim n^{-(s-r)/d}\,,$
where $J:H^{s}(T)\to H^{r}(T)$ denotes the embedding map. By the results mentioned around (18) the same is true for the Kolmogorov numbers and in particular we have $d_{n}(J)\to 0$. Therefore, $J$ is compact and (18) shows that
$s_{n}(J)\sim n^{-(s-r)/d}\,.$
Consequently, we have $\sum_{n\geq 1}s_{n}^{2}(J)<\infty$ if and only if $r<s-d/2$.
To prepare the proof of the general case, we again assume that we have measurable $T_{1}\subset T_{2}\subset\mathbb{R}^{d}$ with $\lambda^{d}(T_{1}),\lambda^{d}(T_{2})\in(0,\infty)$. Moreover, let $Q_{s}:H^{s}(T_{2})\to H^{s}(T_{1})$ and $Q_{r}:H^{r}(T_{2})\to H^{r}(T_{1})$ be the restriction operators, which by construction of the involved spaces are metric surjections. In addition, note that $H^{s}(\mathbb{R}^{d})\subset H^{r}(\mathbb{R}^{d})$ ensures that the embeddings $J_{i}:H^{s}(T_{i})\to H^{r}(T_{i})$ are defined for $i=1,2$. Clearly, these maps enjoy the following commutative diagram
[…]
By the surjectivity of the Kolmogorov numbers we thus find
$d_{n}(J_{1})=d_{n}(J_{1}\circ Q_{s})=d_{n}(Q_{r}\circ J_{2})\leq d_{n}(J_{2})\,,\qquad\qquad n\geq 1.$
With these preparations, the rest of the proof follows the arguments used in the proof of the general case of i). ∎
Proof of Theorem 4.1: ii) $\Rightarrow$ i). By Theorem 3.1 we know that $I_{X,\lambda^{d}}:H_{X}\to{L_{2}(T)}$ is nuclear. Theorem 3.4 then shows that $I_{H^{s}(T),\lambda^{d}}:H^{s}(T)\to{L_{2}(T)}$ is also nuclear. Now the assertion follows by Lemma 6.9.
i) $\Rightarrow$ ii). Let us fix an $d/2<r<s-d/2$. Lemma 6.9 then shows $H^{s}(T)\ll H^{r}(T)$. Since $H_{X}$ is essentially equivalent to $H^{s}(T)$ by assumption, Theorem 3.5 then provides an RKHS $H_{X}^{\dagger}$ on $T$ that is essentially equivalent to $H^{r}(T)$ and satisfies $H_{X}\ll H_{X}^{\dagger}$. Now the assertion follows from Theorem 1.2 for $H:=H_{X}^{\dagger}$. ∎
Proof of Theorem 4.2: Using (30) we first note that it suffices to consider the case $H_{0}=B^{s}_{2,2}(0,1)$. In the following we thus assume that $\hat{k}_{0}$ is the kernel of $\hat{H}_{0}:=B_{2,2}^{s}(\mathbb{R})$ and that $k_{0}$ is its restriction to $(0,1)$.
Now let $S_{2,2}^{s}B(\mathbb{R}^{d})$ denote the Besov space of dominating mixed smoothness $s$, see [35, Appendix A.2] and the references mentioned therein. Then [35, Theorem 2.2] shows
$S_{2,2}^{s}B(\mathbb{R}^{d})\cong B^{s}_{2,2}(\mathbb{R})\otimes_{\alpha_{2}}\dots\otimes_{\alpha_{2}}B^{s}_{2,2}(\mathbb{R})\,,$ (34)
where $\alpha_{2}$ denotes the $2$-nuclear tensor norm, see [35, Definition B.2]. Moreover, the $\alpha_{2}$-tensor norm equals the $g_{2}:=\alpha_{2,1}$-tensor norm in [7, Chapters 12.5 and 12.7], and $g_{2}$ in turn equals the natural Hilbert space tensor norm, see [7, Chapters 26.6 and 26.7], where the latter is called Hilbert-Schmidt norm. Denoting the natural Hilbert space tensor product by $\otimes_{2}$, the Identity (34) thus reads as
$S_{2,2}^{s}B(\mathbb{R}^{d})\cong B^{s}_{2,2}(\mathbb{R})\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(\mathbb{R})\,.$ (35)
Now the tensor product space on the right hand side can be interpreted as the RKHS of the product kernel
$\prod_{i=1}^{d}\hat{k}_{0}(t_{i},t_{i}^{\prime})\,,\qquad\qquad t,t^{\prime}\in\mathbb{R}^{d},$
see e.g. [4, Chapter 1.4.6] or [38, Lemma 4.6]. Restricting both sides of (35) to $T$ thus gives
$S_{2,2}^{s}B(T)\cong B^{s}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(0,1)=H_{X}\,,$ (36)
where the space $S_{2,2}^{s}B(T)$ is defined by restricting the functions of $S_{2,2}^{s}B(\mathbb{R}^{d})$ to $T$ as in [24, Definition 3.8 and Remark 3.9]. Moreover, the embedding $I:S_{2,2}^{s}B(T)\to{L_{2}(T)}$ is defined and compact and its approximation numbers satisfy
$a_{n}(I)\sim n^{-s}(\ln n)^{(d-1)s}$ (37)
see [33] as well as [24, Theorem 2.11] for a presentation of the results of [33] in the notation we use in this proof.
ii) $\Rightarrow$ i). By Theorem 3.1 we know that $I_{X,\lambda^{d}}:H_{X}\to{L_{2}(T)}$ is nuclear. By combining (36), (37), and (18) we thus find $s>1$.
i) $\Rightarrow$ ii). We pick an $r\in(1/2,s-1/2)$ and an ONB $(e_{i})$ of $B^{s}_{2,2}(0,1)$. By Lemma 6.9 and (30) we then know that $B^{s}_{2,2}(0,1)\ll B^{r}_{2,2}(0,1)$, and hence we have
$\sum_{i=1}^{\infty}\|e_{i}\|_{B^{r}_{2,2}(0,1)}^{2}<\infty$
as discussed around (15). Now recall that the collection of elementary tensors $e_{i_{1}}\otimes\dots\otimes e_{i_{d}}$ forms an ONB of $B^{s}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(0,1)=H_{X}$. Moreover, we have
$\sum_{i_{1}=1}^{\infty}\dots\sum_{i_{d}=1}^{\infty}\|e_{i_{1}}\otimes\dots\otimes e_{i_{d}}\|_{B^{r}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{r}_{2,2}(0,1)}^{2}$ $=\sum_{i_{1}=1}^{\infty}\dots\sum_{i_{d}=1}^{\infty}\prod_{j=1}^{d}\|e_{i_{j}}\|_{B^{r}_{2,2}(0,1)}^{2}$
$=\biggl{(}\sum_{i=1}^{\infty}\|e_{i}\|_{B^{r}_{2,2}(0,1)}^{2}\biggr{)}^{d}$
[…]
which shows that the embedding $B^{s}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(0,1)\hookrightarrow B^{r}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{r}_{2,2}(0,1)$ is Hilbert-Schmidt. Now ii) follows by Theorem 1.2. ∎
6.7 Proofs for the Examples
Proof of Example 5.1: Let $\nu:=\lambda_{|[0,1]}$ be the Lebesgue measure on $T$. Then the eigenvalues of $T_{W,\nu}$ are given by
$\mu_{i}(T_{W,\nu})=\Bigl{(}\frac{2}{(2i-1)\pi}\Bigr{)}^{2}\,,\qquad\qquad i\in\mathbb{N},$ (38)
see e.g. [2, Chapter III.4, page 77]. Consequently, (7) is false, and hence $I_{W,\nu}:H_{W}\to{L_{2}(\nu)}$ is not nuclear by Theorem 3.6. Theorem 3.1 then yields the second assertion. ∎
Proof of Example 5.2: Let us recall that the covariance function of the Brownian bridge $X:=(X_{t})_{t\in T}$ is
$k_{X}(t_{1},t_{2})=\min\{t_{1},t_{2}\}-t_{1}t_{2}=k_{W}(t_{1},t_{2})-t_{1}t_{2}\,,\qquad\qquad t_{1},t_{2}\in[0,1],$
where $k_{W}$ is the covariance function of the Wiener process. Now, $(t_{1},t_{2})\mapsto t_{1}t_{2}$ is a kernel on $[0,1]$ whose RKHS is one dimensional, and consequently, $H_{X}$ and $H_{W}$ are essentially equivalent by Lemma 3.3. Moreover, we have seen in Example 5.1 that $I_{W,\lambda}:H_{W}\to{L_{2}(T)}$ is not nuclear. Theorem 3.4 thus shows that $I_{X,\lambda}:H_{X}\to{L_{2}(T)}$ is not nuclear, either. Theorem 3.1 then yields the assertion. ∎
Proof of Example 5.3: We first consider the covariance function $k_{X}^{(1)}$. For the corresponding RKHS $H_{X}^{(1)}$ we then find $H_{X}^{(1)}\cong H^{1}(T)$ by e.g. [4, Example 17 on p. 316] applied for some $a,b\in\mathbb{R}$ with $T\subset(a,b)$ followed by a restriction onto $T$. By Theorem 4.1 we then obtain the assertion.
In the case of the second covariance, we first observe that $(t_{1},t_{2})\mapsto a\exp(-\sigma(t_{1}+t_{2}))$ gives a kernel $k^{*}$ on $[0,1]$ whose RKHS is one dimensional. Moreover, we obviously have $k_{X}^{(2)}=k_{X}^{(1)}-k^{*}$, and therefore the RKHS $H_{X}^{(2)}$ of $k_{X}^{(2)}$ is essentially equivalent to $H_{X}^{(1)}\cong H^{1}(T)$ by Lemma 3.3. Now we again obtain the assertion by Theorem 4.1. ∎
Proof of Example 5.4: In the following we write $H_{\alpha}:=H_{B^{(\alpha)}}$ for the RKHS of $B^{(\alpha)}$. Then [27, Theorem 6.12] shows that a function $h:T\to\mathbb{R}$ is in $H_{\alpha}$ if and only if there are sequences $(a_{n})_{n\geq 0}$ and $(b_{n})_{n\geq 1}$ with
$\sum_{n\geq 0}n^{2\alpha+1}a_{n}^{2}<\infty\qquad\mbox{ and }\qquad\sum_{n\geq 1}n^{2\alpha+1}b_{n}^{2}<\infty$
such that for all $t\in[0,1]$ we have
$h(t)=h(1)t+\sum_{n\geq 0}a_{n}\cos(2\pi nt)+\sum_{n\geq 1}a_{n}\sin(2\pi nt)\,.$
With these preparations we can now turn to the actual proof.
Here, we first note that in the case $\alpha=1/2$ we have $B^{(\alpha)}=W$, and hence there is nothing left to prove by Example 5.1.
Moreover, in the case $20<\alpha<1/$ 2 the above description of $H_{\alpha}$ yields $H_{1/2}\subset H_{\alpha}$ and the corresponding embedding map $J:H_{1/2}\to H_{\alpha}$ is continuous by a simple application of the closed graph theorem, see [34, Lemma 23]. Now assume that there was an RKHS $H\subset{{\cal L}_{\infty}(T)}$ with $\mathrm{P}(\{B_{\bullet}^{(\alpha)}\in H\})=1$. By Theorem 3.1 we would then know that $I_{H_{\alpha},\lambda}:H_{\alpha}\to{L_{2}(T)}$ was nuclear. The natural factorization
then shows that $I_{H_{1/2},\lambda}=I_{W,\lambda}$ would also be nuclear, which by Example 5.1 is not the case.
Finally, in the case $11/2<\alpha<$ 1 it is well known that $\mathrm{P}$-almost all paths of $B^{(\alpha)}$ are $\beta$-Hölder continuous for all $\beta\in(0,\alpha)$, where we note that this can be quickly verified as the stationary increments and the self-similarity of $B^{(\alpha)}$ gives
$\mathbb{E}\bigl{|}B^{(\alpha)}_{t_{1}}-B^{(\alpha)}_{t_{2}}\big{|}^{p}=\mathbb{E}\bigl{|}B^{(\alpha)}_{|t_{1}-t_{2}|}\big{|}^{p}=|t_{1}-t_{2}|^{\alpha p}\cdot\mathbb{E}\bigl{|}B^{(\alpha)}_{1}\big{|}^{p}\,,\qquad\qquad t_{1},t_{2}\geq 0,p>0\,,$
and therefore an application of the Kolmogorov-Chentsov theorem, see e.g. [19, Theorem 21.6] for $p>0$ with $\beta<(\alpha p-1)/p$ gives the $\beta$-Hölder continuity. Let us now fix some $s\in(1/2,\alpha)$ and some $\beta\in(s,\alpha)$. By e.g. [42, (6) and (9) in Chapter 2.5.7 in combination with (1) and (3) in Chapter 2.2.2] we then have
$\operatorname{H\ddot{o}l}^{\beta}(\mathbb{R})\cong B^{\beta}_{\infty,\infty}(\mathbb{R})\,,$
while for open and bounded intervals $I$ we have
$B^{\beta}_{\infty,\infty}(I)\hookrightarrow B^{s}_{2,2}(I)\,,$
see e.g. [42, (7) in Chapter 3.3.1]. Taking such an $I$ with $T\subset I$, the combination of both results restricted to $T$ then leads to
$\operatorname{H\ddot{o}l}^{\beta}(T)\cong B^{\beta}_{\infty,\infty}(T)\hookrightarrow B^{s}_{2,2}(T)\cong H^{s}(T)\,,$
where the last identity was noted in (30). Since we have already seen that almost all paths are $\beta$-Hölder continuous, we thus obtain the assertion. ∎
Proof of Example 5.5: We first consider the case $\alpha\leq 1/2$. Here we recall from [27, Theorem 5.4] that the RKHS $H_{R^{(\alpha)}}$ of the Riemann-Liouville process and the RKHS $H_{B^{(\alpha)}}$ of the fractional Brownian motion are equal up to equivalent norms. Consequently, they are essentially equivalent and therefore the assertion follows by combining Example 5.4 with Theorem 3.4 and Theorem 3.1.
Let us now consider the case $\alpha\in(1/2,1)$. Here, we need to recall that we can jointly realize the processes $R^{(\alpha)}$ and $B^{(\alpha)}$ such that the paths of $R^{(\alpha)}-B^{(\alpha)}$ are $C^{\infty}$, see e.g. [27, Theorem 5.1]. Since for $s\in(1/2,\alpha)$ all $C^{\infty}$-functions are contained in $B_{2,2}^{s}(T)$, see e.g. [1, Paragraph 7.32], the identity $B_{2,2}^{s}(T)\cong H^{s}(T)$, see (30), in combination with
$\mathrm{P}(\{B_{\bullet}^{(\alpha)}\in H^{s}(T)\})=1\,,$
which has been shown in Example 5.4, gives $\mathrm{P}(\{R_{\bullet}^{(\alpha)}\in H^{s}(T)\})=1$. ∎
Proof of Example 5.6: Since $T$ is bounded, there exists an open Euclidean ball $B$ with $T\subset B$. Moreover, the RKHS $H_{\alpha,\sigma}(B)$ of the Matérn kernel on $B$ satisfies $H_{\alpha,\sigma}(B)\cong B_{2,2}^{\alpha+d/2}(B)$, see e.g. [43, Corollary 10.13] in combination with e.g. [32, (4.15)], or alternatively [31, (27)]. In addition, we have $B_{2,2}^{\alpha+d/2}(B)\cong H^{\alpha+d/2}(B)$, see (30), and hence we find
$H_{\alpha,\sigma}(B)\cong H^{\alpha+d/2}(B)\,.$
Restricting both spaces onto $T$ leads to $H_{\alpha,\sigma}(T)\cong H^{\alpha+d/2}(T)$ by [34, Proposition 3]. By Theorem 4.1 and $H_{X}=H_{\alpha,\sigma}(T)$ we then obtain the assertion. ∎
Proof of Example 5.7: We first note that we have $k_{X}(t,t)=k_{X}(t-t,0)=k_{X}(0,0)$ for all $t\in T$. Consequently, $k_{X}$ is bounded, see e.g. [38, (4.14)]. In addition, for $\kappa(r):=k_{X}(r,0)$ we have $k_{X}(t_{1},t_{2})=k_{X}(t_{2},t_{1})=k_{X}(-t_{1}+t_{2},0)=\kappa(-t_{1}+t_{2})$, and hence $k_{X}$ is translation invariant. Finally, recall that $k_{X}$ is measurable by assumption. With the help of these properties of $k_{X}$ [40, Lemma 4.1 and 4.2] then show that $k_{X}$ is continuous and there exist an at most countable family $(e_{i}^{*})_{i\in I}$ of continuous functions $T\to\mathbb{R}$ with $\sup_{i\in I}\|e_{i}^{*}\|_{\infty}\leq\sqrt{2}$ and a summable family $(\mu_{i})_{i\in I}\subset[0,\infty)$ such that $([e_{i}^{*}]_{\sim})_{i\in I}$ is an ONB of ${L_{2}(\lambda)}$ and
$k_{X}(t_{1},t_{2})=\sum_{\mu_{i}>0}\mu_{i}e_{i}^{*}(t_{1})e_{i}^{*}(t_{2})\,,\qquad\qquad t_{1},t_{2}\in T,$ (39)
where the series converges absolutely for all $t_{1},t_{2}\in T$ as well as uniformly in $(t_{1},t_{2})$. Moreover, [40, Lemma 4.2] shows that the positive $\mu_{i}$ are the collection of non-zero eigenvalues of $T_{k,\lambda}$ including geometric multiplicities and the $[e_{i}^{*}]_{\sim}$ are corresponding eigenfunctions. Finally, for the functions defined in (4) with $f_{i}:={e_{i}^{*}}$ and $\mu_{i}>0$ we have
$e_{i}=\mu_{i}^{-1}I_{X,\nu}^{*}[e_{i}^{*}]_{\sim}=\mu_{i}^{-1}\sum_{\mu_{j}>0}\mu_{j}\bigl{\langle}[e_{i}^{*}]_{\sim},[e_{j}^{*}]_{\sim}\bigr{\rangle}_{{L_{2}(\lambda)}}e_{j}^{*}=e_{i}^{*}\,,$
where in the second step we used [39, Theorem 2.11] in combination with (39).
In addition, since $k_{X}$ is continuous, every $h\in H_{X}$ is continuous. Moreover, the Haar measure satisfies $\lambda(O)>0$ for all non-empty and open $O\subset G$. Together, these properties show that $\lambda$ is $H_{X}$-positive.
Using all these properties, the characterization now follows by combining Theorem 3.1 and Theorem 3.6 as discussed after Theorem 3.6. In addition, $\sup_{i\in I}\|e_{i}^{*}\|_{\infty}\leq\sqrt{2}$ in combination with $e_{i}^{*}\in C(T)$ and the summability assumption on the eigenvalues ensures that (6) converges absolutely in $C(T\times T)$, and thus $k_{X}^{1/2}$ is continuous. This in turn implies $H_{X}^{1/2}\subset C(T)$. ∎