### v3

# When does a Gaussian process have its paths in a reproducing kernel Hilbert space?

Abstract
We investigate for which Gaussian processes there do or do not exist reproducing kernel Hilbert spaces (RKHSs) that contain almost all of their paths. In particular, we establish a new result that makes it possible to exclude the existence of such RKHSs in many cases. Moreover, we combine this negative result with some known techniques to establish positive results. Here it turns out that for many classical families of Gaussian processes we can fully characterize for which members of these families there exist RKHSs containing the paths. Similar characterizations are obtained for Gaussian processes, for which the RKHSs of their covariance functions are Sobolev spaces or Sobolev spaces of mixed smoothness.
Mathematical Subject Classification (2010). Primary 60G15, 60G17; Secondary 46E22, 46N30.
1 Introduction
Given a centered Gaussian process $X:=(X_{t})_{t\in T}$ on some (complete) probability space $(\Omega,{\cal A},\mathrm{P})$, a classical task is to describe its path properties by function spaces such as spaces of continuous functions. One particular instance of this general task asks, whether the paths $X_{\bullet}$ of the process are almost surely contained in a suitable reproducing kernel Hilbert space (RKHS) $H$ on $T$. Clearly, the most obvious choice of such a space is the RKHS $H_{X}$ of the the covariance function $k_{X}(s,t):=\mathbb{E}(X_{s}X_{t})$ of $X$. In this case, a classical result, which goes back to Parzen [33], Kallianpur [22], and LePage [32] shows that
$\mathrm{P}(\{X_{\bullet}\in H_{X}\})=1$
if and only if $\dim H_{X}<\infty$. In other words, in basically no situation of interest is $H_{X}$ an RKHS containing the paths of $X$. However, $H_{X}$ is, of course, by no means the only RKHS on $T$ and beginning with Driscoll [12], and subsequently Fortet [18], more general $H$ have been considered. The most general investigation in this direction has been conducted by Lukić and Beder [28] and since their results provide a basis of our analysis, we quickly recall their findings. To this end, we need the following notion.
Definition 1.1.
Let $H_{1}$ and $H_{2}$ be RKHSs on $T$. Then we write $H_{1}\ll H_{2}$ and speak of nuclear dominance of $H_{2}$ over $H_{1}$, if $H_{1}\subset H_{2}$ and the resulting embedding map $J:H_{1}\to H_{2}$ is a Hilbert-Schmidt operator.
Note that formally, [28] defines nuclear dominance differently, see the discussion following Theorem 6.2 below, but Lemma 6.3 shows that our definition is actually equivalent to the one used in [28]. With these preparations, [28, Theorem 7.4] provides the following characterization for a fixed RKHS $H$.
Theorem 1.2.
Let $X:=(X_{t})_{t\in T}$ be a centered Gaussian process on a complete probability space $(\Omega,{\cal A},\mathrm{P})$ and $H$ be an RKHS on $T$. Then the following statements hold true:
  1. i)
    If $H_{X}\ll H$, then there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
  2. ii)
    If $H_{X}\not\ll H$, then for all versions $Y$ of $X$ we have $\mathrm{P}(\{Y_{\bullet}\in H\})=0$.
Here we note that strictly speaking [28, Theorem 7.4] only shows ii) for $Y=X$. However, if we have an arbitrary version $Y$ of $X$, then $Y$ is a centered Gaussian process with $k_{Y}=k_{X}$, which in turn gives $H_{Y}=H_{X}$. Applying the original result [28, Theorem 7.4] to $Y$ then yields ii).
Informally speaking, given a fixed “candidate” RKHS $H$, Theorem 1.2 shows that we have either
$\mathrm{P}(\{X_{\bullet}\in H\})=1\qquad\qquad\mbox{ or }\qquad\qquad\mathrm{P}(\{X_{\bullet}\in H\})=0\,$,
and in addition both situations are characterized by $H_{X}\ll H$, respectively $H_{X}\not\ll H$. In this sense, it provides a “test” for each RKHS $H$ on $T$. However, it does not tell us how to find an $H$ with $H_{X}\ll H$, nor whether $H_{X}\ll H$ is even possible for suitable RKHSs $H$.
In this respect we note that positive, constructive results can sometimes be derived if we already have some knowledge on the paths of $X$. To illustrate this, let $E$ be a Banach space of functions (BSF) on $T$, that is, $E$ is a Banach space consisting of functions $f:T\to\mathbb{R}$ such that the point evaluations $\delta_{t}:E\to\mathbb{R}$ given by $\delta_{t}(f):=f(t)$ for all $t\in T$ and $f\in E$, are continuous. If $(\Omega,{\cal A},\mathrm{P})$ is complete and we know that there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in E\})=1$, then for every “surrounding” RKHS $H$ on $T$, that is $E\subset H$, we obviously also have $\mathrm{P}(\{Y_{\bullet}\in H\})=1$. Unfortunately, however, this technique is limited to “sufficiently benign” $E$, since there are various important BSFs $E$, for which no such surrounding $H$ exists. For example, for the somewhat natural candidate $E=C(T)$, where $C(T)$ denotes the space of continuous functions $T\to\mathbb{R}$ on a compact metric space $T$, we cannot find a surrounding $H$ as soon as $T$ is not countable, see e.g. [45]. Moreover, [42] has recently shown that even for many spaces $E$ of smoother functions no such surrounding $H$ with bounded kernel may exist. For example, if $\operatorname{H\ddot{o}l}^{\beta}([0,1])$ denotes the space of all $\beta$-Hölder continuous functions $f:[0,1]\to\mathbb{R}$, then there exists an RKHS $H$ with bounded kernel and $\operatorname{H\ddot{o}l}^{\beta}([0,1])\subset H$ if $\beta>1/2$, while there is no such RKHS if $\beta<1/2$, see [42, Section 4.5, Theorem 22, and Lemma 24, respectively Theorem 12]. In particular, the approach of taking a surrounding RKHS does not work for the Wiener process, while for the fractional Brownian motions with Hurst index >1/2>1/2 it does work, see also Example 5.4. In summary, the approach of using a surrounding RKHS can only be used if we already know that the paths are contained in a sufficiently small BSF $E$.
A different approach for constructing an RKHS $H$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ for some suitable version $Y$ of $X$ has been taken in [44]. In a nutshell, this approach is based on a refined analysis of the Karhunen-Loève expansion for $X$. Unfortunately, this approach, as well as its recent refinement in [23], again fails to produce positive results in many situations. In particular, it does not provide an RKHS containing the paths of the Wiener process.
The example of the Wiener process suggests that for some Gaussian processes $X$ it may be impossible to find an RKHS containing the paths of $X$. To the best of our knowledge, however, such general impossibility results, even for concrete processes, have not been established yet, despite the fact that Gaussian processes with paths in an RKHS have attracted interests in diverse areas of applied mathematics, see e.g. [37, 17, 20, 49] and the references mentioned in [23]. In particular, Theorem 1.2 cannot be directly used, since we would have to check $H_{X}\ll H$ for all RKHS $H$ on $T$, or at least for a sufficiently rich class of such RKHSs. Similarly, a failure of the described two approaches for finding an $H$ containing the paths of the considered process does, of course, not imply a general impossibility result.
The main goal of this paper is close this gap by characterizing when for a Gaussian process $X$ there exists an RKHS $H$ and a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$. To this end, we will first present a necessary condition for the existence of $H$ with $H_{X}\ll H$. Combining this necessary condition with the two approaches discussed above for finding $Y$ and $H$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ we can then characterize for various families of Gaussian processes, when such $Y$ and $H$ exist.
The rest of this paper is organized as follows: In Section 2 we collect the notions and notations that are necessary for understanding our results. The general theory is then presented in Section 3, and in Section 4 we apply our methods to Gaussian processes $X:=(X_{t})_{t\in T}$ whose RKHSs $H_{X}$ are essentially equal to a fractional Sobolev space on $T\subset\mathbb{R}^{d}$. In addition, the case of Sobolev spaces with dominating mixed smoothness is also investigated. Moreover, some concrete examples are presented in Section 5. Finally, all proofs as well as some additional material needed for these proofs can be found in Section 6.
2 Preliminaries
For sequences $(a_{n})_{n\geq 1}$ and $(b_{n})_{n\geq 1}\subset(0,\infty)$ we write $a_{n}\sim b_{n}$ if there exist constants $c_{1},c_{2}>0$ with $c_{1}a_{n}\leq b_{n}\leq c_{2}a_{n}$ for all $n\geq 1$.
If $E$ and $F$ are Banach spaces, we write $E\hookrightarrow F$, if $E\subset F$ and the resulting embedding map is continuous. In addition, we write $E\circeq F$ if $E\subset F\subset E$, and we write $E\cong F$ if $E\circeq F$ and the norms of $E$ and $F$ are equivalent. In other words, we have $E\cong F$ if and only if $E\hookrightarrow F\hookrightarrow E$. Obviously $E\cong F$ implies $E\circeq F$, and if $E$ and $F$ are BSFs, then the converse implication is also true by a simple application of the closed graph theorem, see e.g. [42, Lemma 23]. For RKHSs $H_{1}$ and $H_{2}$ on $T$ we thus have $H_{1}\circeq H_{2}$ if and only if $H_{1}\cong H_{2}$.
[…]
$I_{\nu}:{{\cal L}_{\infty}({\cal B})}$ $\to{L_{\infty}(\nu)}$ (1)
$f$ $\mapsto[f]_{\sim}\,$,
is linear and bounded with $\|[f]_{\sim}\|_{\infty}\leq\|f\|_{\infty}$ for all $f\in{{\cal L}_{\infty}({\cal B})}$, where $\|[f]_{\sim}\|_{\infty}$ denotes the usual essential supremum norm of $f]_{\sim}$ in $L_{\infty}(\nu)$}. Finally, if $\lambda^{d}$ is the Lebesgue measure on $\mathbb{R}^{d}$ and $T\subset\mathbb{R}^{d}$ is measurable with $\lambda^{d}(T)>0$, then we often write ${L_{2}(T)}:={L_{2}(\lambda^{d}_{|T})}$, where $\lambda^{d}_{|T}$ is the Lebesgue measure on $T$. Analogously, we write ${{\cal L}_{\infty}(T)}:={{\cal L}_{\infty}(\mathfrak{B}^{d}_{|T})}$, where $\mathfrak{B}^{d}_{|T}$ denotes the restriction of the Borel $\sigma$-algebra $\mathfrak{B}^{d}$ on $\mathbb{R}^{d}$ to $T$.
[…]
$I_{k,\nu}:H$ $\to{L_{2}(\nu)}$ (2)
$h$ $\mapsto[h]_{\sim}$
is well-defined and continuous. Finally, we say that $\nu$ is $H$-positive, if $I_{k,\nu}$ is injective, that is, $\nu(\{h\neq 0\})>0$ for all $h\in H$ with $h\neq 0$. For example, if $H$ consists of continuous functions and $\nu(U)>0$ for all non-empty, open $U\subset T$, then $\nu$ is $H$-positive. On the other hand, if e.g. $\nu$ is an empirical measure based upon $n$ points and $\dim H=\infty$, then $I_{k,\nu}$ cannot be injective.
Recall that a probability measure $\mathrm{P}$ on $\mathbb{R}$ is a Gaussian measure, if there are $\mu\in\mathbb{R}$ and $\sigma\in[0,\infty)$ such that the characteristic function of $\mathrm{P}$ is given by
$\varphi_{\mathrm{P}}(t)=\mathrm{e}^{\mathrm{i}t\mu}\cdot\exp\biggl(-\frac{\sigma^{2}t^{2}}{2}\biggr)\,,\qquad\qquad t\in\mathbb{R}.$
Clearly, if $\sigma>0$, then $\mathrm{P}$ is a normal distribution with expectation $\mu$ and variance $\sigma^{2}$. In the case $\sigma=0$, the probability measure $\mathrm{P}$ is a point mass on $\mu$, that is $\mathrm{P}(\{\mu\})=1$.
Now let $X=(X_{t})_{t\in T}$ be an $\mathbb{R}$-valued, stochastic process on some probability space $(\Omega,{\cal A},\mathrm{P})$. As usual, we call the map $X_{\bullet}(\omega):t\mapsto X_{t}(\omega)$, where $\omega\in\Omega$ is fixed, a path of $X$. Moreover, $X$ is called a Gaussian process, if for all $t_{1},\dots,t_{n}\in T$ and $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$, the distribution of $\alpha_{1}X_{t_{1}}+\dots+\alpha_{n}X_{t_{n}}$ is a Gaussian measure on $\mathbb{R}$. Furthermore, the Gaussian process is centered, if $\mathbb{E}X_{t}=0$ for all $t\in T$. In this case, its covariance function $k_{X}:T\times T\to\mathbb{R}$ is given by
$k_{X}(t_{1},t_{2}):=\mathbb{E}(X_{t_{1}}X_{t_{2}})\,,\qquad\qquad t_{1},t_{2}\in T.$
It is well known and easy to verify that $k_{X}$ is a symmetric and positive definite function, and thus a (reproducing) kernel. We write $H_{X}$ for the corresponding RKHS of $k_{X}$ and we further use the short hand $I_{X,\nu}:=I_{k_{X},\nu}$.
3 General Results
In this section we present and investigate necessary and sufficient conditions for the existence of RKHSs containing the paths of a given Gaussian process.
To this end, we first need a notion that controls the growth behavior of the functions in a given RKHS.
Definition 3.1.
Let $H$ be an RKHS on $T$. Then we say that a function $b:T\to[0,\infty)$ is an envelope for $H$ if, for all $h\in H$ and $t\in T$, we have
$|h(t)|\leq\|h\|_{H}\cdot b(t)\,.$ (3)
At first glance, Condition (3) my look restrictive. However, given an arbitrary RKHS $H$ on $T$, a simple estimate yields
$|h(t)|=\bigl|\langle h,k(\,\cdot\,,t)\rangle_{H}\bigr|\leq\|h\|_{H}\cdot\|k(\,\cdot\,,t)\|_{H}=\|h\|_{H}\cdot\sqrt{k(t,t)}$ (4)
for all $h\in H$ and $t\in T$. In other words, every RKHS has an envelope function, namely the diagonal function $t\mapsto\sqrt{k(t,t)}$. In general, however, the measurability of this diagonal function does not follow from the measurability of all $h\in H$, and this missing implication may cause some technical issues if we would directly work with the diagonal function. Moreover, the diagonal function is closely related to the considered RKHS, whereas choosing an envelope function is to some extend independent of a specific RKHS. For example, by considering all RKHSs $H$ on $T$ having the constant envelope function $b=1$, we actually work with all RKHSs $H$ satisfying $\|\operatorname{id}:H\to\ell_{\infty}(T)\|\leq 1$. For more information on the relation between diagonal and envelope functions we refer to Lemma 6.7.
[…]
Theorem 3.2.
Let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ with RKHS $H_{X}$. Moreover, let $H$ be an RKHS on $T$ and $Y$ be a version of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
[…]
Note that there is a trade-off between the considered class of RKHS $H$ in Theorem 3.2 and the strength of the derived necessary condition. Indeed, if we have two measurable envelope functions $b_{1}$ and $b_{2}$ with $b_{1}\leq b_{2}$, then every RKHS $H$ having envelope $b_{1}$ also has envelope $b_{2}$. In contrast, not every measure $\nu$ with $b_{1}\in{{\cal L}_{2}(\nu)}$ also needs to satisfy $b_{2}\in{{\cal L}_{2}(\nu)}$. In other words, by increasing the envelope function we potentially shrink the class of considered operators $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$. An extreme case in this trade-off are the constant envelope functions that describe RKHSs containing bounded functions. The following corollary presents the results of Theorem 3.2 for such envelopes.
Corollary 3.3.
[…]
$\mathrm{P}(\{Y_{\bullet}\in H\})>0$
then, for all finite measures $\nu$ on $(T,{\cal B})$, the operator $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is well-defined and nuclear.
Note that there is a conceptual difference between Theorems 1.2 and 3.2, respectively Corollary 3.3. Indeed, Theorem 1.2 provides a “test” for every single candidate RKHS $H$ as discussed in the introduction. In contrast, Theorem 3.2 and its Corollary 3.3 provide a general impossibility result: For example, if for a given $H_{X}\subset{{\cal L}_{\infty}({\cal B})}$ we find a finite measure $\nu$ on $(T,{\cal B})$ for which $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is not nuclear, then there exist no RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ and no version $Y$ of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
While at first glance finding such a $\nu$ may seem to be a hard task, we will see in Sections 4 and 5 that for e.g. bounded, measurable $T\subset\mathbb{R}^{d}$ with non-empty interior the Lebesgue measure is typically a sufficient choice. In fact, we will leverage there from well-established results on the “embedding” $I_{X,\nu}:H_{X}\to{L_{2}(T)}$ if $H_{X}$ belongs to particular classes of RKHSs such as fractional Sobolev spaces.
We finish this discussion with another corollary showing that a certain “localization” reduces the problem to bounded RKHS $H$ even if we are interested in more general RKHSs.
Corollary 3.4.
Let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ with RKHS $H_{X}$. Moreover, let $H$ be an RKHS on $T$ and $Y$ be a version of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
[…]
$I_{X_{|T^{*}},\nu}:H_{X_{|T^{*}}}\to{L_{2}(\nu)}$
is nuclear, where $H_{X_{|T^{*}}}\subset{{\cal L}_{\infty}({\cal B}_{|T^{*}})}$ denotes the RKHS of the restricted process $X_{|T^{*}}:=(X_{t})_{t\in T^{*}}$.
Note that on $T^{*}$, every function $f\in H$ is bounded, namely we have $|f(t)|\leq\|f\|_{H}\cdot M$ for all $t\in T^{*}$. In other words, the restriction of $H$ onto $T^{*}$ consists of bounded, measurable functions, and therefore we are actually in the situation of Corollary 3.3. The proof of Corollary 3.4 uses this observation.
In some other situations, however, the RKHS $H_{X}$ and the operator $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ may not be as amenable as desired. In the following, we will thus present two equivalent characterizations of $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ being nuclear, which sometimes make it easier to work with this property.
We will begin with a simple yet useful characterization that makes it possible to alter the space $H_{X}$ by some finite dimensional space. To this end, we need the following definition.
Definition 3.5.
Let $H_{1}$ and $H_{2}$ be RKHSs on $T$. Then we say that $H_{1}$ and $H_{2}$ are essentially equivalent, if there exist RKHSs $H_{1}^{\star}$ and $H_{2}^{\star}$ on $T$ with $\dim H_{1}^{\star}<\infty$, $\dim H_{2}^{\star}<\infty$, and
$H_{1}+H_{1}^{\star}\circeq H_{2}+H_{2}^{\star}\,.$ (5)
It is straightforward to check that the notion of being essentially equivalent gives an equivalence relation on the set of all RKHSs on $T$, but in the following we do not need this. In addition, another simple calculation shows that if $H_{1}$ and $H_{2}$ are essentially equivalent, then both $H_{1}$ and $H_{2}$ are also essentially equivalent to $H_{1}+H_{2}$. Furthermore, RKHSs $H_{1}$ and $H_{2}$ with $H_{1}\circeq H_{2}$ are obviously also essentially equivalent. Finally, the following lemma shows that in some situations the essential equivalence can be easily verified with the help of the kernels of the involved RKHSs.
Lemma 3.6.
Let $H_{1}$ and $H_{2}$ be RKHSs on $T$ with kernels $k_{1}$ and $k_{2}$, and $k_{1}^{\star}$ and $k_{2}^{\star}$ be kernels on $T$ with finite dimensional RKHSs. Then $H_{1}$ and $H_{2}$ are essentially equivalent, if we have
$k_{1}-k_{2}=k_{2}^{\star}-k_{1}^{\star}\,.$
Now, the following result shows that $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is nuclear if and only if $I_{H,\nu}:H\to{L_{2}(\nu)}$ is nuclear for all RKHS $H$ that are essentially equivalent to $H_{X}$. This will make it possible to switch from $H_{X}$ to an essentially equivalent $H$ for which $I_{H,\nu}:H\to{L_{2}(\nu)}$ is better understood.
Theorem 3.7.
Let $(T,{\cal B},\nu)$ be a measure space and $H_{1}$ and $H_{2}$ be essentially equivalent RKHSs on $T$ with $H_{1}\subset{{\cal L}_{0}({\cal B})}$ and $H_{1}\subset{{\cal L}_{0}({\cal B})}$. If, in addition, $b\in{{\cal L}_{2}(\nu)}$ is an envelope function for both $H_{1}$ and $H_{2}$, then $I_{H_{1},\nu}:H_{1}\to{L_{2}(\nu)}$ is nuclear if and only if $I_{H_{2},\nu}:H_{2}\to{L_{2}(\nu)}$ is nuclear.
Theorem 3.7 can be used to establish non-existence results with the help of Theorem 3.2 and its Corollary 3.3, see e.g. Example 5.2. Combining the next result with Theorem 1.2 shows that in some situations the essential equivalence of RKHSs can also be used to find RKHSs containing the paths of a given Gaussian process, see e.g. Theorem 4.1.
Theorem 3.8.
Let $X:=(X_{t})_{t\in T}$ be a centered Gaussian process on a probability space $(\Omega,{\cal A},\mathrm{P})$ and $H$ be an RKHS that is essentially equivalent to $H_{X}$. If there exists an RKHS $H^{\dagger}$ on $T$ with $H\ll H^{\dagger}$, then there also exists an RKHS $H_{X}^{\dagger}$ on $T$ that is essentially equivalent to $H^{\dagger}$ and satisfies both $H_{X}^{\dagger}\subset H_{X}+H^{\dagger}$ and
$H_{X}\ll H_{X}^{\dagger}\,.$
Our next goal is to present another equivalent way to express that $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is nuclear as well as to construct an $H$ with $H_{X}\ll H$ in some situations. To this end, we assume that $b\in{{\cal L}_{2}(\nu)}$ is an envelope function for $H_{X}\subset{{\cal L}_{0}({\cal B})}$, where we note that in the situation of Theorem 3.2, we do have $H_{X}\subset H$ by Theorem 1.2, and therefore the envelope function of $H$ in Theorem 3.2 is, modulo a constant, also an envelope function for $H_{X}$. In this sense, the assumption of $b$ being an envelope for $H_{X}$ fits to our previous results. Let us now consider the operator $T_{X,\nu}:=I_{X,\nu}\circ I_{X,\nu}^{*}:{L_{2}(\nu)}\to{L_{2}(\nu)}$, where $I_{X,\nu}^{*}:{L_{2}(\nu)}\to H_{X}$ denotes the adjoint of $I_{X,\nu}$. Then some elementary calculations, see Lemma 6.8, show that $T_{X,\nu}$ is given by
$T_{X,\nu}f=\int_{T}k_{X}(\,\cdot\,,t)f(t)\,\mathrm{d}\nu(t)\,,\qquad\qquad f\in{L_{2}(\nu)},$
where the right-hand side is viewed as a $\nu$-equivalence class in $L_{2}(\nu)$}. By construction, $T_{X,\nu}$ is positive and self-adjoint, and it is also compact, since $I_{X,\nu}$ is Hilbert-Schmidt by our assumption of $b\in{{\cal L}_{2}(\nu)}$ being an envelope for $H_{X}$, see again Lemma 6.8. Consequently, the spectral theorem, see Section 6.1, shows that $T_{X,\nu}$ has at most countably many non-zero eigenvalues $(\mu_{i}(T_{X,\nu}))_{i\in I}$, where we have set either $I:=\{1,\dots,n\}$ or $I:=\mathbb{N}$, and where we have included the geometric multiplicities. Moreover, these eigenvalues are positive and there exists an orthonormal system (ONS) $([f_{i}]_{\sim})_{i\in I}\subset{L_{2}(\nu)}$ such that each $f_{i}]_{\sim}$ is an eigenfunction of $T_{X,\nu}$ with eigenvalue $\mu_{i}:=\mu_{i}(T_{X,\nu})$. We now define
$e_{i}:=\mu_{i}^{-1}I_{X,\nu}^{*}[f_{i}]_{\sim}\,,\qquad\qquad i\in I.$ (6)
This gives $e_{i}\in H_{X}$ and $e_{i}]_{\sim}=I_{X,\nu}e_{i}=\mu_{i}^{-1}T_{X,\nu}[f_{i}]_{\sim}=[f_{i}]_{\sim}$ for all $i\in I$. In other words, we may choose representatives $e_{i}$ of $f_{i}]_{\sim}$ that are contained in $H$. Finally, $T_{X,\nu}$ is nuclear since $I_{X,\nu}$ is Hilbert-Schmidt, see Lemma 6.1, and therefore $(\mu_{i}(T_{X,\nu}))_{i\in I}$ is summable.
Now recall from [47] that in some cases the eigenfunctions and -values can be used to construct RKHSs surrounding $H_{X}$. To briefly describe this construction, let us fix a $\beta>0$ for which we have
$\sum_{i\in I}\mu_{i}^{\beta}\,e_{i}^{2}(t)<\infty\,,\qquad\qquad t\in T.$ (7)
Then [47, Lemma 2.6] shows that
$k_{X}^{\beta}(t,t^{\prime}):=\sum_{i\in I}\mu_{i}^{\beta}e_{i}(t)e_{i}(t^{\prime})\,,\qquad\qquad t,t^{\prime}\in T$ (8)
defines a kernel on $T$ whose RKHS is given by
$H_{X}^{\beta}:=\biggl\{\sum_{i\in I}a_{i}\mu_{i}^{\beta/2}e_{i}:(a_{i})_{i\in I}\in\ell_{2}(I)\biggr\}\,$,
where the series converges pointwise and $\ell_{2}(I)$ denotes the space of all 2-summable families $(a_{i})_{i\in I}\subset\mathbb{R}$. Moreover, $(\mu_{i}^{\beta/2}e_{i})_{i\in I}$ is an ONB of $H_{X}^{\beta}$ and we have $H_{X}^{\beta}\subset{{\cal L}_{0}({\cal B})}$. Obviously, Condition (7) is monotone in the sense that if it holds for some $\beta$ then it also holds for all exponents $\alpha>\beta$. In addition, for such pairs we have $H_{X}^{\alpha}\subset H_{X}^{\beta}$. Finally, if $\nu$ is $H$-positive, then $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ is an ONB of $H_{X}$, see [47, Theorem 3.1], and thus we have $H_{X}^{1}=H_{X}$. Consequently, if in this case we have found a $\beta\in(0,1)$ with (7) then $H_{X}\subset H_{X}^{\beta}$.
With these preparations we can now present our last result of this section, where we note that ii) essentially simplifies similar results of [44, Section 5].
Theorem 3.9.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,{\cal B},\nu)$ be a measure space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ such that $b\in{{\cal L}_{2}(\nu)}$ is an envelope for $H_{X}$ and $H_{X}\subset{{\cal L}_{0}({\cal B})}$. Moreover, let $(\mu_{i})_{i\in I}$ be the non-zero eigenvalues of $T_{X,\nu}$ and $(e_{i})_{i\in I}\subset H_{X}$ be a family such that $([e_{i}]_{\sim})_{i\in I}\subset{L_{2}(\nu)}$ is an ONS in $L_{2}(\nu)$} and each $e_{i}]_{\sim}$ is an eigenfunction of $T_{X,\nu}$ to the eigenvalue $\mu_{i}$. Then the following statements hold true:
  1. i)
    The operator $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is nuclear if and only if
    $\sum_{i\in I}\sqrt{\mu_{i}}<\infty.$ (9)
  2. ii)
    Assume that $\nu$ is $H_{X}$-positive and we have a $\beta\in(0,1)$ satisfying (7). Then $H_{X}\ll H_{X}^{\beta}$ if and only if
    $\sum_{i\in I}\mu_{i}^{1-\beta}<\infty.$ (10)
    In this case, there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H_{X}^{\beta}\})=1$.
To illustrate Theorem 3.9 we assume for a moment that $\nu$ is $H_{X}$-positive and $\sup_{i\in I}\|e_{i}\|_{\infty}<\infty$. Then, if (9) is not satisfied, Theorem 3.9 shows that $I_{X,\nu}:H_{X}\to{L_{2}(\nu)}$ is not nuclear, and therefore Corollary 3.3 shows that there exist no RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ and no version $Y$ of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
Conversely, if (9) is satisfied, then both (7) and (10) hold true for $\beta:=1/2$, and hence there exists a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H_{X}^{1/2}\})=1$. Furthermore, note that in this case $k_{X}^{1/2}$ is bounded, and hence we have $H_{X}^{1/2}\subset{{\cal L}_{\infty}({\cal B})}$. In other words, if we have an $H_{X}$-positive, finite measure $\nu$ with $\sup_{i\in I}\|e_{i}\|_{\infty}<\infty$, then (9) characterizes the existence of an RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ containing the paths of a suitable version of $X$. We refer to Example 5.7 for such a situation.
Finally, recall from [47, Theorem 4.6] that the spaces $H_{X}^{\beta}$ can in some sense be identified with the interpolation spaces ${L_{2}(\nu)},[H_{X}]_{\sim}]_{\beta,2}$ of the real method. In particular, if under some mild technical assumptions we have ${L_{2}(\nu)},[H_{X}]_{\sim}]_{1/2,2}\hookrightarrow{L_{\infty}(\nu)}$, then $H_{X}^{1/2}$ exists and is an RKHS version of ${L_{2}(\nu)},[H_{X}]_{\sim}]_{1/2,2}$, see [44, Theorem 5.5], or its recent refinement [6, Theorem 3.4], in combination with [47, (36) and Theorem 4.6], while [47, Theorem 5.3] ensures (9). In such situations it may thus be possible to construct $H_{X}^{1/2}$ modulo equivalent norms without using the eigenfunctions and -values. In this respect note that for certain kernels $k_{X}$ [17] presents another way to find $H_{X}^{1/2}$ without the eigenfunctions and -values.
Finally, [44, Section 1] lists various articles in which the eigenfunctions and -values are computed for specific Gaussian processes. For such processes, we can then apply Theorem 3.9 directly.
4 Results for RKHSs of Sobolev Type
In this section we investigate how our main results can be used to investigate Gaussian processes whose RKHSs are essentially equivalent to a fractional Sobolev space, or equal a Sobolev space of dominating mixed smoothness. To this end, recall that for $s\in(0,\infty)$, the fractional Sobolev spaces, or Bessel potential spaces, are defined by
$H^{s}(\mathbb{R}^{d}):=\bigl\{f\in{L_{2}(\mathbb{R}^{d})}:(1+\|\cdot\|^{2})^{s/2}\hat{f}\in{L_{2}(\mathbb{R}^{d})}\bigr\}\,$,
where $\hat{f}$ is the Fourier transformation of $f$, see e.g. [30, Section 3.1] or [50, Section 2.2.2]. By setting
$\|f\|_{H^{s}(\mathbb{R}^{d})}:=\bigl\|\,(1+\|\cdot\|^{2})^{s/2}\hat{f}\,\bigr\|_{{L_{2}(\mathbb{R}^{d})}}$
the space $H^{s}(\mathbb{R}^{d})$ becomes a Hilbert space and for $m\in\mathbb{N}$ we have $H^{m}(\mathbb{R}^{d})\cong W^{m}_{2}(\mathbb{R}^{d})$, where $W^{m}_{2}(\mathbb{R}^{d})$ denotes the classical Sobolev space with smoothness $m$, see e.g. [50, (3) in Chapter 2.5.6] or [1, Chapter 7, Paragraph 7.62]. Moreover, the definition of the fractional Sobolev spaces immediately ensures $H^{s}(\mathbb{R}^{d})\subset H^{r}(\mathbb{R}^{d})$ for all $0<r<s<\infty$, and for $s>d/2$ we have
$H^{s}(\mathbb{R}^{d})\hookrightarrow C_{b}(\mathbb{R}^{d})$ (11)
by Sobolev’s embedding theorem, see the beginning of Section 6.7 for details and references. In other words, in the case $s>d/2$ the space $H^{s}(\mathbb{R}^{d})$ can be viewed as an RKHS with bounded and continuous kernel. Finally, for bounded, measurable $T\subset\mathbb{R}^{d}$ and $s>d/2$ we define $H^{s}(T):=\{f_{|T}:f\in H^{s}(\mathbb{R}^{d})\}$ and equip this space with the norm
$\|h\|_{H^{s}(T)}:=\inf\bigl\{\|f\|_{H^{s}(\mathbb{R}^{d})}:f\in H^{s}(\mathbb{R}^{d})\mbox{ with }f_{|T}=h\bigr\}\,.$ (12)
Here, we followed [15, Chapter 2.5.1] and in particular we refer to the discussion there, when such restricted spaces also have an intrinsic description. In any case, as inclusions between BSFs $E$ and $F$ on some set $T_{0}$ remain unchanged when both spaces are restricted to some subset $T\subset T_{0}$, see e.g. [42, Proposition 3] for a general statement, the inclusions above lead to $H^{s}(T)\hookrightarrow H^{r}(T)\hookrightarrow C_{b}(T)\hookrightarrow{{\cal L}_{\infty}(T)}$ for all $s>r>d/2$. In this case, both $H^{s}(T)$ and $H^{r}(T)$ can thus be viewed to be RKHSs.
With these preparations, we can now exactly describe those Gaussian processes that have an RKHS that is essentially equivalent to a fractional Sobolev space and whose paths are contained in some RKHS.
Theorem 4.1.
Let $T\subset\mathbb{R}^{d}$ be bounded and measurable with non-empty interior. Moreover, let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ whose RKHS $H_{X}$ consists of continuous functions and is essentially equivalent to $H^{s}(T)$ for some $s>d/2$. Then the following statements are equivalent:
  1. i)
    We have $s>d$.
  2. ii)
    There exist an RKHS $H\subset{{\cal L}_{\infty}(T)}$ and a version $Y$ of $X$ such that $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
  3. iii)
    There exist an RKHS $H$ on $T$ and a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
In this case, for all $r\in(d/2,s-d/2)$ there actually exists such an $H$ that is essentially equivalent to $H^{r}(T)$.
As the implication from ii) to i) is based upon Corollary 3.3 we see that the necessary condition derived there is sharp in the setup of Theorem 4.1. Below in Example 5.6 we apply Theorem 4.1 to the class of Matérn kernels that are popular in many applications, see e.g. [39, 38].
Theorem 4.1 shows that we require a strong, dimension dependent regularity to ensure that the paths are contained in an RKHS. Interestingly, such a dimension dependent regularity disappears if we switch to (tensor) product kernels, which are used in e.g. efficient implementations of Gaussian processes for machine learning [19, 16].
Theorem 4.2.
Let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space, $T=(0,1)^{d}$, and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ whose covariance function is of the form
$k_{X}(t,t^{\prime})=\prod_{i=1}^{d}k_{0}(t_{i},t_{i}^{\prime})\,,\qquad\qquad t,t^{\prime}\in T,$
where $k_{0}:(0,1)\times(0,1)\to\mathbb{R}$ is a kernel with RKHS $H_{0}\circeq H^{s}(0,1)$ for some $s>1/2$. Then the following statements are equivalent:
  1. i)
    We have $s>1$.
  2. ii)
    There exist an RKHS $H\subset{{\cal L}_{\infty}(T)}$ and a version $Y$ of $X$ such that $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
In this case, we may choose a Sobolev space of dominating mixed smoothness $H\cong B^{r}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{r}_{2,2}(0,1)$ for any $r\in(1/2,s-1/2)$.
5 Some Concrete Examples
In this section we present some explicit examples of Gaussian processes that do or do not have their paths in a suitable RKHS. To this end, we assume throughout this section that the considered Gaussian processes are defined on some complete probability space $(\Omega,{\cal A},\mathrm{P})$. Moreover, $\lambda$ and $\lambda^{d}$ denote the Lebesgue measures on the considered subsets $T\subset\mathbb{R}$, respectively $T\subset\mathbb{R}^{d}$.
With these preparations, our first example considers the Wiener process.
Example 5.1.
Let $T:=[0,1]$ and $W:=(W_{t})_{t\in T}$ be the Wiener process. Then $I_{W,\lambda}:H_{W}\to{L_{2}(T)}$ is not nuclear and there exists no RKHS $H$ and $\mathrm{P}(\{W_{\bullet}\in H\})=1$.
The proof of Example 5.1 shows that $I_{W,\lambda}:H_{W}\to{L_{2}(T)}$ fails to be nuclear by a thin margin. This observation will be further investigated in Example 5.4, which considers the fractional Brownian motion.
Since the Brownian bridge $(B_{t})_{t\in T}$ on $T:=[0,1]$ can be defined by $B_{t}:=(1-t)W_{t}$, it is not surprising that the negative result of Example 5.1 carries over to the Brownian bridge. The details are formulated in the following example, whose proof also illustrates the usefulness of the notion of essentially equivalent RKHSs.
Example 5.2.
Let $T:=[0,1]$ and $(B_{t})_{t\in T}$ be the Brownian bridge. Then there exists no RKHS $H$ such that $\mathrm{P}(\{B_{\bullet}\in H\})=1$.
Like the Brownian bridge, the Ornstein-Uhlenbeck process is closely related to the Wiener process, and hence it is again not surprising that the negative result of Example 5.1 carries over to the Ornstein-Uhlenbeck process. The following example presents the details for two variants of this process.
Example 5.3.
Let $T:=[0,1]$ and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process whose covariance function is given by either
$k_{X}^{(1)}(t_{1},t_{2})$ $:=a\exp(-\sigma|t_{1}-t_{2}|)\,$,
or
$k_{X}^{(2)}(t_{1},t_{2})$ $:=a\exp(-\sigma|t_{1}-t_{2}|)-a\exp(-\sigma(t_{1}+t_{2}))\,$,
where $a,\sigma>0$ are some constants. Then there exist no RKHS $H$ and no version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$.
Our next example generalizes Example 5.1 to the fractional Brownian motion. Here it turns out that the results depend on the Hurst parameter.
Example 5.4.
Let $T:=[0,1]$ and $\alpha\in(0,1)$. Moreover, let $B^{(\alpha)}=(B^{(\alpha)}_{t})_{t\in T}$ be the fractional Brownian motion with Hurst index $\alpha$, that is, the centered Gaussian process with continuous paths and covariance function
$k_{B^{(\alpha)}}(t_{1},t_{2}):=\frac{1}{2}\bigl(|s|^{2\alpha}+|t|^{2\alpha}-|t-s|^{2\alpha}\bigr)\,,\qquad\qquad t_{1},t_{2}\in[0,1]\,.$
Then there exists an RKHS $H$ with $\mathrm{P}(\{B_{\bullet}^{(\alpha)}\in H\})=1$ if and only if $\alpha>1/2$. Moreover, in the case $\alpha>1/2$ we may choose $H=H^{s}(T)\subset{{\cal L}_{\infty}(T)}$ for any $s\in(1/2,\alpha)$.
In the case $\alpha<1/2$ the paths of the fractional Brownian motion are “rougher” than the paths of the Wiener process. Since we have already seen in Example 5.1 that the paths of the Wiener process are not contained in an RKHS, it is therefore not surprising that the paths of $B^{(\alpha)}$ are not contained in an RKHS, either. In the case $\alpha>1/2$, it is well-known that the paths of $B^{(\alpha)}$ are $\beta$-Hölder-continuous for all $\beta\in(0,\alpha)$. The positive result then follows by standard embeddings between Besov spaces, see the proof of Example 5.4 for details.
The fractional Brownian motion is closely related to the Riemann-Liouville process, see e.g. [27, Example 3.4] or [34], which is considered in the following example.
Example 5.5.
Let $T:=[0,1]$ and $\alpha\in(0,1)$. Moreover, let $R^{(\alpha)}=(R^{(\alpha)}_{t})_{t\in T}$ be the Riemann-Liouville process with index $\alpha$, that is, the centered Gaussian process with continuous paths given by the Itô-Integral
$R^{(\alpha)}_{t}:=\frac{1}{\Gamma{(\alpha+1/2})}\int_{0}^{t}(t-s)^{\alpha-1/2}\,\mathrm{d}W_{s}\,,\qquad\qquad t\in[0,1].$
Then there exists an RKHS $H$ with $\mathrm{P}(\{R_{\bullet}^{(\alpha)}\in H\})=1$ if and only if $\alpha>1/2$. Moreover, in the case $\alpha>1/2$ we may choose $H=H^{s}(T)\subset{{\cal L}_{\infty}(T)}$ for any $s\in(1/2,\alpha)$.
The next example, which considers Gaussian processes on a higher dimensional $T$, uses our result for RKHSs of Sobolev type formulated in Theorem 4.1. Analogous results can thus be obtained for other kernels of Sobolev type, such as the Wendland kernels, which are often attractive from a computational point of view, see [51, Chapter 9].
Example 5.6.
Let $T\subset\mathbb{R}^{d}$ be bounded and measurable with non-empty interior. Moreover, let $X:=(X_{t})_{t\in T}$ be a centered Gaussian process whose covariance function is a Matérn kernel of order $\alpha>0$ and width $\sigma>0$, that is
$k_{X}(t_{1},t_{2})=\bigl(\sigma\|t_{1}-t_{2}\|\bigr)^{\alpha}h_{\alpha}\bigl(\sigma\|t_{1}-t_{2}\|\bigr)\,,\qquad\qquad t_{1},t_{2}\in T,$
where $h_{\alpha}$ denotes the modified Bessel function of the second type of order $\alpha$. Then there exist an RKHS $H$ and a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ if and only if $\alpha>d/2$. Moreover, in the case $\alpha>d/2$ we may choose $H=H^{r}(T)$ for any $r\in(d/2,\alpha)$.
To compare the smoothness of the paths and the elements of the RKHS in Example 5.6, we quickly assume that $\alpha>d/2$. Then the paths of $X$ are contained in $H^{\alpha-\varepsilon}(T)$ for all sufficiently small $\varepsilon>0$. Moreover, the proof of Example 5.6 uses that the RKHS $H_{\alpha,\sigma}$ of the Matérn kernel equals $H^{\alpha+d/2}$ up to equivalent norms. Consequently, there is a $d/2+\varepsilon$-gap, when the smoothness is measured in the Sobolev sense. This gap coincides with the one observed in [44], where it is also shown that it is essentially sharp, see also [23] for a more detailed analysis and [41] for similar results for weakly stationary processes. Finally, [44, Example 4.8] shows that in the case $d=1$ these Sobolev results improve classical results on the continuous differentiability of the paths of $X$, of e.g. [9].
Our last example considers stationary processes on compact Abelian groups $T$ with Haar measure $\lambda$. It mostly relies on the fact that for translation invariant, $\mathbb{C}$-valued kernels the characters of $T$ form on ONB of $L_{2}(\lambda,\mathbb{C})$} consisting of eigenvectors of $T_{k,\lambda}$.
Example 5.7.
Let $(T,+)$ be a compact Abelian group with Haar measure $\lambda$ and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process whose covariance function $k_{X}$ is measurable and satisfies
$k_{X}(t_{1},t_{2})=k_{X}(t_{1}-t_{2},0)\,,\qquad\qquad t_{1},t_{2}\in T.$
Then there exist an RKHS $H\subset{{\cal L}_{\infty}({\cal B})}$ and a version $Y$ of $X$ with $\mathrm{P}(\{Y_{\bullet}\in H\})=1$ if and only if
$\sum_{i\in I}\sqrt{\mu_{i}(T_{X,\lambda})}<\infty.$
Moreover, in this case we may choose $H=H_{X}^{1/2}$ and we have $H_{X}^{1/2}\subset C(T)$.
Since Example 5.7 is an application of Theorem 3.9, it can be easily extended to situations in which the spectral properties of $k_{X}$ are well understood in the sense of the discussion following Theorem 3.9. Such situations include e.g. isotropic Gaussian random fields on $\mathbb{S}^{2}$, and with some more work also on $\mathbb{S}^{d}$, if $\lambda$ is the surface measure. We refer to [26] for the details on the required spectral properties and to [21] for an application for neural networks.
6 Proofs
6.1 Hilbert-Schmidt and Nuclear Operators
In this subsection we briefly recall some aspects of compact operators between Hilbert spaces. To this end, let $H$ be a Hilbert space and $R:H\to H$ be a compact, positive, and self-adjoint operator. Then the spectral theorem for such operators, e.g. [24, Chapter V.2.3] or [52, Theorem VI.3.2], shows that we have at most countably many non-zero eigenvalues $(\mu_{i}(R))_{i\in I}$, where we have set either $I:=\{1,\dots,n\}$ or $I:=\mathbb{N}$, and where we have included the geometric multiplicities. Moreover, there exists an ONS $(e_{i})_{i\in I}$ in $H$ such that $e_{i}$ is an eigenvector of $R$ to the eigenvalue $\mu_{i}(R)$ for all $i\in I$. Finally, these eigenvalues are positive and in the case $I=\mathbb{N}$ they can be ordered to form a decreasing sequence converging to 0. In the following, we always assume a decreasing order. Finally, the spectral theorem leads to the following representation
$Rx=\sum_{i\in I}\mu_{i}(R)\langle x,e_{i}\rangle e_{i}\,,\qquad\qquad x\in H.$ (13)
In particular, we have $H=\ker R\oplus\overline{\operatorname{span}\{e_{i}:i\in I\}}^{H}$.
If $S:H_{1}\to H_{2}$ is a compact operator between two Hilbert spaces $H_{1}$ and $H_{2}$, we write $S^{*}:H_{2}\to H_{1}$ for the adjoint of $S$. Then the operator $S\circ S^{*}:H_{2}\to H_{2}$ is compact, positive, and self-adjoint, and hence we can apply the spectral theorem to $S\circ S^{*}$. By doing so, we can define the singular numbers of $S$ by
$s_{i}(S):=\sqrt{\mu_{i}(S\circ S^{*})}\,,\qquad\qquad i\in I,$
and $s_{i}(S):=0$ if $I:=\{1,\dots,n\}$ and $i>n$. Recall that for compact, positive, and self-adjoint $R:H\to H$ we have
$s_{i}(R)=\sqrt{\mu_{i}(R\circ R^{*})}=\sqrt{\mu_{i}(R^{2})}=\mu_{i}(R)\,$, (14)
where the last identity directly follows from the spectral series representation, see e.g. [24, Equation (2.21) in Chapter V.2.3] or [52, Theorem VI.3.2]. For compact $S:H_{1}\to H_{2}$ this implies
$s_{i}^{2}(S)=\mu_{i}(S\circ S^{*})=s_{i}(S\circ S^{*})\,$, (15)
and in addition, there exist ONSs $(e_{i})_{i\in I}$ in $H_{1}$ and $(f_{i})_{i\in I}$ in $H_{2}$ with
$Sx=\sum_{i\in I}s_{i}(S)\langle x,e_{i}\rangle_{H_{1}}f_{i}\,,\qquad\qquad x\in H_{1},$ (16)
see e.g. [24, Equation (2.23) in Chapter V.2.3] or [52, Theorem VI.3.6]. Finally, recall that a compact $S:H_{1}\to H_{2}$ is nuclear if and only if we have
$\sum_{i\in I}s_{i}(S)<\infty\,$, (17)
see e.g. [5, Chapter 11.2]. Moreover, by definition $S$ is a Hilbert-Schmidt operator if $\sum_{j\in J}\|Se_{j}\|_{H_{2}}^{2}<\infty$ for some (and then all) ONB(s) $(e_{j})_{j\in J}$ of $H_{1}$, or equivalently, if and only if
$\sum_{i\in I}s_{i}^{2}(S)<\infty\,$, (18)
see e.g. [5, Chapter 11.3, Theorem 1]. By (15) we then immediately obtain the following elementary yet useful lemma.
Lemma 6.1.
Let $S:H_{1}\to H_{2}$ be a compact operator between two Hilbert spaces $H_{1}$ and $H_{2}$. Then $S$ is a Hilbert-Schmidt operator if and only if $S\circ S^{*}$ is nuclear.
Finally, it is sometimes helpful to compare the singular numbers with some other $s$-number scales in the sense of [36, Chapter 2]. To this end, let us first recall that for a given bounded linear operator $S:E\to F$ between two Banach spaces $E$ and $F$, the $n$-th approximation number is defined by
$a_{n}(S):=\inf\bigl\{\|S-A\|\,\bigl|\,A:E\to F\mbox{ bounded and linear with }\operatorname{rank}A<n\bigr\}\,.$ (19)
Furthermore, the $n$-th Kolmogorov number is defined by
$d_{n}(S):=\inf\bigl\{\varepsilon>0:\exists F_{\varepsilon}\mbox{ subspace of $F$ with $\dim F_{\varepsilon}<n$ and }SB_{E}\subset F_{\varepsilon}+\varepsilon B_{F}\bigr\}\,$, (20)
where, $B_{E}$ and $B_{F}$ denote the closed unit balls in $E$ and $F$ and we followed the exposition in [7, Chapter 2.2]. It is well known that the approximation and Kolmogorov numbers form an $s$-scale in the sense of [36, Chapter 2.2.1]. We refer to [36, Chapters 2.3 and 2.5] and [7, Chapters 2.1 and 2.2] for corresponding results. In addition, the Kolmogorov numbers are a surjective $s$-scale, see e.g. [36, Chapter 2.5]. Moreover, $S:E\to F$ is compact if and only if $d_{n}(S)\to 0$, see e.g. [7, page 49]. Furthermore, for operators $S:H_{1}\to H_{2}$ between Hilbert spaces we have $a_{n}(S)=d_{n}(S)$, see e.g. [35, Theorem 11.3.4], and if $S:H_{1}\to H_{2}$ is compact we even have
$s_{n}(S)=a_{n}(S)=d_{n}(S)\,,\qquad\qquad n\geq 1,$ (21)
by [35, Proposition 11.3.3] as (16) is a so-called Schmidt representation in the sense of [35, D.3.2]. In other words, we may freely switch between these quantities as soon as we know that $S$ is compact.
6.2 Nuclear Dominance
In this subsection we show that the notion of nuclear dominance used in [28] is equivalent to ours. To this end, we begin with recalling [28, Theorem 1.1].
Theorem 6.2.
Let $H_{1}$ and $H_{2}$ be RKHSs over $T$ with $H_{1}\subset H_{2}$. Then the resulting embedding map
$J:H_{1}$ $\to H_{2}$ (22)
$h_{1}$ $\mapsto h_{1}$
is linear and bounded. Moreover, there exists a unique linear $D_{H_{1},H_{2}}:H_{2}\to H_{2}$ that satisfies both $\operatorname{ran}D_{H_{1},H_{2}}\subset H_{1}$ and
$\langle h_{2},h_{1}\rangle_{H_{2}}=\langle D_{H_{1},H_{2}}h_{2},h_{1}\rangle_{H_{1}}\,,\qquad\qquad h_{1}\in H_{1},h_{2}\in H_{2}.$ (23)
In addition, $D_{H_{1},H_{2}}$ is bounded, positive and self-adjoint.
Recall that [28] speaks of nuclear dominance if $D_{H_{1},H_{2}}$ is nuclear. The following simple lemma shows that this notion is equivalent to our Definition 1.1.
Lemma 6.3.
Let $H_{1}$ and $H_{2}$ be RKHSs over $T$ with $H_{1}\subset H_{2}$ and $J:H_{1}\to H_{2}$ be the corresponding embedding map (22). Then we have
$D_{H_{1},H_{2}}=J\circ J^{*}\,.$
In particular, $D_{H_{1},H_{2}}$ is nuclear if and only if $J$ is a Hilbert-Schmidt operator. If, in this case, $(\mu_{i})_{i\in I}$ are the at most countably many non-zero eigenvalues (including geometric multiplicities) of $D_{H_{1},H_{2}}$ and $(f_{i})_{i\in I}\subset H_{2}$ is an ONS of corresponding eigenvectors, then we have
$H_{2}=\ker D_{H_{1},H_{2}}\oplus\overline{\operatorname{span}\{f_{i}:i\in I\}}^{H_{2}}=\ker D_{H_{1},H_{2}}\oplus\overline{\operatorname{ran}J}^{H_{2}}\,.$ (24)
Moreover, the vectors $e_{i}:=\mu_{i}^{-1}J^{*}f_{i}\in H_{1}$ satisfy $Je_{i}=f_{i}$ and $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ is an ONS in $H_{1}$.
Proof of Lemma 6.3: Let us write $R:=J\circ J^{*}:H_{2}\to H_{2}$. By construction, its range satisfies $\operatorname{ran}R\subset H_{1}$. Moreover, for $h_{1}\in H_{1}$ and $h_{2}\in H_{2}$ we have
$\langle h_{2},h_{1}\rangle_{H_{2}}=\langle h_{2},Jh_{1}\rangle_{H_{2}}=\langle J^{*}h_{2},h_{1}\rangle_{H_{1}}=\langle(J\circ J^{*})h_{2},h_{1}\rangle_{H_{1}}=\langle Rh_{2},h_{1}\rangle_{H_{1}}\,.$
Consequently, $R$ satisfies (23), and by uniqueness stated in Theorem 6.2 we can thus conclude $R=D_{H_{1},H_{2}}$. The equivalence is a direct consequence of Lemma 6.1 and $D_{H_{1},H_{2}}=J\circ J^{*}$.
Moreover, the first identity in (24) holds as noted below (13). To establish the second identity we first show
$\ker J^{*}=\ker D_{H_{1},H_{2}}\,.$ (25)
Here the inclusion $\subset$” follows from $D_{H_{1},H_{2}}=J\circ J^{*}$, and the converse inclusion follows from
$\|J^{*}h_{2}\|_{H_{1}}=\langle J^{*}h_{2},J^{*}h_{2}\rangle_{H_{1}}=\langle D_{H_{1},H_{2}}h_{2},h_{2}\rangle_{H_{2}}=0\,,\qquad\qquad h_{2}\in\ker D_{H_{1},H_{2}}.$
Now, combining the first identity in (24) with (25) yields
$\overline{\operatorname{span}\{f_{i}:i\in I\}}^{H_{2}}=(\ker D_{H_{1},H_{2}})^{\perp}=(\ker J^{*})^{\perp}=\overline{\operatorname{ran}J}^{H_{2}}\,$, (26)
where the last equation is a textbook formula, see e.g. [8, Theorem II.2.19 with Corollary I.2.10].
Finally, the definition of the $e_{i}$ immediately yields
$Je_{i}=\mu_{i}^{-1}(J\circ J^{*})f_{i}=\mu_{i}^{-1}D_{H_{1},H_{2}}f_{i}=f_{i}\,$,
as well as
$\langle e_{i},e_{j}\rangle_{H_{1}}=\mu_{i}^{-1}\mu_{j}^{-1}\langle J^{*}f_{i},J^{*}f_{j}\rangle_{H_{1}}=\mu_{i}^{-1}\mu_{j}^{-1}\langle f_{i},D_{H_{1},H_{2}}f_{j}\rangle_{H_{2}}=\mu_{i}^{-1}\langle f_{i},f_{j}\rangle_{H_{2}}\,.$
With the help of this identity we easily see that $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ is an ONS in $H_{1}$. ∎
6.3 Choosing Small RKHSs
In this subsection, we show how we can “shrink” the RKHS containing the paths of a suitable version of a Gaussian process in the situation of Theorem 1.2.
To this end, let us recall that a map $\xi:\Omega\to H$ from a measurable space $(\Omega,{\cal A})$ into a Hilbert space $H$ is weakly measurable, if $\langle\xi,f\rangle_{H}:\Omega\to\mathbb{R}$ is measurable for all $f\in H$. Moreover, it is called strongly measurable if it is weakly measurable and $\xi(\Omega)$ is separable in $H$. In this case, the function $\omega\to\|\xi(\omega)\|_{H}$ is also measurable.
Now, our first result shows that we can always assume that the space $H$ is separable as long as we are willing to potentially switch the versions of $X$.
Lemma 6.4.
Let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ with RKHS $H_{X}$. Moreover, let $H$ be an RKHS on $T$ with kernel $k$ and $Y$ be a version of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
Then there exists a separable RKHS $H^{\dagger}\subset H$ and another version $Y^{\dagger}$ of $X$ with $Y_{\bullet}^{\dagger}(\omega)\in H^{\dagger}$ for all $\omega\in\Omega$ such that the resulting map $\Omega\to H^{\dagger}$ given by $\omega\mapsto Y_{\bullet}^{\dagger}(\omega)$ is strongly measurable and satisfies
$\mathbb{E}_{\mathrm{P}}|\langle Y_{\bullet}^{\dagger},f\rangle_{H^{\dagger}}|^{2}=\langle D_{H_{X},H^{\dagger}}f,f\rangle_{H}\,,\qquad\qquad f\in H^{\dagger}.$
Proof of Lemma 6.4: Theorem 1.2 gives $H_{X}\ll H$ and a version $\tilde{Y}$ of $X$ with $\mathrm{P}(\{\tilde{Y}_{\bullet}\in H\})=1$. Obviously, we may assume without loss of generality that this version actually satisfies $\tilde{Y}_{\bullet}(\omega)\in H$ for all $\omega\in\Omega$.
Let us now define $\xi:\Omega\to H$ by $\xi(\omega):=\tilde{Y}_{\bullet}(\omega)$ for all $\omega\in\Omega$. Then [28, Lemma 2.1] shows that $\xi$ is weakly measurable and the reproducing property of $k$ yields
$\tilde{Y}_{t}(\omega)=\langle\xi(\omega),k(t,\,\cdot\,)\rangle_{H}\,,\qquad\qquad\omega\in\Omega,t\in T.$ (27)
Moreover, since $\tilde{Y}$ is a version of $X$ we have $H_{\tilde{Y}}=H_{X}\ll H$. Consequently [28, Corollary 3.1] gives a strongly measurable $\xi^{\dagger}:\Omega\to H$ with $\|\xi^{\dagger}\|_{H}\in{{\cal L}_{2}(\mathrm{P})}$ and
$\mathrm{P}\bigl(\bigl\{\omega\in\Omega:\langle\xi^{\dagger}(\omega),f\rangle_{H}=\langle\xi(\omega),f\rangle_{H}\bigr\}\bigr)=1\,,\qquad\qquad f\in H.$ (28)
Alternatively, one can obtain $\xi^{\dagger}$ by [14, Theorem 5.1], see also [11, page 89] for a more detailed proof, as $\xi$ is Pettis integrable and Hilbert spaces have the Radon-Nikodym property.
We now define $H^{\dagger}:=\overline{\xi^{\dagger}(\Omega)}^{H}$. Since $\xi^{\dagger}$ is strongly measurable, $H^{\dagger}$ is separable. In addition, as a closed subspace of an RKHS it is an RKHS. Let us now consider process $Y^{\dagger}:=(Y^{\dagger}_{t})_{t\in T}$ defined by
$Y^{\dagger}_{t}(\omega):=\langle\xi^{\dagger}(\omega),k(t,\,\cdot\,)\rangle_{H}\,,\qquad\qquad t\in T,\omega\in\Omega.$
Then our construction gives $Y_{\bullet}^{\dagger}(\omega)=\xi^{\dagger}(\omega)\in H^{\dagger}$ for all $\omega\in\Omega$. Moreover, combining this definition with (28) and (27) we see that $Y^{\dagger}$ is a version of $\tilde{Y}$, and thus also a version of $X$.
Now, since $Y^{\dagger}$ is a version of $X$ with $\mathrm{P}(\{Y_{\bullet}^{\dagger}\in H^{\dagger}\})=1$ another application of Theorem 1.2 yields $H_{Y^{\dagger}}=H_{X}\ll H^{\dagger}$. Applying [28, Corollary 3.1], see also [28, Equation (3.2)], to $Y^{\dagger}$ then gives the last assertion. ∎
The next result, which is the main theorem of this subsection, refines the “shrinking” to a somewhat minimal space.
Theorem 6.5.
Let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ with RKHS $H_{X}$. Moreover, let $H$ be an RKHS on $T$ and $Y$ be a version of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
Then there exist a version $Y^{*}$ of $X$ and a separable RKHS $H^{*}$ on $T$ such that $H_{X}\subset H^{*}$ is dense in $H^{*}$ and
$\mathrm{P}(\{Y_{\bullet}^{*}\in H^{*}\})=1\,.$
Proof of Theorem 6.5: Without loss of generality we may assume that $Y=Y^{\dagger}$ and $H=H^{\dagger}$, where $Y^{\dagger}$ and $H^{\dagger}$ are according to Lemma 6.4. In particular, we may assume that $H$ is separable. In the following we consider $Y^{*}:=Y$, so that it remains to find $H^{*}$.
Since Theorem 1.2 ensures $H_{X}\ll H$, Lemma 6.3 gives us an at most countable ONS $(f_{i})_{i\in I}\subset H$ consisting of eigenvectors of $D_{H_{X},H}$ with non-zero eigenvalues $(\mu_{i})_{i\in I}$ such that
$H=\ker D_{H_{X},H}\oplus\overline{\operatorname{span}\{f_{i}:i\in I\}}^{H}\,.$ (29)
We define $H^{*}:=\overline{\operatorname{span}\{f_{i}:i\in I\}}^{H}$ and equip it with the norm of $H$. Then $H^{*}$ is a separable RKHS and the second identity of (24) in Lemma 6.3 shows
$H^{*}=\overline{\operatorname{span}\{f_{i}:i\in I\}}^{H}=\overline{\operatorname{ran}J}^{H}=\overline{H_{X}}^{H}\,$,
where the third identity is true since $J:H_{X}\to H$ is the embedding map. In particular, we have $H_{X}\subset H^{*}$, and for every $h\in H^{*}$ there exists a sequence $(f_{n})\subset H_{X}$ with $\|f_{n}-h\|_{H}\to 0$. Since we know $f_{n}\in H^{*}$ we conclude that $\|f_{n}-h\|_{H^{*}}\to 0$. In other words, we have shown $H^{*}\subset\overline{H_{X}}^{H^{*}}$.
Now, since $H$ is separable, so is its subspace $\ker D_{H_{X},H}$. Consequently, there exists a countable dense subset $\mathfrak{D}\subset\ker D_{H_{X},H}$. For $f\in\mathfrak{D}$ the last assertion of Lemma 6.4 yields
$\mathbb{E}_{\mathrm{P}}|\langle Y_{\bullet},f\rangle_{H}|^{2}=\langle D_{H_{X},H}f,f\rangle_{H}=0\,$,
and hence there exists an $N_{f}\in{\cal A}$ with $\mathrm{P}(N_{f})=0$ and $\langle Y_{\bullet}(\omega),f\rangle_{H}=0$ for all $\omega\in\Omega\setminus N_{f}$. Let us define
$N:=\bigcup_{f\in\mathfrak{D}}N_{f}\,.$
Clearly, we have $N\in{\cal A}$ with $\mathrm{P}(N)=0$ and $\langle Y_{\bullet}(\omega),f\rangle_{H}=0$ for all $f\in\mathfrak{D}$ and $\omega\in\Omega\setminus N$.
With these preparations we can now show $Y_{\bullet}(\omega)\in H^{*}$ for all $\omega\in\Omega\setminus N$. To this end, we pick an $f\in\ker D_{H_{X},H}$. Then there exists a sequence $(f_{n})\subset\mathfrak{D}$ with $f_{n}\to f$ in $H$. For $\omega\in\Omega\setminus N$ our construction then shows
$\langle Y_{\bullet}(\omega),f\rangle_{H}=\lim_{n\to\infty}\langle Y_{\bullet}(\omega),f_{n}\rangle_{H}=0\,.$
In other words, we have found $Y_{\bullet}(\omega)\in(\ker D_{H_{X},H})^{\perp}=H^{*}$, where the last identity follows from (29) and the definition of $H^{*}$. ∎
6.4 Proof of Theorem 3.2 and its Corollaries
Before we can prove Theorem 3.2 and its corollaries we need to establish a few auxiliary results. We begin with an elementary lemma describing the effect of multiplying a kernel by some function.
Lemma 6.6.
Let $k$ be a kernel on $T$ with RKHS $H$ and $\alpha:T\to\mathbb{R}$ be a function. Then the RKHS of the kernel $\alpha\star k:T\times T\to\mathbb{R}$ defined by
$(\alpha\star k)(t,t^{\prime}):=\alpha(t)\alpha(t^{\prime})k(t,t^{\prime})\,,\qquad\qquad t,t^{\prime}\in X,$
is $H_{\alpha\star k}:=\{\alpha h:h\in H\}$ with norm given by $\|f\|_{H_{\alpha\star k}}=\inf\bigl\{\|h\|_{H}:h\in H\mbox{ with }f=\alpha h\bigr\}$.
Proof of Lemma 6.6: Let $\Phi:T\to H$ be the canonical feature map of $k$, that is $\Phi(t)=k(\,\cdot\,,t)$ for all $t\in T$. For $t_{1},t_{2}\in T$ an elementary calculation then shows
$\bigl\langle\alpha(t_{1})\Phi(t_{1}),\alpha(t_{2})\Phi(t_{2})\bigr\rangle_{H}=\alpha(t_{1})\alpha(t_{2})\langle\Phi(t_{1}),\Phi(t_{2})\rangle_{H}=\alpha(t_{1})\alpha(t_{2})k(t_{1},t_{2})\,.$
Consequently, $\alpha\Phi:T\to H$ is a feature map of $\alpha\star k$. Moreover, for $h\in H$ and $t\in T$ another simple calculation shows
$\bigl\langle h,(\alpha\Phi)(t)\bigr\rangle_{H}=\alpha(t)\,\langle h,\Phi(t)\rangle_{H}=\alpha(t)h(t)\,.$
Now the assertion follows from the fundamental theorem of RKHSs, see e.g. [46, Theorem 4.21]. ∎
The next auxiliary results describes the relationship between envelope functions and the behavior of the kernel on its diagonal.
Lemma 6.7.
Let $H$ be an RKHS on $T$ with kernel $k$. Then $b:T\to[0,\infty)$ is an envelope for $H$, if and only if
$\sqrt{k(t,t)}\leq b(t)\,,\qquad\qquad t\in T.$ (30)
Proof of Lemma 6.7: If (30) is true, then the assertion immediately follows from (4).
To establish the converse implication, we define $\alpha:T\to[0,\infty)$ by
$\alpha(t):=\begin{cases}1/b(t),&\mbox{ if }b(t)>0\,,\\ 0,&\mbox{ otherwise.}\end{cases}$
Let us now fix an $f\in H_{\alpha\star k}$, where $H_{\alpha\star k}$ is the RKHS considered in Lemma 6.6. According to Lemma 6.6 there then exists an $h\in H$ with $f=\alpha h$.
Our next, intermediate goal is to show that there is only one such $h$. To this end, we note that for $t\in T$ with $b(t)>0$ we obviously have $h(t)=b(t)f(t)$ by the definition of $\alpha$. Moreover, for $t\in T$ with $b(t)=0$ the assumed envelope condition (3) ensures $h(t)=0$. In other words, the function $h$ is uniquely determined for all $t\in T$, and therefore Lemma 6.6 shows
$\|f\|_{H_{\alpha\star k}}=\|h\|_{H}\,.$
With these preparations we now observe that, for all $t\in T$, the envelope condition gives
$|f(t)|=|\alpha(t)|\cdot|h(t)|\leq|\alpha(t)|\cdot\|h\|_{H}\cdot b(t)\leq\|h\|_{H}=\|f\|_{H_{\alpha\star k}}\,.$
Since $f\in H_{\alpha\star k}$ was arbitrary, we have thus found $H_{\alpha\star k}\hookrightarrow\ell_{\infty}(T)$ with $\|\operatorname{id}:H_{\alpha\star k}\to\ell_{\infty}(T)\|\leq 1$. By standard RKHS theory, see e.g. [46, Lemma 4.23], we then obtain
$\alpha(t)\alpha(t)k(t,t)=(\alpha\star k)(t,t)\leq 1\,,\qquad\qquad t\in T.$
For $t\in T$ with $b(t)>0$ this obviously implies (30). Moreover, for $t\in T$ with $b(t)=0$ the envelope condition applied to $h:=k(\,\cdot\,,t)\in H$ gives
$|h(t^{\prime})|\leq\|h\|_{H}\cdot b(t^{\prime})\,,\qquad\qquad t^{\prime}\in T,$
and for $t^{\prime}:=t$ the latter implies $k(t,t)=h(t)=0$. Consequently, (30) is also true for $t\in T$ with $b(t)=0$. ∎
Our next and final auxiliary result investigates the operator $I_{k,\nu}:H\to{L_{2}(\nu)}$ in more detail.
Lemma 6.8.
[…]
$I_{k,\nu}:H$ $\to{L_{2}(\nu)}$ (31)
$h$ $\mapsto[h]_{\sim}$
is linear and Hilbert-Schmidt. Moreover, its adjoint $I^{*}_{k,\nu}:{L_{2}(\nu)}\to H$ is given by
$(I_{k,\nu}^{*}f)(s)=\int_{T}k(s,t)f(t)\,\mathrm{d}\nu(t)\,,\qquad\qquad f\in{L_{2}(\nu)},\,s\in T.$
Proof of Lemma 6.8: We first note that the envelope condition (3) gives $|h(t)|^{2}\leq\|h\|_{H}^{2}\cdot b^{2}(t)$ for all $h\in H$ and $t\in T$. By the assumed $b\in{{\cal L}_{2}(\nu)}$ we thus find $H\subset{{\cal L}_{2}(\nu)}$. This in turn shows that $I_{k,\nu}$ is well-defined. Moreover, the linearity of $I_{k,\nu}$ is obvious.
To verify that $I_{k,\nu}$ is Hilbert-Schmidt, we fix an ONB $(e_{j})_{j\in J}$ of $H$. For finite $A\subset J$ we then find
$\sum_{j\in A}\|I_{k,\nu}e_{j}\|_{{L_{2}(\nu)}}^{2}=\int_{T}\sum_{j\in A}e^{2}_{j}(t)\,\mathrm{d}\nu(t)=\int_{T}\sum_{j\in A}\bigl|\langle e_{j},k(\,\cdot\,,t)\rangle_{H}\bigr|^{2}\,\mathrm{d}\nu(t)\,.$
Moreover, Bessel’s inequality gives
$\sum_{j\in A}\bigl|\langle e_{j},k(\,\cdot\,,t)\rangle_{H}\bigr|^{2}\leq\|k(\,\cdot\,,t)\|_{H}^{2}=k(t,t)\leq b^{2}(t)\,$,
where in the last step we used Lemma 6.7. Combining this estimate with our initial identity then yields
$\sum_{j\in A}\|I_{k,\nu}e_{j}\|_{{L_{2}(\nu)}}^{2}=\int_{T}\sum_{j\in A}\bigl|\langle e_{j},k(\,\cdot\,,t)\rangle_{H}\bigr|^{2}\,\mathrm{d}\nu(t)\leq\|b\|_{{{\cal L}_{2}(\nu)}}^{2}\,.$
Since this is true for all finite $A\subset J$ it is also true for $A:=J$, that is, $I_{k,\nu}$ is Hilbert-Schmidt.
Let us finally derive the formula for the adjoint operator. To this end, we fix an $f\in{L_{2}(\nu)}$ and an $s\in T$. We then obtain
$(I_{k,\nu}^{*}f)(s)=\bigl\langle I_{k,\nu}^{*}f,k(s,\,\cdot\,)\bigr\rangle_{H}=\bigl\langle f,I_{k,\nu}(k(s,\,\cdot\,))\bigr\rangle_{{L_{2}(\nu)}}=\bigl\langle f,[k(s,\,\cdot\,)]_{\sim}\bigr\rangle_{{L_{2}(\nu)}}=\int_{T}k(s,t)f(t)\,\mathrm{d}\nu(t)\,$,
where in the first step we used the reproducing property $k$. ∎
Proof of Theorem 3.2: By Theorem 6.5 we may assume without loss of generality that $H_{X}\subset H$ is dense in $H$. Since convergence in $H$ implies pointwise convergence and we assumed $H_{X}\subset{{\cal L}_{0}({\cal B})}$, we therefore also find $H\subset{{\cal L}_{0}({\cal B})}$. In turn, $b\in{{\cal L}_{2}(\nu)}$ in combination with Lemma 6.8 ensures that $I_{k,\nu}$ is a Hilbert-Schmidt operator, where $k$ is the kernel of $H$ and $I_{k,\nu}$ is defined by (31).
Moreover, Theorem 1.2 ensures $H_{X}\ll H$, that is the embedding map $J:H_{X}\to H$ is a Hilbert-Schmidt operator. In summary, both factors $J$ and $I_{k,\nu}$ in the natural factorization
[…]
are Hilbert-Schmidt, and therefore their product $I_{X,\nu}$ is nuclear, see e.g. [5, Chapter 11.3, Theorem 3]. ∎
Proof of Corollary 3.3: Since $H\subset\ell_{\infty}(T)$ we may consider the constant envelope function $b$ given by $b(t)=\|\operatorname{id}:H\to\ell_{\infty}(T)\|$ for all $t\in T$. Then all finite measures $\nu$ on $(T,{\cal B})$ satisfy $b\in{{\cal L}_{2}(\nu)}$, and therefore the assertion follows from Theorem 3.2. ∎
Proof of Corollary 3.4: As in the proof of Theorem 3.2 we may assume without loss of generality that $H\subset{{\cal L}_{0}({\cal B})}$.
Now, let $k$ be the kernel of $H$ and $k_{|T^{*}}:=T^{*}\times T^{*}\to\mathbb{R}$ be its restriction onto $T^{*}$, that is, $k_{|T^{*}}(s,t)=k(s,t)$ for all $s,t\in T^{*}$. It is well-known, see e.g. [4, Theorem 6 on page 25], that the RKHS of $k_{|T^{*}}$ is $H_{|T^{*}}:=\{h_{|T^{*}}:h\in H\}$ with norm given by
$\|f\|_{H_{|T^{*}}}=\inf\bigl\{\|h\|_{H}:h\in H\mbox{ with }h_{|T^{*}}=f\bigr\}$
for all $f\in H_{|T^{*}}$. Note that $H\subset{{\cal L}_{0}({\cal B})}$ immediately gives $H_{|T^{*}}\subset{{\cal L}_{0}({\cal B}_{|T^{*}})}$. In fact, even the inclusion $H_{|T^{*}}\subset{{\cal L}_{\infty}({\cal B}_{|T^{*}})}$ is true, since for $f\in H_{|T^{*}}$ and $h\in H$ with $h_{|T^{*}}=f$ we have
$\|f\|_{\infty}=\sup_{t\in T^{*}}|f(t)|=\sup_{t\in T^{*}}|h(t)|\leq\|h\|_{H}\cdot\sup_{t\in T^{*}}b(t)\leq\|h\|_{H}\cdot M<\infty\,.$
Since $\mathrm{P}(\{Y_{\bullet}\in H\})>0$ implies $H_{X}\subset H$ by Theorem 1.2, and this in turn guarantees $H_{X_{|T^{*}}}\subset H_{|T^{*}}$, we then also find the claimed inclusion $H_{X_{|T^{*}}}\subset{{\cal L}_{\infty}({\cal B}_{|T^{*}})}$.
If we now write $Y^{*}:=(Y_{t})_{t\in T^{*}}$ for the restriction of $Y$ onto $T^{*}$, then $Y^{*}$ is a version of the centered Gaussian process $X_{|T^{*}}$. Moreover, restricting the paths of $Y$ onto $T^{*}$ gives the paths of $Y^{*}$, that is,
$Y_{\bullet}^{*}(\omega)=\bigl(Y_{\bullet}(\omega)\bigr)_{|T^{*}}\,,\qquad\qquad\omega\in\Omega,$
and therefore we find $\mathrm{P}(\{Y_{\bullet}^{*}\in H_{|T^{*}}\})\geq\mathrm{P}(\{Y_{\bullet}\in H\})>0$. Applying Corollary 3.3 to $X_{|T^{*}}$, $H_{|T^{*}}$, and the version $Y^{*}=(Y_{t})_{t\in T^{*}}$ then gives the assertion. ∎
6.5 Proofs Related to Essentially Equivalent RKHSs
In this subsection we prove Lemma 3.6 as well as Theorems 3.7 and 3.8.
To begin with, let us assume that $H_{1}$ and $H_{2}$ are RKHSs on $T$ with kernels $k_{1}$ and $k_{2}$. Then their sum $H_{1}+H_{2}$ can be equipped with the natural norm
$\|h\|_{H_{1}+H_{2}}^{2}=\inf\bigl\{\|h_{1}\|_{H_{1}}^{2}+\|h_{2}\|_{H_{2}}^{2}:h_{1}\in H_{1},h_{2}\in H_{2}\mbox{ with }h=h_{1}+h_{2}\bigr\}\,.$ (32)
It turns out that $H_{1}+H_{2}$ equipped with this norm is the RKHS of the kernel $k_{1}+k_{2}$, see e.g. [3, Theorem on page 353]. Consequently, (5) implies $H_{1}+H_{1}^{\star}\cong H_{2}+H_{2}^{\star}$ as noted at the beginning of Section 2. With these preparations, the proof of Lemma 3.6 is straightforward:
Proof of Lemma 3.6: By our assumption we have $k_{1}+k_{1}^{\star}=k_{2}+k_{2}^{\star}$. As discussed above this gives $H_{1}+H_{1}^{\star}=H_{2}+H_{2}^{\star}$, and hence also $H_{1}+H_{1}^{\star}\circeq H_{2}+H_{2}^{\star}$. ∎
In the following, we adopt the usual notation $H_{1}\oplus H_{2}:\circeq H_{1}+H_{2}$ if we additionally have $H_{1}\cap H_{2}=\{0\}$. Moreover, if we equip $H_{1}\oplus H_{2}$ with the norm given by (32), then the resulting RKHS is denoted by $H_{1}\oplus_{2}H_{2}$. In this case, taking the infimum in (32) is, of course, superfluous, and $H_{1}$ and $H_{2}$ are closed subspaces of $H_{1}\oplus_{2}H_{2}$ with $H_{1}^{\perp}=H_{2}$ and $H_{2}^{\perp}=H_{1}$.
Now, our first auxiliary result shows that sums of RKHSs are suitable direct sums if one of the involved spaces is finite dimensional.
Lemma 6.9.
Let $H$ and $H_{0}$ be RKHSs on $T$ with $\dim H_{0}<\infty$. Then there exists an RKHS $H^{\star}$ on $T$ with $H^{\star}\subset H_{0}$ and
$H\oplus H^{\star}\circeq H+H_{0}\,.$ (33)
Proof of Lemma 6.9: Since $H\cap H_{0}$ is a subspace of $H_{0}$ and $\dim H_{0}<\infty$, there exists a subspace $H^{\star}\subset H_{0}$ with
$H_{0}\circeq(H\cap H_{0})\oplus H^{\star}\,.$ (34)
Moreover, $\dim H_{0}<\infty$ shows that $H^{\star}$ is closed in $H_{0}$ and thus $H^{\star}$ is an RKHS.
We first show that $H\cap H^{\star}=\{0\}$. This this end, we pick an $h\in H\cap H^{\star}$. By $H^{\star}\subset H_{0}$ we find $h\in H\cap H_{0}$ and hence we have $h\in(H\cap H_{0})\cap H^{\star}$. By (34) we then obtain $h=0$.
For the verification of (33) we first note that $H^{\star}\subset H_{0}$ implies $H\oplus H^{\star}\subset H+H_{0}$. For the converse inclusion, we fix some $h\in H$ and $h_{0}\in H_{0}$. By (34) there then exist a $g\in H\cap H_{0}$ and an $h^{\star}\in H^{\star}$ with $h_{0}=g+h^{\star}$. This gives
$h+h_{0}=h+h_{0}-h^{\star}+h^{\star}\in H+H^{\star}\circeq H\oplus H^{\star}\,$,
where in the second step we used $h_{0}-h^{\star}=g\in H\cap H_{0}\subset H$. ∎
The next auxiliary lemma shows that we may use direct sums in the definition of essentially equivalent RKHSs and that we have some additional control over the spaces $H_{1}^{\star}$ and $H_{2}^{\star}$.
Lemma 6.10.
Let $H_{1}$ and $H_{2}$ be essentially equivalent RKHSs on $T$. If $E$ is a vector space consisting of functions $T\to\mathbb{R}$ with $H_{1}\cup H_{2}\subset E$, then there exist finite dimensional RKHSs $H_{1}^{\star}$ and $H_{2}^{\star}$ on $T$ with $H_{1}^{\star}\cup H_{2}^{\star}\subset E$ and
$H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}\,.$
Proof of Lemma 6.10: By our assumption and Lemma 6.9 there exist finite dimensional RKHSs $H_{1}^{\times}$ and $H_{2}^{\times}$ on $T$ with
$H_{1}\oplus H_{1}^{\times}\circeq H_{2}\oplus H_{2}^{\times}\,.$ (35)
We equip both direct sums with the $\oplus_{2}$-norm and consider the two orthogonal projections $P_{i}:H_{i}\oplus_{2}H_{i}^{\times}\to H_{i}\oplus_{2}H_{i}^{\times}$ onto $H_{i}$ as well as the two orthogonal projections $Q_{i}:H_{i}\oplus_{2}H_{i}^{\times}\to H_{i}\oplus_{2}H_{i}^{\times}$ onto $H_{i}^{\times}$. Moreover, we define $H:\circeq H_{1}\oplus H_{1}^{\times}\circeq H_{2}\oplus H_{2}^{\times}$ and the subspace
$H^{\dagger}:\circeq\bigl\{(Q_{1}-Q_{2})h:h\in H\bigr\}\circeq\bigl\{(Q_{2}-Q_{1})h:h\in H\bigr\}\,.$
Clearly, $H^{\dagger}$ is a subspace of $H_{1}^{\times}+H_{2}^{\times}$ and thus a finite dimensional RKHS on $T$. Moreover, for $h\in H$ our construction yields
$P_{1}h+Q_{1}h=h=P_{2}h+Q_{2}h\,.$ (36)
In particular, for $h_{1}\in H_{1}$ we find $h_{1}=P_{1}h_{1}=P_{2}h_{1}+(Q_{2}-Q_{1})h_{1}\in H_{2}+H^{\dagger}$, and hence $H_{1}\subset H_{2}+H^{\dagger}$. The latter implies $H_{1}+H^{\dagger}\subset H_{2}+H^{\dagger}$. Since analogously we find $H_{2}+H^{\dagger}\subset H_{1}+H^{\dagger}$ by the symmetry of $H^{\dagger}$ in $Q_{1}$ and $Q_{2}$, we have
$H_{1}+H^{\dagger}\circeq H_{2}+H^{\dagger}\,.$
Moreover, (36) also shows $(Q_{1}-Q_{2})h=P_{2}h-P_{1}h\in H_{2}+H_{1}\subset E$, and hence we obtain $H^{\dagger}\subset E$. Applying Lemma 6.9 on both sides then yields the assertions. ∎
Lemma 6.11.
Let $H_{1},H_{2}$ and $H_{1}^{\star},H_{2}^{\star}$ be RKHSs on $T$ with $H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}$ and $m_{i}:=\dim H_{i}^{\star}<\infty$. For $i=1,2$ we write $\overline{H}_{i}:=H_{i}\oplus_{2}H_{i}^{\star}$ and consider the embedding maps $J_{i}:H_{i}\to\overline{H}_{i}$. Finally, let $E$ be a Banach space and $S_{i}:\overline{H}_{i}\to E$ be bounded and linear operators with $S_{1}h=S_{2}h$ for all $h\in\overline{H}_{1}\circeq\overline{H}_{2}$. Then there exists a constant $c\in(0,\infty)$ with
$d_{n+m_{2}}(S_{1}\circ J_{1})$ $\leq c\,d_{n}(S_{2}\circ J_{2})\,$,
$d_{n+m_{1}}(S_{2}\circ J_{2})$ $\leq c\,d_{n}(S_{1}\circ J_{1})$
for all $n\geq 1$.
Proof of Lemma 6.11: We first note that we have $\overline{H}_{1}\cong\overline{H}_{2}$ as mentioned at the beginning of Section 2, and hence there exists a constant $c\in(0,\infty)$ with
$c^{-1}\,d_{n}(S_{1})\leq d_{n}(S_{2})\leq c\,d_{n}(S_{1})\,,\qquad\qquad n\geq 1.$ (37)
Let us further write $P_{i}:\overline{H}_{i}\to\overline{H}_{i}$ for the orthogonal projection onto $H_{i}$ and $P^{\perp}_{i}:\overline{H}_{i}\to\overline{H}_{i}$ for the orthogonal projection onto its complement $H_{i}^{\perp}=H_{i}^{\star}$, that is $P^{\perp}_{i}=\operatorname{id}_{\overline{H}_{i}}-P_{i}$, where $\operatorname{id}_{\overline{H}_{i}}$ is the identity map on $\overline{H}_{i}$. Moreover, we define $Q_{i}:\overline{H}_{i}\to H_{i}$ by $Q_{i}h:=P_{i}h$ for all $h\in\overline{H}_{i}$. In other words, the only difference between $P_{i}$ and $Q_{i}$ are the spaces the operators formally map into. Now, by construction the operators $Q_{i}$ are metric surjections and we have
$P_{i}=J_{i}\circ Q_{i}\qquad\qquad\mbox{ and }\qquad\qquad Q_{i}\circ J_{i}=\operatorname{id}_{H_{i}}\,.$
By the additivity of the Kolmogorov numbers we then find for all $n\geq 1$
$d_{n+m_{i}}(S_{i})\leq d_{m_{i}+1}(S_{i}-S_{i}\circ P_{i})+d_{n}(S_{i}\circ P_{i})$ $=d_{m_{i}+1}(S_{i}\circ P^{\perp}_{i})+d_{n}(S_{i}\circ J_{i}\circ Q_{i})$
$\leq\|S_{i}\|\cdot d_{m_{i}+1}(P^{\perp}_{i})+d_{n}(S_{i}\circ J_{i}\circ Q_{i})$
$=d_{n}(S_{i}\circ J_{i})\,$,
where in the last step we used $\operatorname{rank}P_{i}^{\perp}=m_{i}<m_{i}+1$ and the surjectivity of the Kolmogorov numbers. Combining this inequality with $\|J_{i}\|=1$ and (37) yields
$d_{n+m_{2}}(S_{1}\circ J_{1})\leq d_{n+m_{2}}(S_{1})\leq c\,d_{n+m_{2}}(S_{2})\leq c\,d_{n}(S_{2}\circ J_{2})\,$,
and the second inequality can be shown analogously. ∎
Proof of Theorem 3.7: Since $H_{1}$ and $H_{2}$ are essentially equivalent with $H_{1},H_{2}\subset\color[rgb]{0,0,0}{{\cal L}_{2}(\nu)}$ by Lemma 6.8, Lemma 6.10 shows that there are finite dimensional RKHSs $H_{1}^{\star}\subset\color[rgb]{0,0,0}{{\cal L}_{2}(\nu)}\color[rgb]{0,0,0}$ and $H_{2}^{\star}\subset\color[rgb]{0,0,0}{{\cal L}_{2}(\nu)}\color[rgb]{0,0,0}$ on $T$ with
$H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}\,.$
We write $m_{i}:=\dim H_{i}^{\star}$ and observe that $\overline{H}_{i}:=H_{i}\oplus_{2}H_{i}^{\star}\subset\color[rgb]{0,0,0}{{\cal L}_{2}(\nu)}$ for $i=1,2$. Consequently, we can consider the “embedding” operators $S_{i}:=I_{\overline{H}_{i},\nu}:\overline{H}_{i}\to{L_{2}(\nu)}$ given by $h\mapsto[h]_{\sim}$. Clearly this shows $S_{1}h=S_{2}h$ for all $h\in\overline{H}_{1}\circeq\overline{H}_{2}$ and for the embedding maps $J_{i}:H_{i}\to\overline{H}_{i}$ we have
$S_{i}\circ J_{i}=I_{H_{i},\nu}\,.$
Before be can apply Lemma 6.11 we finally need to verify that both $S_{1}$ and $S_{2}$ are continuous: To this end, we note that both operators $I_{H_{i},\nu}:H_{i}\to{L_{2}(\nu)}$ are continuous by our assumptions and Lemma 6.8, and the operators $I_{H_{i}^{\star},\nu}:H_{i}^{\star}\to{L_{2}(\nu)}$ are continuous because of $\dim H_{i}^{\star}<\infty$. For $h\in\overline{H}_{i}:=H_{i}\oplus_{2}H_{i}^{\star}$ and its unique decomposition $h=h_{i}+h_{i}^{\star}$ with $h_{i}\in H_{i}$ and $h_{i}^{\star}\in H_{i}^{\star}$ we then find
$\|S_{i}h\|_{{L_{2}(\nu)}}\leq\bigl\|\,[h_{i}]_{\sim}\,\bigr\|_{{L_{2}(\nu)}}+\bigl\|\,[h_{i}^{\star}]_{\sim}\,\bigr\|_{{L_{2}(\nu)}}$ $\leq\|I_{H_{i},\nu}\|\cdot\|h_{i}\|_{H_{i}}+\|I_{H_{i}^{\star},\nu}\|\cdot\|h_{i}^{\star}\|_{H_{i}^{\star}}$
$\leq\sqrt{2}\cdot\max\{\|I_{H_{i},\nu}\|,\|I_{H_{i}^{\star},\nu}\|\}\cdot\bigl(\|h_{i}\|_{H_{i}}^{2}+\|h_{i}^{\star}\|_{H_{i}^{\star}}^{2}\bigr)^{1/2}$
$=\sqrt{2}\cdot\max\{\|I_{H_{i},\nu}\|,\|I_{H_{i}^{\star},\nu}\|\}\cdot\|h\|_{\overline{H}_{i}}\,$,
that is, the operators $S_{i}:\overline{H}_{i}\to{L_{2}(\nu)}$ are indeed continuous. Applying Lemma 6.11 thus gives a constant $c\in(0,\infty)$ with
$d_{n+m_{2}}(I_{H_{1},\nu})\leq c\,d_{n}(I_{H_{2},\nu})\qquad\qquad\mbox{ and }\qquad\qquad d_{n+m_{1}}(I_{H_{2},\nu})\leq c\,d_{n}(I_{H_{1},\nu})$ (38)
for all $n\geq 1$. Let us now assume that $I_{H_{1},\nu}$ is nuclear. Then $I_{H_{1},\nu}$ is compact and (21) together with the discussion around it shows both $s_{n}(I_{H_{1},\nu})=d_{n}(I_{H_{1},\nu})$ and $d_{n}(I_{H_{1},\nu})\to 0$. By (38) we conclude that $d_{n}(I_{H_{2},\nu})\to 0$ and hence $I_{H_{2},\nu}$ is also compact as discussed around (21). Consequently, we have $s_{n}(I_{H_{2},\nu})=d_{n}(I_{H_{2},\nu})$, and therefore the second inequality in (38) shows
$s_{n+m_{1}}(I_{H_{2},\nu})\leq c\,s_{n}(I_{H_{1},\nu})$
for all $n\geq 1$. Since we assumed that $I_{H_{1},\nu}$ is nuclear, we then see that $I_{H_{2},\nu}$ is also nuclear. The converse implication can be shown analogously. ∎
Proposition 6.12.
Let $H_{1},H_{2}$ be RKHS on $T$ and $H_{1}^{\star},H_{2}^{\star}$ be finite dimensional RKHSs on $T$ with $H_{1}\oplus H_{1}^{\star}\circeq H_{2}\oplus H_{2}^{\star}$. Moreover, let $H$ be an RKHS on $T$ with $H_{1}\ll H$. Then there exists an RKHS $H_{1}^{\dagger}\subset H_{1}^{\star}$ with
$H_{2}\ll H\oplus_{2}H_{1}^{\dagger}\,.$
Proof of Proposition 6.12: For $i=1,2$ we consider the embedding maps $J_{i}:H_{i}\to H_{i}\oplus_{2}H_{i}^{\star}$. Now, we first note that by Lemma 6.9 there exists an $H_{1}^{\dagger}\subset H_{1}^{\star}$ with
$H\oplus H_{1}^{\dagger}\circeq H+H_{1}^{\star}\,.$
Moreover, by our assumption $H_{1}\ll H$ we have $H_{1}\subset H$ and the resulting embedding $J:H_{1}\to H$ is Hilbert-Schmidt. Since $H_{1}\subset H$ gives
$H_{2}\oplus H_{2}^{\star}\circeq H_{1}\oplus H_{1}^{\star}\subset H+H_{1}^{\star}\circeq H\oplus H_{1}^{\dagger}\,$,
we can thus consider the two resulting embedding maps $S_{i}:H_{i}\oplus_{2}H_{i}^{\star}\to H\oplus_{2}H_{1}^{\dagger}$. Clearly, this gives $S_{i}h=S_{2}h$ for all $h\in H_{1}\oplus_{2}H_{1}^{\star}\circeq H_{2}\oplus_{2}H_{2}^{\star}$. Moreover, the operators $S_{i}$ are continuous since embedding maps between RKHSs are automatically continuous by a simple application of the closed graph theorem, see e.g. [42, Lemma 23]. By Lemma 6.11 there then exists a constant $c\in(0,\infty)$ such that for $m_{1}:=\dim H_{1}^{\dagger}$ we have
$d_{n+m_{1}}(S_{2}\circ J_{2})$ $\leq c\,d_{n}(S_{1}\circ J_{1})\,,\qquad\qquad n\geq 1.$
Let us now consider the embedding map $I:H\to H\oplus_{2}H_{1}^{\dagger}$, which by construction satisfies $\|I\|\leq 1$. Clearly, our construction gives
$H_{1}$$_{1}\oplus_{2}H_{1}^{\star}HH$$H\oplus_{2}H_{1}^{\dagger}$$_{1}JJ$$S_{1}$II
and hence we obtain
$d_{n+m_{1}}(S_{2}\circ J_{2})\leq c\,d_{n}(S_{1}\circ J_{1})=c\,d_{n}(I\circ J)\leq c\,d_{n}(J)$
for all $n\geq 1$. Now, $J$ is Hilbert-Schmidt, and hence compact. This shows both $d_{n}(J)\to 0$ and $d_{n}(J)=s_{n}(J)$ as discussed around (21). Consequently, we find $d_{n}(S_{2}\circ J_{2})\to 0$ and thus also $s_{n}(S_{2}\circ J_{2})=d_{n}(S_{2}\circ J_{2})$, where for the second assertion we again used the properties mentioned around (21). In summary, we thus have $s_{n+m_{1}}(S_{2}\circ J_{2})\leq c\,s_{n}(J)$, and since $J$ is Hilbert-Schmidt, so ist $S_{2}\circ J_{2}$. However, $S_{2}\circ J_{2}:H_{2}\to H\oplus_{2}H_{1}^{\dagger}$ is the embedding map, and hence we have found $H_{2}\ll H\oplus_{2}H_{1}^{\dagger}$. ∎
Proof of Theorem 3.8: Let us consider the RKHS $E:=H^{\dagger}+H_{X}$. By our assumptions we have $H_{X}\cup H\subset H_{X}\cup H^{\dagger}\subset E$, and therefore Lemma 6.10 provides us with finite dimensional RKHSs $H^{\star}$ and $H_{X}^{\star}$ on $T$ with $H^{\star}\cup H_{X}^{\star}\subset E$ and
$H\oplus H^{\star}\circeq H_{X}\oplus H_{X}^{\star}\,.$
Combining this with the assumed $H\ll H^{\dagger}$, Proposition 6.12 gives us an RKHS $H^{\ddagger}\subset H^{\star}$ with
$H_{X}\ll H^{\dagger}\oplus_{2}H^{\ddagger}\,.$
In addition, note that $H_{X}^{\dagger}:=H^{\dagger}\oplus_{2}H^{\ddagger}$ is essentially equivalent to $H^{\dagger}$ and our construction ensures $H_{X}^{\dagger}=H^{\dagger}\oplus_{2}H^{\ddagger}\subset H^{\dagger}+H^{\star}\subset H^{\dagger}+H^{\dagger}+H_{X}=H^{\dagger}+H_{X}$, where for the second inclusion we used $H^{*}\subset E=H^{\dagger}+H_{X}$. ∎
6.6 Proof of Theorem 3.9
Proof of Theorem 3.9: i). We first recall that Lemma 6.8 ensures the compactness of $I_{X,\nu}$ as since $b\in{{\cal L}_{2}(\nu)}$ is an envelope function for $H_{X}$. The definition of singular numbers then yields
$s_{i}(I_{X,\nu})=\sqrt{\mu_{i}(I_{X,\nu}\circ I_{X,\nu}^{*})}=\sqrt{\mu_{i}(T_{X,\nu})}\,.$
By (17) we thus obtain the equivalence.
ii). Let $J:H_{X}\to H_{X}^{\beta}$ be the embedding map. For $i\in I$ we then have
$\|J(\sqrt{\mu_{i}}e_{i})\|_{H_{X}^{\beta}}=\mu_{i}^{1/2-\beta/2}\|\mu_{i}^{\beta/2}e_{i}\|_{H_{X}^{\beta}}=\mu_{i}^{1/2-\beta/2}\,$,
and this in turn shows
$\sum_{i\in I}\|J(\sqrt{\mu_{i}}e_{i})\|_{H_{X}^{\beta}}^{2}=\sum_{i\in I}\mu_{i}^{1-\beta}\,.$
Since the $H_{X}$-positivity of $\nu$ ensures that $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ is an ONB of $H_{X}$, see [47, Theorem 3.1], we can then conclude that (10) indeed characterizes the Hilbert-Schmidt property of $J$, see the discussion around (18). The second assertion directly follows from $H_{X}\ll H_{X}^{\beta}$ and Theorem 1.2. ∎
6.7 Proofs Related to RKHSs of Sobolev Type
Before we can prove Theorem 4.1 some preparations are required. We begin by providing details for (11). To this end, let $B^{s}_{p,q}(\mathbb{R}^{d})$ and $F^{s}_{p,q}(\mathbb{R}^{d})$ denote the Besov spaces, respectively Triebel-Lizorkin spaces, see e.g. [1, 15]. Then for $s>d/2$ we have
$H^{s}(\mathbb{R}^{d})\cong F_{2,2}^{s}(\mathbb{R}^{d})\cong B_{2,2}^{s}(\mathbb{R}^{d})\hookrightarrow C_{b}(\mathbb{R}^{d})\,$, (39)
where the first identity follows from e.g. [1, Chapter 7, Paragraphs 7.62, 7.65, and 7.57], the second identity is shown in e.g. [1, Chapter 7, Paragraph 7.67], and the inclusion can be found in e.g. [1, Theorem 7.34].
In addition, Triebel-Lizorkin and Besov spaces on open Euclidean balls $T\subset\mathbb{R}^{d}$ are defined analogously to (12), that is $A^{s}_{p,q}(T):=\{f_{|T}:f\in A^{s}_{p,q}(\mathbb{R}^{d})\}$ with norm
$\|h\|_{A^{s}_{p,q}(T)}:=\inf\bigl\{\|f\|_{A^{s}_{p,q}(\mathbb{R}^{d})}:f\in A^{s}_{p,q}(\mathbb{R}^{d})\mbox{ with }f_{|T}=h\bigr\}\,$
where $A^{s}_{p,q}$ stands for either $F^{s}_{p,q}$ or $B^{s}_{p,q}$, see e.g. [15, Chapter 2.5]. Note that if $F_{p,q}^{s}(\mathbb{R}^{d})$ or $B_{p,q}^{s}(\mathbb{R}^{d})$ are BSFs then the same is true for $F_{p,q}^{s}(T)$ or $B_{p,q}^{s}(T)$. In addition, continuous inclusions between such spaces on $\mathbb{R}^{d}$ carry over to the restricted spaces on $T$, see e.g. [15, Chapter 2.5] or [42, Proposition 3] for a general BSF statement. For $s>d/2$, Equation (39) thus implies
$H^{s}(T)\cong F_{2,2}^{s}(T)\cong B_{2,2}^{s}(T)\hookrightarrow C_{b}(T)\,.$ (40)
For the proof of Theorem 4.1 we need the following auxiliary result that investigates both the operators $I_{H^{s}(T),\lambda^{d}}:H^{s}(T)\to{L_{2}(T)}$ and the nuclear dominance between fractional Sobolev spaces.
Lemma 6.13.
Let $T\subset\mathbb{R}^{d}$ be bounded and measurable with non-empty interior and $s>d/2$. Then the following statements hold true:
  1. i)
    The operator $I_{H^{s}(T),\lambda^{d}}:H^{s}(T)\to{L_{2}(T)}$ is nuclear if and only if $s>d$.
  2. ii)
    If $s>d$ and $r>d/2$, then $H^{s}(T)\ll H^{r}(T)$ if and only if r<sd/2r<s-d/2.
Proof of Lemma 6.13: i). We first consider the case, when $T$ is an open Euclidean ball. Then we have $B_{2,2}^{0}(\mathbb{R}^{d})\cong F_{2,2}^{0}(\mathbb{R}^{d})\cong{L_{2}(\mathbb{R}^{d})}$, where the first identity follows from e.g. [15, (1) in Chapter 2.3.3] and the second identity can be found in e.g. [50, (5) in Chapter 2.5.6]. Restricting to $T$ leads to $B_{2,2}^{0}(T)\cong{L_{2}(T)}$, and therefore [15, (10) in Chapter 2.5.1] shows that the embedding $B_{2,2}^{s}(T)\hookrightarrow{L_{2}(T)}$ is compact. By (40) we can thus conclude that $I_{H^{s}(T),\lambda^{d}}:H^{s}(T)\to{L_{2}(T)}$ is also compact, and by (21) we hence find
$a_{n}(I_{H^{s}(T),\lambda^{d}})=d_{n}(I_{H^{s}(T),\lambda^{d}})=s_{n}(I_{H^{s}(T),\lambda^{d}})\,,\qquad\qquad n\geq 1.$ (41)
Moreover, the already observed $H^{s}(T)\cong B_{2,2}^{s}(T)$ and $B_{2,2}^{0}(T)\cong{L_{2}(T)}$ in combination with the results in [15, Chapter 3.3.4] show that
$a_{n}(I_{H^{s}(T),\lambda^{d}})\sim n^{-s/d}\,.$ (42)
Combining (41) with (42) then yields the characterization.
To prepare the proof of the general case, let us assume that we have measurable $T_{1}\subset T_{2}\subset\mathbb{R}^{d}$ with $\lambda^{d}(T_{1}),\lambda^{d}(T_{2})\in(0,\infty)$. Moreover, let $Q:H^{s}(T_{2})\to H^{s}(T_{1})$ and $R:{L_{2}(T_{2})}\to{L_{2}(T_{1})}$ be the restriction operators. Then $Q$ and $R$ are metric surjections and we have the commutative diagram
[…]
By the surjectivity of the Kolmogorov numbers we thus find
$d_{n}(I_{H^{s}(T_{1}),\lambda^{d}})=d_{n}(I_{H^{s}(T_{1}),\lambda^{d}}\circ Q)=d_{n}(R\circ I_{H^{s}(T_{2}),\lambda^{d}})\leq d_{n}(I_{H^{s}(T_{2}),\lambda^{d}})\,,\qquad\qquad n\geq 1.$ (43)
With these preparations we can now establish the assertion in the general case. To this end, we first note that by our assumptions on $T$ there exist open balls $B_{0}\subset T\subset B_{1}$. Now, we have already seen at the very beginning that $I_{H^{s}(B_{i}),\lambda^{d}}$ is compact, and hence we have $d_{n}(I_{H^{s}(B_{i}),\lambda^{d}})\to 0$ for $i=0,1$. Applying (43) we further find $d_{n}(I_{H^{s}(T),\lambda^{d}})\to 0$, and therefore $I_{H^{s}(T),\lambda^{d}}$ is also compact. By (21) we conclude that
$s_{n}(I_{H^{s}(T),\lambda^{d}})=d_{n}(I_{H^{s}(T),\lambda^{d}})\,,\qquad\qquad n\geq 1,$
while our initial considerations (41) with (42) showed
$d_{n}(I_{H^{s}(B_{i}),\lambda^{d}})=a_{n}(I_{H^{s}(B_{i}),\lambda^{d}})\sim n^{-s/d}\,,\qquad\qquad n\geq 1$
for $i=0,1$. Combining these identities with (43) yields $s_{n}(I_{H^{s}(T),\lambda^{d}})\sim n^{-s/d}$.
ii). Again, we first consider the case, when $T$ is an open ball. Here, we note that $H^{s}(T)\cong B_{2,2}^{s}(T)$ and $H^{r}(T)\cong B_{2,2}^{r}(T)$ in combination with the results in [15, Chapter 3.3.4] show that
$a_{n}(J)\sim n^{-(s-r)/d}\,$,
where $J:H^{s}(T)\to H^{r}(T)$ denotes the embedding map. By the results mentioned around (21) the same is true for the Kolmogorov numbers and in particular we have $d_{n}(J)\to 0$. Therefore, $J$ is compact and (21) shows that
$s_{n}(J)\sim n^{-(s-r)/d}\,.$
Consequently, we have $\sum_{n\geq 1}s_{n}^{2}(J)<\infty$ if and only if r<sd/2r<s-d/2.
To prepare the proof of the general case, we again assume that we have measurable $T_{1}\subset T_{2}\subset\mathbb{R}^{d}$ with $\lambda^{d}(T_{1}),\lambda^{d}(T_{2})\in(0,\infty)$. Moreover, let $Q_{s}:H^{s}(T_{2})\to H^{s}(T_{1})$ and $Q_{r}:H^{r}(T_{2})\to H^{r}(T_{1})$ be the restriction operators, which by construction of the involved spaces are metric surjections. In addition, note that $H^{s}(\mathbb{R}^{d})\subset H^{r}(\mathbb{R}^{d})$ ensures that the embeddings $J_{i}:H^{s}(T_{i})\to H^{r}(T_{i})$ are defined for $i=1,2$. Clearly, these maps enjoy the following commutative diagram
[…]
By the surjectivity of the Kolmogorov numbers we thus find
$d_{n}(J_{1})=d_{n}(J_{1}\circ Q_{s})=d_{n}(Q_{r}\circ J_{2})\leq d_{n}(J_{2})\,,\qquad\qquad n\geq 1.$
With these preparations, the rest of the proof follows the arguments used in the proof of the general case of i). ∎
In addition, the proof of Theorem 4.1 as well as some of the examples in Section 5 need the following result that refines the analysis of Theorem 6.5.
Lemma 6.14.
Let $T\subset\mathbb{R}^{d}$ be bounded and measurable with non-empty interior. Moreover, let $(\Omega,{\cal A},\mathrm{P})$ be a complete probability space and $X:=(X_{t})_{t\in T}$ be a centered Gaussian process over $(\Omega,{\cal A},\mathrm{P})$ whose RKHS $H_{X}$ consists of continuous functions. Finally, let $H$ be an RKHS $H$ on $T$ and $Y$ be a version of $X$ with
$\mathrm{P}(\{Y_{\bullet}\in H\})>0\,.$
Then there exist a version $Y^{*}$ of $X$, a separable RKHS $H^{*}$ on $T$ with kernel $k^{*}$, and an $M>0$ such that $H_{X}\subset H^{*}$ is dense in $H^{*}$, the set $T^{*}:=\{t\in T:k(t,t)\leq M\}$ is measurable and has non-empty interior, and
$\mathrm{P}(\{Y_{\bullet}^{*}\in H^{*}\})=1\,.$
Proof of Lemma 6.14: By Theorem 6.5 we may assume without loss of generality that $H$ is separable and $H_{X}\subset H$ is dense in $H$. In the following we thus show the assertions for $H$ and $Y$.
Under these assumptions, the embedding $J:H_{X}\to H$ satisfies
$\overline{\operatorname{ran}J}^{H}=\overline{H_{X}}^{H}=H\,$,
and therefore Lemma 6.3 gives us an at most countable family $(e_{i})\subset H_{X}$ such that the functions $f_{i}:=Je_{i}=e_{i}$ form an ONB $(f_{i})_{i\in I}$ of $H$. By e.g. [46, Theorem 4.20] we can thus conclude that the kernel $k$ of $H$ is given by
$k(s,t)=\sum_{i\in I}f_{i}(s)f_{i}(t)\,,\qquad\qquad s,t\in T.$
Now, since $H_{X}$ consists of continuous functions, all $f_{i}$ are continuous, and therefore the function $b:T\to[0,\infty)$ defined by b(t):=k(t,t)b(t):={k(t,t)} is both measurable and lower semi-continuous.
We now fix an open and non-empty $O\subset T$ and write $O_{M}:=\{t\in O:b_{|O}(t)>M\}$. Obviously, this yields
$\emptyset=\bigcap_{M\in\mathbb{N}}O_{M}$ (44)
and all sets $O_{M}$ are relatively open in $O$ by the lower semi-continuity of $b$. Moreover, (44) shows that their countable intersection is not dense in $O$. In addition, since $\mathbb{R}^{d}$ is a complete metric space, there exists a complete metric on $O$ describing the topology on $O$, see e.g. [13, Theorem 2.5.4]. Consequently, Baire’s theorem, see e.g. [29, Theorem 1.5.4], shows that there is an $M\geq 1$ such that $O_{M}$ is not dense in $O$. This in turn means that its complement $O\setminus O_{M}=\{t\in O:b_{|O}(t)\leq M\}$ has non-empty relative interior in $O$. Consequently, there exists an relatively $O$-open set $O^{*}\neq\emptyset$ with
$O^{*}\subset\{t\in O:b_{|O}(t)\leq M\}\subset\{t\in T:b(t)\leq M\}\,.$
Since $O$ open in $\mathbb{R}^{d}$, so is $O^{*}$. In other words, the set $T^{*}$ has non-empty interior. ∎
Proof of Theorem 4.1: ii) $\Rightarrow$ iii). Trivial.
iii) $\Rightarrow$ i). By Lemma 6.14 we may assume without loss of generality that there is an $M>0$ such that $T^{*}:=\{t\in T:k(t,t)\leq M\}$ is measurable and has non-empty interior, where $k$ is the kernel of $H$. Consequently, the Lebesgue measure on $T^{*}$ is non-trivial and finite. By Corollary 3.4 we then know that
$I_{X_{|T^{*}},\lambda^{d}}:H_{X_{|T^{*}}}\to{L_{2}(T^{*})}$
is nuclear. Moreover, since $H_{X}$ is essentially equivalent to $H^{s}(T)$, the spaces $H_{X_{|T^{*}}}$ and $H^{s}(T^{*})$ are also essentially equivalent. In addition, we know $H_{X_{|T^{*}}}\subset{{\cal L}_{\infty}(T^{*})}$ by Corollary 3.4, and since we also have $H^{s}(T^{*})\subset{{\cal L}_{\infty}(T^{*})}$ both spaces have a joint constant envelope. Consequently, Theorem 3.7 shows that $I_{H^{s}(T^{*}),\lambda^{d}}:H^{s}(T^{*})\to{L_{2}(T^{*})}$ is also nuclear. Now the assertion follows by Lemma 6.13.
i) $\Rightarrow$ ii). Let us fix an d/2<r<sd/2d/2<r<s-d/2. Lemma 6.13 then shows $H^{s}(T)\ll H^{r}(T)$. Since $H_{X}$ is essentially equivalent to $H^{s}(T)$ by assumption, Theorem 3.8 then provides an RKHS $H_{X}^{\dagger}$ on $T$ that is essentially equivalent to $H^{r}(T)$ and satisfies $H_{X}\ll H_{X}^{\dagger}$. Now the assertion follows from Theorem 1.2 for $H:=H_{X}^{\dagger}$. ∎
Proof of Theorem 4.2: Using (40) we first note that it suffices to consider the case $H_{0}=B^{s}_{2,2}(0,1)$. In the following we thus assume that $\hat{k}_{0}$ is the kernel of $\hat{H}_{0}:=B_{2,2}^{s}(\mathbb{R})$ and that $k_{0}$ is its restriction to $(0,1)$.
Now let $S_{2,2}^{s}B(\mathbb{R}^{d})$ denote the Besov space of dominating mixed smoothness $s$, see [43, Appendix A.2] and the references mentioned therein. Then [43, Theorem 2.2] shows
$S_{2,2}^{s}B(\mathbb{R}^{d})\cong B^{s}_{2,2}(\mathbb{R})\otimes_{\alpha_{2}}\dots\otimes_{\alpha_{2}}B^{s}_{2,2}(\mathbb{R})\,$, (45)
where $\alpha_{2}$ denotes the $2$-nuclear tensor norm, see [43, Definition B.2]. Moreover, the $\alpha_{2}$-tensor norm equals the $g_{2}:=\alpha_{2,1}$-tensor norm in [10, Chapters 12.5 and 12.7], and $g_{2}$ in turn equals the natural Hilbert space tensor norm, see [10, Chapters 26.6 and 26.7], where the latter is called Hilbert-Schmidt norm. Denoting the natural Hilbert space tensor product by $\otimes_{2}$, the Identity (45) thus reads as
$S_{2,2}^{s}B(\mathbb{R}^{d})\cong B^{s}_{2,2}(\mathbb{R})\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(\mathbb{R})\,.$ (46)
Now the tensor product space on the right hand side can be interpreted as the RKHS of the product kernel
$\prod_{i=1}^{d}\hat{k}_{0}(t_{i},t_{i}^{\prime})\,,\qquad\qquad t,t^{\prime}\in\mathbb{R}^{d},$
see e.g. [4, Chapter 1.4.6] or [46, Lemma 4.6]. Restricting both sides of (46) to $T$ thus gives
$S_{2,2}^{s}B(T)\cong B^{s}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(0,1)=H_{X}\,$, (47)
where the space $S_{2,2}^{s}B(T)$ is defined by restricting the functions of $S_{2,2}^{s}B(\mathbb{R}^{d})$ to $T$ as in [31, Definition 3.8 and Remark 3.9]. Moreover, the embedding $I:S_{2,2}^{s}B(T)\to{L_{2}(T)}$ is defined and compact and its approximation numbers satisfy
$a_{n}(I)\sim n^{-s}(\ln n)^{(d-1)s}$ (48)
see [40] as well as [31, Theorem 2.11] for a presentation of the results of [40] in the notation we use in this proof.
ii) $\Rightarrow$ i). By Corollary 3.3 we know that $I_{X,\lambda^{d}}:H_{X}\to{L_{2}(T)}$ is nuclear. By combining (47), (48), and (21) we thus find $s>1$.
i) $\Rightarrow$ ii). We pick an $r\in(1/2,s-1/2)$ and an ONB $(e_{i})$ of $B^{s}_{2,2}(0,1)$. By Lemma 6.13 and (40) we then know that $B^{s}_{2,2}(0,1)\ll B^{r}_{2,2}(0,1)$, and hence we have
$\sum_{i=1}^{\infty}\|e_{i}\|_{B^{r}_{2,2}(0,1)}^{2}<\infty$
as discussed around (18). Now recall that the collection of elementary tensors $e_{i_{1}}\otimes\dots\otimes e_{i_{d}}$ forms an ONB of $B^{s}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(0,1)=H_{X}$. Moreover, we have
$\sum_{i_{1}=1}^{\infty}\dots\sum_{i_{d}=1}^{\infty}\|e_{i_{1}}\otimes\dots\otimes e_{i_{d}}\|_{B^{r}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{r}_{2,2}(0,1)}^{2}$ $=\sum_{i_{1}=1}^{\infty}\dots\sum_{i_{d}=1}^{\infty}\prod_{j=1}^{d}\|e_{i_{j}}\|_{B^{r}_{2,2}(0,1)}^{2}$
$=\biggl(\sum_{i=1}^{\infty}\|e_{i}\|_{B^{r}_{2,2}(0,1)}^{2}\biggr)^{d}$
$<\infty\,$,
which shows that the embedding $B^{s}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{s}_{2,2}(0,1)\hookrightarrow B^{r}_{2,2}(0,1)\otimes_{2}\dots\otimes_{2}B^{r}_{2,2}(0,1)$ is Hilbert-Schmidt. Now ii) follows by Theorem 1.2. ∎
6.8 Proofs for the Examples
Proof of Example 5.1: It is well-known, see e.g. [2, Chapter III.2, page 69], [4, Example 1, page 71], or [27, page 28], that the RKHS $H_{W}$ of the Wiener process is given
$H_{W}\circeq\{h\in H^{1}([0,1]):h=0\}\,.$ (49)
Therefore, $H_{W}$ is essentially equivalent to $H^{1}([0,1])$ and the non-existence of $H$ follows from Theorem 4.1.
For the proof of the second assertion we note that $I_{H^{1}([0,1],\lambda}:H^{1}([0,1]\to{L_{2}(\lambda)}$ is not nuclear by Lemma 6.13. Moreover, we obviously have $H_{W}\subset H^{1}([0,1])\subset{{\cal L}_{\infty}(T)}$, and therefore Theorem 3.7 shows that $I_{W,\lambda}:H_{W}\to{L_{2}(\lambda)}$ is not nuclear, either. ∎
Proof of Example 5.2: Let us recall that the covariance function of the Brownian bridge $X:=(X_{t})_{t\in T}$ is
$k_{X}(t_{1},t_{2})=\min\{t_{1},t_{2}\}-t_{1}t_{2}=k_{W}(t_{1},t_{2})-t_{1}t_{2}\,,\qquad\qquad t_{1},t_{2}\in[0,1],$
where $k_{W}$ is the covariance function of the Wiener process. Now, $(t_{1},t_{2})\mapsto t_{1}t_{2}$ is a kernel on $[0,1]$ whose RKHS is one dimensional, and consequently, $H_{B}$ and $H_{W}$ are essentially equivalent by Lemma 3.6. Moreover, we know by (49) that $H_{W}$ is essentially equivalent to $H^{1}([0,1])$, and therefore $H_{B}$ is also essentially equivalent to $H^{1}([0,1])$. Now the assertion follows from Theorem 4.1 since $H_{B}$ consists of continuous functions. ∎
Proof of Example 5.3: We first consider the covariance function $k_{X}^{(1)}$. For the corresponding RKHS $H_{X}^{(1)}$ we then find $H_{X}^{(1)}\cong H^{1}(T)$ by e.g. [4, Example 17 on p. 316] applied for some $a,b\in\mathbb{R}$ with $T\subset(a,b)$ followed by a restriction onto $T$. By Theorem 4.1 we then obtain the assertion.
In the case of the second covariance, we first observe that $(t_{1},t_{2})\mapsto a\exp(-\sigma(t_{1}+t_{2}))$ gives a kernel $k^{*}$ on $[0,1]$ whose RKHS is one dimensional. Moreover, we obviously have $k_{X}^{(2)}=k_{X}^{(1)}-k^{*}$, and therefore the RKHS $H_{X}^{(2)}$ of $k_{X}^{(2)}$ is essentially equivalent to $H_{X}^{(1)}\cong H^{1}(T)$ by Lemma 3.6. Now we again obtain the assertion by Theorem 4.1. ∎
Proof of Example 5.4: In the following we write $H_{\alpha}:=H_{B^{(\alpha)}}$ for the RKHS of $B^{(\alpha)}$. Then [34, Theorem 6.12] shows that a function $h:T\to\mathbb{R}$ is in $H_{\alpha}$ if and only if there are sequences $(a_{n})_{n\geq 0}$ and $(b_{n})_{n\geq 1}$ with
$\sum_{n\geq 0}n^{2\alpha+1}a_{n}^{2}<\infty\qquad\mbox{ and }\qquad\sum_{n\geq 1}n^{2\alpha+1}b_{n}^{2}<\infty$
such that for all $t\in[0,1]$ we have
$h(t)=h(1)t+\sum_{n\geq 0}a_{n}\cos(2\pi nt)+\sum_{n\geq 1}a_{n}\sin(2\pi nt)\,.$
With these preparations we can now turn to the actual proof.
Here, we first note that in the case $\alpha=1/2$ we have $B^{(\alpha)}=W$, and hence there is nothing left to prove by Example 5.1.
Moreover, in the case $0<\alpha<1/2$ the above description of $H_{\alpha}$ yields $H_{1/2}\subset H_{\alpha}$ and the corresponding embedding map $J:H_{1/2}\to H_{\alpha}$ is continuous by a simple application of the closed graph theorem, see [42, Lemma 23]. Now assume that there was an RKHS $H$ on $T$ with $\mathrm{P}(\{B_{\bullet}^{(\alpha)}\in H\})=1$. By Lemma 6.14 we may then assume without loss of generality that there is an $M>0$ such that $T^{*}:=\{t\in T:k(t,t)\leq M\}$ is measurable and has non-empty interval, where $k$ is the kernel of $H$. Corollary 3.4 then shows that the operator $I_{(H_{\alpha})_{|T^{*}},\lambda}:(H_{\alpha})_{|T^{*}}\to{L_{2}(T^{*})}$ is nuclear. The natural factorization
[…]
then shows that $I_{(H_{1/2})_{{\color[rgb]{0,0,0}|T^{*}}},\lambda}$ would also be nuclear. However, we have seen in (49) that $H_{1/2}=H_{W}$ is essentially equivalent to $H^{1}([0,1])$ and therefore $(H_{1/2})_{{\color[rgb]{0,0,0}|T^{*}}}$ is essentially equivalent to $H^{1}(T^{*})$. Now, Lemma 6.13 shows that $I_{H^{1}(T^{*}),\lambda}:H^{1}(T^{*})\to{L_{2}(T^{*})}$ is not nuclear, and since we have both $H^{1}(T^{*})\subset{{\cal L}_{\infty}(T^{*})}$ and $(H_{1/2})_{{\color[rgb]{0,0,0}|T^{*}}}\subset{{\cal L}_{\infty}(T^{*})}$, Theorem 3.7 shows that $I_{(H_{1/2})_{{\color[rgb]{0,0,0}|T^{*}}},\lambda}$ is not nuclear, either. In other words, we have found a contradiction.
Finally, in the case $1/2<\alpha<1$ it is well known that $\mathrm{P}$-almost all paths of $B^{(\alpha)}$ are $\beta$-Hölder continuous for all $\beta\in(0,\alpha)$, where we note that this can be quickly verified as the stationary increments and the self-similarity of $B^{(\alpha)}$ gives
$\mathbb{E}\bigl|B^{(\alpha)}_{t_{1}}-B^{(\alpha)}_{t_{2}}\big|^{p}=\mathbb{E}\bigl|B^{(\alpha)}_{|t_{1}-t_{2}|}\big|^{p}=|t_{1}-t_{2}|^{\alpha p}\cdot\mathbb{E}\bigl|B^{(\alpha)}_{1}\big|^{p}\,,\qquad\qquad t_{1},t_{2}\geq 0,p>0\,$,
and therefore an application of the Kolmogorov-Chentsov theorem, see e.g. [25, Theorem 21.6] for $p>0$ with $\beta<(\alpha p-1)/p$ gives the $\beta$-Hölder continuity. Let us now fix some $s\in(1/2,\alpha)$ and some $\beta\in(s,\alpha)$. By e.g. [50, (6) and (9) in Chapter 2.5.7 in combination with (1) and (3) in Chapter 2.2.2] we then have
$\operatorname{H\ddot{o}l}^{\beta}(\mathbb{R})\cong B^{\beta}_{\infty,\infty}(\mathbb{R})\,$,
while for open and bounded intervals $I$ we have
$B^{\beta}_{\infty,\infty}(I)\hookrightarrow B^{s}_{2,2}(I)\,$,
see e.g. [50, (7) in Chapter 3.3.1]. Taking such an $I$ with $T\subset I$, the combination of both results restricted to $T$ then leads to
$\operatorname{H\ddot{o}l}^{\beta}(T)\cong B^{\beta}_{\infty,\infty}(T)\hookrightarrow B^{s}_{2,2}(T)\cong H^{s}(T)\,$,
where the last identity was noted in (40). Since we have already seen that almost all paths are $\beta$-Hölder continuous, we thus obtain the assertion. ∎
Proof of Example 5.5: We first consider the case $\alpha\leq 1/2$. Here we recall from [34, Theorem 5.4] that the RKHS $H_{R^{(\alpha)}}$ of the Riemann-Liouville process and the RKHS $H_{B^{(\alpha)}}$ of the fractional Brownian motion are equal up to equivalent norms. Consequently, they are essentially equivalent and we have
$H_{B^{(1/2)}}\subset H_{B^{(\alpha)}}\circeq H_{R^{(\alpha)}}\,.$
Let $J:H_{B^{(1/2)}}\to H_{R^{(\alpha)}}$ be the corresponding embedding map, which is bounded by e.g. [42, Lemma 23]. As in the proof of Example 5.4, we now assume that there was an RKHS $H$ on $T$ with $\mathrm{P}(\{R_{\bullet}^{(\alpha)}\in H\})=1$. We then choose $T^{*}$ as in the proof of Example 5.4, so that Corollary 3.4 again ensures that
$I_{(H_{R^{(\alpha)}})_{|T^{*}},\lambda}:(H_{R^{(\alpha)}})_{|T^{*}}\to{L_{2}(T^{*})}$
is nuclear. Since $J$ is bounded, a diagram analogous to that in the proof of Example 5.4 then shows that $I_{(H_{1/2})_{{\color[rgb]{0,0,0}|T^{*}}},\lambda}$ is nuclear. However, we have already seen in the proof of Example 5.4 that is false.
Let us now consider the case $\alpha\in(1/2,1)$. Here, we need to recall that we can jointly realize the processes $R^{(\alpha)}$ and $B^{(\alpha)}$ such that the paths of $R^{(\alpha)}-B^{(\alpha)}$ are $C^{\infty}$, see e.g. [34, Theorem 5.1]. Since for $s\in(1/2,\alpha)$ all $C^{\infty}$-functions are contained in $B_{2,2}^{s}(T)$, see e.g. [1, Paragraph 7.32], the identity $B_{2,2}^{s}(T)\cong H^{s}(T)$, see (40), in combination with
$\mathrm{P}(\{B_{\bullet}^{(\alpha)}\in H^{s}(T)\})=1\,$,
which has been shown in Example 5.4, gives $\mathrm{P}(\{R_{\bullet}^{(\alpha)}\in H^{s}(T)\})=1$. ∎
Proof of Example 5.6: Since $T$ is bounded, there exists an open Euclidean ball $B$ with $T\subset B$. Moreover, the RKHS $H_{\alpha,\sigma}(B)$ of the Matérn kernel on $B$ satisfies $H_{\alpha,\sigma}(B)\cong B_{2,2}^{\alpha+d/2}(B)$, see e.g. [51, Corollary 10.13] in combination with e.g. [39, (4.15)], or alternatively [38, (27)]. In addition, we have $B_{2,2}^{\alpha+d/2}(B)\cong H^{\alpha+d/2}(B)$, see (40), and hence we find
$H_{\alpha,\sigma}(B)\cong H^{\alpha+d/2}(B)\,.$
Restricting both spaces onto $T$ leads to $H_{\alpha,\sigma}(T)\cong H^{\alpha+d/2}(T)$ by [42, Proposition 3]. By Theorem 4.1 and $H_{X}=H_{\alpha,\sigma}(T)$ we then obtain the assertion. ∎
Proof of Example 5.7: We first note that we have $k_{X}(t,t)=k_{X}(t-t,0)=k_{X}(0,0)$ for all $t\in T$. Consequently, $k_{X}$ is bounded, see e.g. [46, (4.14)]. In addition, for $\kappa(r):=k_{X}(r,0)$ we have $k_{X}(t_{1},t_{2})=k_{X}(t_{2},t_{1})=k_{X}(-t_{1}+t_{2},0)=\kappa(-t_{1}+t_{2})$, and hence $k_{X}$ is translation invariant. Finally, recall that $k_{X}$ is measurable by assumption. With the help of these properties of $k_{X}$ [48, Lemma 4.1 and 4.2] then show that $k_{X}$ is continuous and there exist an at most countable family $(e_{i}^{*})_{i\in I}$ of continuous functions $T\to\mathbb{R}$ with $\sup_{i\in I}\|e_{i}^{*}\|_{\infty}\leq\sqrt{2}$ and a summable family $(\mu_{i})_{i\in I}\subset[0,\infty)$ such that $([e_{i}^{*}]_{\sim})_{i\in I}$ is an ONB of $L_{2}(\lambda)$} and
$k_{X}(t_{1},t_{2})=\sum_{\mu_{i}>0}\mu_{i}e_{i}^{*}(t_{1})e_{i}^{*}(t_{2})\,,\qquad\qquad t_{1},t_{2}\in T,$ (50)
where the series converges absolutely for all $t_{1},t_{2}\in T$ as well as uniformly in $(t_{1},t_{2})$. Moreover, [48, Lemma 4.2] shows that the positive $\mu_{i}$ are the collection of non-zero eigenvalues of $T_{k,\lambda}$ including geometric multiplicities and the $e_{i}^{*}]_{\sim}$ are corresponding eigenfunctions. Finally, for the functions defined in (6) with $f_{i}:={e_{i}^{*}}$ and $\mu_{i}>0$ we have
$e_{i}=\mu_{i}^{-1}I_{X,\nu}^{*}[e_{i}^{*}]_{\sim}=\mu_{i}^{-1}\sum_{\mu_{j}>0}\mu_{j}\bigl\langle[e_{i}^{*}]_{\sim},[e_{j}^{*}]_{\sim}\bigr\rangle_{{L_{2}(\lambda)}}e_{j}^{*}=e_{i}^{*}\,$,
where in the second step we used [47, Theorem 2.11] in combination with (50).
In addition, since $k_{X}$ is continuous, every $h\in H_{X}$ is continuous. Moreover, the Haar measure satisfies $\lambda(O)>0$ for all non-empty and open $O\subset G$. Together, these properties show that $\lambda$ is $H_{X}$-positive.
Using all these properties, the characterization now follows by combining Corollary 3.3 and Theorem 3.9 as discussed after Theorem 3.9. In addition, $\sup_{i\in I}\|e_{i}^{*}\|_{\infty}\leq\sqrt{2}$ in combination with $e_{i}^{*}\in C(T)$ and the summability assumption on the eigenvalues ensures that (8) converges absolutely in $C(T\times T)$, and thus $k_{X}^{1/2}$ is continuous. This in turn implies $H_{X}^{1/2}\subset C(T)$. ∎